### v1

# Normal forms for ordinary differential operators, I

Abstract
In this paper we develop the generalised Schur theory offered in the recent paper by the second author in dimension one case, and apply it to obtain two applications in different directions of algebra/algebraic geometry.
The first application is a new explicit parametrisation of torsion free rank one sheaves on projective irreducible curves with vanishing cohomology groups.
The second application is a commutativity criterion for operators in the Weyl algebra or, more generally, in the ring of ordinary differential operators, which we prove in the case when operators have a normal form with the restriction top line (for details see Introduction).
Both applications are obtained with the help of normal forms. Namely, considering the ring of ordinary differential operators $D_{1}=K[[x]][\partial]$ as a subring of a certain complete non-commutative ring $\hat{D}_{1}^{sym}$, the normal forms of differential operators mentioned here are obtained after conjugation by some invertible operator (”Schur operator”), calculated using one of the operators in a ring. Normal forms of commuting operators are polynomials with constant coefficients in the differentiation, integration and shift operators, which have a finite order in each variable, and can be effectively calculated for any given commuting operators.
1 Introduction
In this paper we develop the generalised Schur theory offered in [42] in dimension one case and apply it to obtain two applications in different directions of algebra/algebraic geometry.
The first application is a new convenient explicit parametrisation of torsion free rank one sheaves on projective irreducible curves with vanishing cohomology groups. It is well known (see e.g. [4] or review [43]) that such sheaves are exactly the spectral sheaves of commutative subrings of ordinary differential operators of rank one. This result is the first step in establishing similar parametrisation for spectral sheaves of arbitrary rank, and also for spectral sheaves of commutative subrings of operators in higher dimensions (cf. [5], [41]). It is motivated by an important problem that appears in algebraic-geometric classification of commutative subrings of operators in higher dimensions – a description of torsion free sheaves with specific conditions on cohomology groups (see [41]), in particular with fixed Hilbert polynomial and some extra conditions. In the work [5] a description of Cohen-Macaulay sheaves on the spectral surface of quantum Calogero-Moser systems was given with the help of matrix problem approach due to I. Burban and Y. Drozd [2], [3] (Cohen-Macaulay sheaves form an open part of the moduli space of torsion free sheaves with fixed Hilbert polynomial, and by that reason it is important to describe them at first), however this approach meets with a difficulty to describe sheaves with specific cohomological properties (cf. [5, Sec. 6]). We expect that our new approach will help to solve this problem in an effective way in any dimension.
The second application is a commutativity criterion for operators in the Weyl algebra or, more generally, in the ring of ordinary differential operators. It is motivated by the following natural question from the Burchnall-Chaundy theory. The famous Burchnall-Chaundy lemma ([6]) says that if $P,Q\in D_{1}$ are differential operators of coprime orders $p$ and $q$, then they are algebraically dependent, and satisfy a polynomial relation of the form $F(X,Y)=\alpha X^{q}\pm Y^{p}+\ldots=0$, where $GCD(p,q)=1$ and $\lambda\neq 0$ (here the weighted degree $v_{q,p}$ from [10] of $F(X,Y)$ is $pq$, and $\ldots$ mean terms of lower weighted degree, where the weight of $X$ is $p$, and the weight of $Y$ is $q$). [^2] A similar result for operators of rank $r$ was established in [38], cf. [31, Th. 2.11] Vice versa, if $P,Q\in D$ is a solution of such polynomial $f(X,Y)$ [^3] A solution of the equation $f(X,Y)=0$ is a pair $(P,Q)\in D$ such that $\sum_{i,j=0}^{n}\alpha_{ij}P^{i}Q^{j}=0$., then $[P,Q]=0$. Now a natural question whether $F(P,Q)=0\Rightarrow[P,Q]=0$ for generic polynomial $F$ appears. This question appears to be surprisingly difficult in general case. We give a partial affirmative answer on this question in the case when the normal form has the restriction top line (see discussion below).
The necessity of further development of the Schur theory was not restricted only by the above mentioned applications. Let’s recall two major theorems of this theory from [42] (we formulate them here in a simplified form):
Theorem 1.1 (A generalized Schur theorem, Th. 7.1 ).
Let $P_{1},\ldots,P_{n}\in\hat{D}_{n}^{sym}$ be commuting operators with $\mathop{\rm\bf ord}(P_{i})=k$ for all $i=1,\ldots,n$. Assume that the module $F$ of the ring $K[\sigma(P_{1}),\ldots,\sigma(P_{n})]$ is finitely generated and free.
Then there exists an invertible operator $S\in\hat{D}_{n}^{sym}$ with $\mathop{\rm\bf ord}(S)=0$ such that
$S^{-1}\partial_{i}^{k}S=P_{i},\quad i=1,\ldots,n.$
If $n=1$ then the conditions of the generalized Schur theorem are automatically satisfied for any monic operator $P\in\hat{D}_{1}^{sym}$. So, $P=S\partial^{q}S^{-1}$ for some $S\in\hat{D}_{1}^{sym}$.
Theorem 1.2 (A centralizer theorem, Prop.7.1 ).
Assume $[\partial_{q}^{k},Q]=0$ for $q=1,\ldots,n$, where $Q\in\hat{D}_{n}^{sym}$. Then
$Q=\sum_{j_{1}=0}^{k-1}\ldots\sum_{j_{n}=0}^{k-1}c_{j_{1},\ldots,j_{n}}A_{k;j_{1},1}\ldots A_{k;j_{n},n},$
where all coefficients $c_{j_{1},\ldots,j_{n}}$ are polynomials in $\partial_{q}$, $\int_{q}$, $q=1,\ldots,n$ with constant coefficients and the degree of these polynomials with respect to $\partial_{q}$ is not greater than $\mathop{\rm\bf ord}(Q)$ and the degree of these polynomials with respect to $\int_{q}$ is not greater than $k-1$.
If the operators $P_{i}$ from the first theorem are differential, i.e. $P_{i}\in D_{n}$, a natural question appears: what is the shape of the operator $S$? In dimension one case it is well known (see e.g. [26], [18], [34] or the book [43] and references therein) that, if the centraliser of $P_{1}$ is non-trivial, then $S$ can be expressed via the Baker-Akhieser function (and vice versa), though in quite non-trivial way, if the rank of the centraliser is greater than one. Starting from works [14], [15], [16] it is known that the Baker-Akhieser function plays an important role in many problems of mathematical physics, in particular it played a key role in constructing explicit examples of commuting operators in many works (cf. [25], [24]). Analogously, in higher dimension the Schur-Sato operator $S$ determines the common eigenfunction of commuting operators (of different nature), cf. [5, Sec. 6], [41], and the knowledge of its shape could help to prove the classification conjecture [41, Conj. 7.11] about commuting partial differential operators. Besides, in higher dimension, the operator $S$ determines an order-preserving endomorphism of the Weyl algebra ([42, Cor. 2.1]), thus giving hint to the Dixmier and Jacobian conjectures.
Analysing the shape of the Schur operator $S$, we find that all its homogeneous components are non-commutative polynomials with constant coefficients in the differentiation, integration, shift operators $A_{i}$ (see below) and the operator $\Gamma:=x\partial$ (we call such polynomials as HCPC for short), which have a finite order in each variable and additionally satisfy a specific property of being totally free of $B_{j}$ (see definition 2.6). All HCPC form a subring $Hcpc(k)$, which occasionally looks very similar to the algebra of polynomial integro-differential operators studied in the paper [1] (though the shift operators are not included into this algebra). We establish estimates on the degrees of this polynomials in section 2.4. As a result we encode all necessary properties of the operator $S$ in a condition $A_{q}(k)$ of section 2.5 (see definition 2.8 or the list of notations below). With the help of this condition, we gave a criterion of an operator $P\in\hat{D}_{1}^{sym}$ to be a differential operator (see theorem 2.2).
A normal form of a pair of operators $P,Q\in\hat{D}_{1}^{sym}$ is a pair $P^{\prime},Q^{\prime}\in\hat{D}_{1}^{sym}$ obtained after conjugation by a Schur operator $S$ as above, calculated using one of the operators in a pair $(P,Q)$ (or, more generally, in the ring $\hat{D}_{1}^{sym}$, see definitions 2.7, 3.6 and remark 2.7). The normal form is not uniquely defined, but up to conjugation with invertible $S\in\hat{D}_{1}^{sym}$ from the centralizer $C(\partial^{k})$ with $\mathop{\rm\bf ord}(S)=0$. By centralizer theorem 1.2 such $S$ is a polynomial of restricted degree. Notably, the whole centraliser $C(\partial^{k})$ is naturally isomorphic to a matrix $k\times k$ algebra over a polynomial ring, see remark 3.3. The normal form of commuting operators can be normalized in some way (see section 3). By the centralizer theorem $C(\partial^{k})$ consists of (non-commutative) polynomials, so normalised normal form can be calculated for any such operators. If the operators do not commute, the normal forms will be series in general, for which, however, it is possible to calculate any given number of terms. For a pair of differential operators normal forms satisfy condition $A_{q}(0)$ (see corollary 2.4).
Normal forms of a commutative subring $B\subset D_{1}$ appear to be a very effective tool to describe the moduli space of spectral sheaves, i.e. torsion free sheaves on the spectral curve with certain conditions on cohomology groups, cf. [4, §1.3] and theorem 3.1 below. Roughly speaking, the set of coefficients of a normalised normal form determines such a sheaf up to an isomorphism. This set depends on a choice of normalisation, and can be thought of as a system of local coordinates on a chart of a manifold – the moduli space of spectral sheaves. Precise statement about this parametrisation in case of sheaves of rank one is formulated in theorem 3.3.
To study normal forms of non-commuting operators we develop a technique of Newton regions (see section 4) – this is a natural generalisation of the technique of Newton polygons widely used for study of operators in the Weyl algebra (cf. [10], [13], [19]). Since normal forms of non-commuting operators are usually infinite series, the convex hull of all monomials may not be a restricted domain. However, in this case it is possible to define relevant notions of weights and top lines (generalisations of corresponding notions from [10]). In section 4 we study normal forms of a pair of non-commuting monic differential operators $P,Q\in D_{1}$. After conjugating this pair by a Schur operator of, say, operator $Q$, we obtain a monic operator $P^{\prime}\in\hat{D}_{1}^{sym}$ satisfying condition $A_{q}(0)$ (where $q$ is the order of $Q$). It is possible to define a weight function $v_{\sigma,\rho}$ and a notion of related top line for such operators. We distinguish 2 principal cases of top lines: the restriction top line and the asymptotic top line, both lines are uniquely defined (see definitions 4.3 and 4.4). Lemma 4.1 says that there are only two possibilities for a non commuting with $\partial^{q}$ operator $P^{\prime}$: it has either the restriction top line or the asymptotic top line. In section 4.3 we give the affirmative answer on the question whether $F(P,Q)=0\Rightarrow[P,Q]=0$ in the case when the normal form $P^{\prime}$ of the pair $P,Q$ has the restriction top line.
We will consider the remaining case of the asymptotic top line as well as a similar description of the moduli space for sheaves of rank $>1$ in the next article, since these cases require much more detail. We expect that further study of normal forms (both for commuting and non-commuting operators) is reasonable not only for differential operators, but also for operators of other type, like difference, integro-differential, etc. In particular, it seems to be promising to study normal forms of recently discovered examples from [21], [20].
The structure of this article is the following. In section 2 we develop generic theory of normal forms for ordinary differential operators. Namely, in section 2.1 we review the Schur theory from [42] in the case of dimension one, strengthening some specific statements useful in the rest of the paper. In section 2.2 we deduce a list of useful formulae, in section 2.3 we introduce an important technical notion of homogeneous canonical polynomials (HCP) and study their basic properties. This section contains important estimations and formulae useful for fast explicit calculations of concrete examples of normal forms and Schur operators. In section 2.4 we develop the Schur theory further by studying necessary conditions on the Schur conjugating operators for ordinary differential operators. In section 2.5 we introduce the main subject of this section – normal forms for differential operators and study basic properties of them. The major result of this section is theorem 2.2, a criterion of an operator from $\hat{D}_{1}^{sym}$ to be a differential operator.
In section 3 we study normal forms of commuting differential operators with the help of theory from section 2. We give a convenient description of the centraliser and of normal forms of a pair of commuting operators. With the help of this description we give a new parametrisation of torsion free sheaves of rank one with vanishing cohomology groups on a projective curve.
In section 4 we study normal forms of non-commuting differential operators. In section 4.1 we introduce the notion of Newton region – a natural generalisation of the Newton polygon – for operators from $\hat{D}_{1}^{sym}$ and study its basic properties for operators satisfying condition $A_{k}(0)$ (all normal forms of differential operators satisfy this condition). In section 4.2 we prove one general combinatorial lemma, and in section 4.3 we prove the main theorem of section 4 – a commutativity criterion of a pair of operators in the case when the normal form of this pair has the restriction top line.
In Appendix we collect all necessary basic technical assertions about the weight function $v_{\sigma,\rho}$ and the homogeneous highest terms $f_{\sigma,\rho}$ used in section 4, with detailed proofs.
Acknowledgements. A. Zheglov is grateful to the Sino-Russian Mathematics Center at Peking University for hospitality and excellent working conditions while preparing this paper.
1.1 List of notations
Since this work uses quite different techniques, for convenience of the reader we introduce now the most important notations used in this paper.
  1. 1.
    $K$ is a field of characteristic zero. Recall some notation from [42]: $\hat{R}:=K[[x_{1},\ldots,x_{n}]]$, the $K$-vector space
[…]
    for any element $0\neq P:=\sum\limits_{\underline{k}\geq\underline{0}}a_{\underline{k}}\underline{\partial}^{\underline{k}}\in{{\cal M}}_{n}$
[…]
    $\hat{D}_{n}^{sym}:=\bigl{\{}Q\in{{\cal M}}_{n}\,\big{|}\,\mathop{\rm\bf ord}(Q)<\infty\bigr{\}};$
    $P_{m}:=\sum\limits_{|\underline{i}|-|\underline{k}|=m}\alpha_{\underline{k},\underline{i}}\,\underline{x}^{\underline{i}}\underline{\partial}^{\underline{k}}$ – the $m$-th homogeneous component of $P$,
    $\sigma(P):=P_{\mathop{\rm\bf ord}(P)}=P_{-d}$ – the highest symbol.
  2. 2.
    In this paper we use: $\hat{R}:=K[[x]]$, $D_{1}:=\hat{R}[\partial]$,
    $\hat{D}_{1}^{sym}:=\{Q=\sum_{k\geq 0}a_{k}\partial^{k}|\mathop{\rm\bf ord}(Q)<\infty\}.$
    Operators: $\delta:=\exp((-x)\ast\partial)$, $\int:=(1-exp((-x)\ast\partial))\cdot\partial^{-1}$, $A_{k;i}:=\exp((\xi^{i}-1)x\ast\partial)\in\hat{D}^{sym}_{1}\hat{\otimes}_{K}\tilde{K}$ (in the case when $k$ is fixed, simply written as $A_{i}$), $\Gamma_{i}=(x\partial)^{i}$. $B_{n}=\frac{1}{(n-1)!}x^{n-1}\delta\partial^{n-1}$.
    Operators written in the (Standard) form as
    $H=[\sum_{0\leq i<k}f_{i;r}(x,A_{k;i},\partial)+\sum_{0<j\leq N}g_{j;r}B_{j}]D^{r}$
    are called HCP and form a sub-ring $Hcpc(k)$. They can be written also in G-form:
    $H=(\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f^{\prime}_{l,i;r}\Gamma_{l}A_{i}+\sum_{0<j\leq N}g_{j;r}B_{j})D^{r}$
    The $A$ and $B$ Stable degrees of HCP are defined as
    $Sdeg_{A}(H)=\max\{d_{i}|\quad 0\leq i<k\}\quad\mbox{or $-\infty$, if all $f_{l,i;r}=0$ }$
    and
    $Sdeg_{B}(H)=\max\{j|\quad g_{j;r}\neq 0\}\quad\mbox{or $-\infty$, if all $g_{j;r}=0$}$
    We denote $D^{i}=\partial^{i}$ if $i\geq 0$ and $\int^{i}$ if $i<0$.
    In the case when $Sdeg_{B}(HD^{p})=-\infty,\forall p\in\mathbf{Z}$ $H$ is called totally free of $B_{j}$.
[…]
    1. (a)
      $P_{t}$ is a HCP from $Hcpc(q)$ for all $t$;
    2. (b)
      $P_{t}$ is totally free of $B_{j}$ for all $t$;
    3. (c)
      $Sdeg_{A}(P_{\mathop{\rm\bf ord}(P)-i})<i+k$ for all $i>0$;
    4. (d)
      $\sigma(P)$ does not contain $A_{q;i}$, $Sdeg_{A}(\sigma(P))=k$.
  3. 3.
    In section 3, $\mathfrak{B}={{\cal R}}_{S}$ is the right quotient ring of ${{\cal R}}=\tilde{K}^{\oplus k}[D,\sigma]$ by $S=\{D^{k}|k\geq 0\}$. And the ring of skew pseudo-differential operators
    $E_{k}:=\tilde{K}[\Gamma_{1},A_{1}]((\tilde{D}^{-1}))=\{\sum_{l=M}^{\infty}P_{l}\tilde{D}^{-l}|\quad P_{l}\in\tilde{K}[\Gamma_{1},A_{1}]\}\simeq\tilde{K}^{\oplus k}[\Gamma_{1}]((\tilde{D}^{-1}))$
    $\widehat{Hcpc}_{B}(k)$ is the $\tilde{K}$-subalgebra in $\hat{D}_{1}^{sym}\hat{\otimes}\tilde{K}$ consisting of operators whose homogeneous components are HCPs totally free of $B_{j}$.
    Suppose $B$ is a commutative sub-algebra of $D_{1}$, then $(C,p,{{\cal F}})$ stands for the spectral data of $B$ (the spectral curve, point at infinity and the spectral sheaf with vanishing cohomologies).
  4. 4.
    In section 4, suppose $H$ is an operator whose components are all HCP. Then $E(H)$ denotes the point set where $f_{l,i;r}\neq 0$, $v_{\sigma,\rho}$ stands for the weight degree of $H$, and $f_{\sigma,\rho}$ for the highest terms associated to $(\sigma,\rho)$:
    $v_{\sigma,\rho}(H)=\sup\{\sigma l+\rho j|(l,j)\in E(H)\}\quad f_{\sigma,\rho}(H)=\sum_{(l,j)\in E(H,\sigma,\rho)}\sum_{i}f_{l,i;j}\Gamma_{l}A_{k,i}D^{j}$
    The up-edge of the Newton region of $P$ is the set
    $Edg_{u}(P):=\{(a,b)\in E(P)|\quad a=Sdeg_{A}(P_{b})\mbox{\quad and\quad}\forall b^{\prime}>b\quad Sdeg_{A}(P_{b^{\prime}})<a\}.$
    And $H_{d;(\sigma,\rho)}(H),HS_{d;(\sigma,\rho)}^{m}(H)$ stands for
    $H_{d;(\sigma,\rho)}(L):=\sum_{\sigma l+\rho j\geq d}\sum_{i=0}^{k-1}\alpha_{l,i;j}\Gamma_{l}A_{i}\partial^{j}$
    and
    $HS_{d;(\sigma,\rho)}^{m}(L):=\sum_{\sigma l+\rho j\geq d;l\leq m}\sum_{i=0}^{k-1}\alpha_{l,i;j}\Gamma_{l}\partial^{j}$
2 Generic theory of normal forms
2.1 Preliminary statements from the Schur theory for the ring $\hat{D}_{1}^{sym}$
Suppose $K$ is a field of characteristic zero.
Following the notations in [42], denote $\hat{R}:=K[[x]]$, $D_{1}:=\hat{R}[\partial]$, define the $K$-vector space
$\mathcal{M}_{1}:=\hat{R}[[\partial]]=\{\sum_{k\geq 0}a_{k}\partial^{k}|a_{k}\in\hat{R}\quad\forall k\in\mathbb{N}_{0}\},$
where $v:\hat{R}\rightarrow\mathbb{N}_{0}\cup\infty$ is the discrete valuation defined by the unique maximal ideal of $\hat{R}$, for any element $0\neq P:=\sum_{k\geq 0}a_{k}\partial^{k}\in\mathcal{M}_{1}$ define the order function
${}.\mathop{\rm\bf ord}(P):=\sup\{k-v(a_{k})|\quad k\in\mathbb{N}_{0}\}\in\mathbb{Z}\cup\{\infty\}.$
Define the space
$\hat{D}^{sym}_{1}:=\{Q\in\mathcal{M}_{1}|\quad\mathop{\rm\bf ord}(Q)<\infty\}.$
By definition, any element $P\in\hat{D}^{sym}_{1}$ is written in the canonical form
$P:=\sum_{k-i\leq\mathop{\rm\bf ord}(P)}\alpha_{k,i}x^{i}\partial^{k}.$
We call $P_{m}:=\sum_{k-i=m}\alpha_{k,i}x^{i}\partial^{k}$ as the $m$-th homogeneous component of $P$, we call $\sigma(P):=P_{\mathop{\rm\bf ord}(P)}$ as the highest symbol of $P$. Then we have the (uniquely defined) homogeneous decomposition for any $P\in\hat{D}^{sym}_{1}$:
$P=\sum_{m=-\infty}^{\mathop{\rm\bf ord}}P_{m}.$
Denote by $A_{1}:=K[x][\partial]$ the first Weyl algebra. Clearly, $A_{1}\subset D_{1}\subset\hat{D}_{1}^{sym}$.
Remark 2.1.
Note that the order function $\mathop{\rm\bf ord}$ coincide with the weight function $v_{1,-1}$ on the ring $A_{1}$ from the paper [10]. For any operator $P\in D_{1}$, we define the usual order (or degree) of $P$ as $\deg(P):=v_{0,1}(P)$. If $P\in D_{1}$ has an invertible highest coefficient (with respect to the usual order), then it is easy to see that $\deg(P)=\mathop{\rm\bf ord}(P)$.
Theorem 2.1.
([42],Theorem 2.1) The following statement are properties of $\hat{D}^{sym}_{1}$
  1. 1.
    $\hat{D}^{sym}_{1}$ is a ring (with natural operations $\cdot$ ,+ coming from $D_{1}$); $\hat{D}^{sym}_{1}\supset D_{1}$.
  2. 2.
    $\hat{R}$ has a natural structure of a left $\hat{D}^{sym}_{1}$-module, which extends its natural structure of a left $D_{1}$-module.
  3. 3.
    We have a natural isomorphism of $K$-vector spaces
    $F:=\hat{D}^{sym}_{1}/\mathfrak{m}\hat{D}^{sym}_{1}\rightarrow K[\partial]$
    where $\mathfrak{m}=(x)$ is maximal ideal of $\hat{R}$.
  4. 4.
    Operators from $\hat{D}^{sym}_{1}$ can realise arbitrary endomorphisms of the $K$-algebra $\hat{R}$ which are continuous in the $\mathfrak{m}$-adic topology.
  5. 5.
    There are Dirac delta functions, operators of integration, difference opertors.
The Dirac delta is given by the series $\delta:=exp((-x)\ast\partial):=1-x\partial+\frac{1}{2!}x^{2}\partial^{2}-\ldots$, which satisfies $\delta\circ f(x)=f(0)$, and the operator of integration is given by the series
$\int:=(1-exp((-x)\ast\partial))\cdot\partial^{-1}=\sum_{k=0}^{\infty}\frac{x^{k+1}}{(k+1)!}(-\partial)^{k}$
which satisfies
$\int\circ x^{m}=\frac{x^{m+1}}{m+1}.$
Note that $\int$ is only the right inverse of $\partial$, because $\partial\int=1$ and $\int\cdot\partial=((1-exp((-x)\ast\partial))\cdot\partial^{-1}\cdot\partial)=1-\delta$.
Remark 2.2.
Unlike the usual ring of PDOs the ring $\hat{D}_{1}^{sym}$ contains zero divisors. There are the following obvious properties of the order function (cf. the proof of [5, Th. 5.3]):
  1. 1.
    $\mathop{\rm\bf ord}(P\cdot Q)\leq\mathop{\rm\bf ord}(P)+\mathop{\rm\bf ord}(Q)$, and the equality holds if $\sigma(P)\cdot\sigma(Q)\neq 0$,
  2. 2.
    $\sigma(P\cdot Q)=\sigma(P)\cdot\sigma(Q)$, provided $\sigma(P)\cdot\sigma(Q)\neq 0$,
  3. 3.
    $\mathop{\rm\bf ord}(P+Q)\leq\max\{\mathop{\rm\bf ord}(P),\mathop{\rm\bf ord}(Q)\}$.
In particular, the function $-\mathop{\rm\bf ord}$ determines a discrete pseudo-valuation on the ring $\hat{D}_{1}^{sym}$.
Definition 2.1.
Let $\xi$ be a primitive $k$-th root of unity, $\tilde{K}=K[\xi]$. For any $i\in\mathbb{Z}$, we define operators
$A_{k;i}:=\exp((\xi^{i}-1)x\ast\partial)\in\hat{D}^{sym}_{1}\hat{\otimes}_{K}\tilde{K},$
where $\hat{D}^{sym}_{1}\hat{\otimes}_{K}\tilde{K}$ means the same ring $\hat{D}^{sym}_{1}$, but defined over the base field $\tilde{K}$.
Further, if it will be clear from the context, we’ll omit index $k$ and use notation $A_{i}:=A_{k;i}$.
Lemma 2.1.
(cf. [42],Lemma 7.2) The sum
$A:=c_{0}+c_{1}A_{k,1}+\dots+c_{k-1}A_{k;k-1},\quad c_{i}\in\tilde{K}$
is equal to zero iff $c_{i}=0,i=0,\dots,k-1$. If it is not equal to zero, then it is of order zero.
Moreover, $A$ is a polynomial in $\partial$ iff $c_{1}=\ldots=c_{k-1}=0$.
Proof.
The first part of lemma coincides with [42, Lem. 7.2]. The last assertion follows easily from the proof of [42, Lem. 7.2]. Namely, $A$ is a polynomial iff the infinite system of linear equations hold
[…]
for an appropriate $n_{0}$. But by the well known property of the Vandermonde matrix this system has the unique solution $c_{1}=\ldots=c_{k-1}=0$.
The following claim is a partial case of [42, Prop. 7.1].
Proposition 2.1.
In the notation of definition 2.1 we have
  1. 1.
    For any $i,j,p$, we have
    $A_{i}A_{j}=A_{i+j}$
    $\partial^{p}A_{i}=\xi^{pi}A_{i}\partial^{p}$
    $A_{i}x^{p}=\xi^{pi}x^{p}A_{i}$
    $\int^{p}A_{i}=\xi^{-pi}A_{i}\int^{p}$
  2. 2.
    For a given $Q\in\hat{D}^{sym}_{1}$ assume that $[\partial^{k},Q]=0$.
    Then $Q=c_{0}+c_{1}A_{1}+\dots+c_{k-1}A_{k-1}\in\hat{D}^{sym}_{1}\hat{\otimes}_{K}\tilde{K}$, where $c_{i}\in\hat{D}^{sym}_{1}\hat{\otimes}_{K}\tilde{K}$ are given by the following formula:
    $c_{i}=\sum_{m=0}^{\mathop{\rm\bf ord}(Q)}c_{i,m}\partial^{m}+c_{i,-1}\int+\dots+c_{i,-k+1}\int^{k-1}$
    where $c_{i,j}\in\tilde{K}$ (so that $\mathop{\rm\bf ord}(c_{i})\leq\mathop{\rm\bf ord}(Q)$). Besides, the coefficients $c_{i,j}$ are uniquely defined.
Proof.
For convenience of the reader we’ll give here a proof of item 2 which is easier in our case that the proof of [42, Prop. 7.1] (and we’ll use some of its arguments later).
The identity $[\partial^{k},Q]=0$ can be rewritten as
$\sum_{i=1}^{k}\binom{k}{i}\partial^{i}(Q)\partial^{k-i}=0.$ (2.1)
Note that any solution $Q\in\hat{D}_{1}^{sym}$ of this equation gives a solution $Q^{\prime}\in\tilde{K}[[x,\tilde{\partial}]]$ of the equation
$\sum_{i=1}^{k}\binom{k}{i}\partial^{i}(Q^{\prime})\tilde{\partial}^{k-i}=0,$
where $\tilde{\partial}$ means a new formal variable (commutative with $x$). Namely, we just replace $\partial$ by $\tilde{\partial}$ in the series $Q$.
On the other hand, the last equation can be written in the form
$\prod_{i=1}^{k}(\partial+(1-\xi_{k}^{i})\tilde{\partial})(Q^{\prime})=0.$
Any solution of the last equation in the commutative ring $\tilde{R}:=\tilde{K}[[x]]((\tilde{\partial}))$ has the form
$Q^{\prime}=c_{0}+c_{1}\exp((\xi_{k}-1)x\tilde{\partial})+\ldots+c_{k-1}\exp((\xi_{k}^{k-1}-1)x\tilde{\partial}),$
where $c_{i}\in\tilde{R}$ don’t depend on $x$ (as it follows from elementary differential algebra) [^4] First note that $\ker\{\partial:\tilde{R}\rightarrow\tilde{R}\}=\tilde{K}((\tilde{\partial}))$. Suppose there is a solution $H$ such that it is linearly independent with $1,\exp((\xi_{k}-1)x\tilde{\partial}),\ldots,\exp((\xi_{k}^{k-1}-1)x\tilde{\partial})$ over $\tilde{K}((\tilde{\partial}))$. Then $H_{0}:=\partial(H)=H^{\prime}$ is not equal to zero and is linearly independent with $\exp((\xi_{k}-1)x\tilde{\partial}),\ldots,\exp((\xi_{k}^{k-1}-1)x\tilde{\partial})$ over $\tilde{K}((\tilde{\partial}))$. By induction, $H_{i}:=(\partial+(1-\xi_{k}^{i})\tilde{\partial})(H_{i-1})=H_{i-1}^{\prime}+(1-\xi_{k}^{i})\tilde{\partial}H_{i-1}$ is not equal to zero and is linearly independent with $\exp((\xi_{k}^{i+1}-1)x\tilde{\partial}),\ldots,\exp((\xi_{k}^{k-1}-1)x\tilde{\partial})$ over $\tilde{K}((\tilde{\partial}))$ for all $i\leq k-1$, in particular $H_{k-1}\neq 0$. On the other hand, $H_{k-1}=\prod_{i=1}^{k}(\partial+(1-\xi_{k}^{i})\tilde{\partial})(H)=0$, a contradiction..
If we choose a canonical representation form of elements in $\tilde{R}$ such that each monomial has the form $x^{j}\tilde{\partial}^{q}$ (with $x$ on the left and $\tilde{\partial}$ on the right), then the right hand side of the last formula can be rewritten as
$Q^{\prime\prime}:=c_{0}+\exp((\xi_{k}-1)x\tilde{\partial})c_{1}+\ldots+\exp((\xi_{k}^{k-1}-1)x\tilde{\partial})c_{k-1}=c_{0}+\tilde{A}_{1}c_{1}+\ldots+\tilde{A}_{k-1}c_{k-1}.$ (2.2)
Then $Q^{\prime}=Q^{\prime\prime}$ also as elements written in this representation. Note that, since $Q^{\prime}$ contains only non-negative powers of $\partial$, the series $Q^{\prime\prime}$ in (2.2) contains only non-negative powers of $\partial$ (and if we replace $\tilde{\partial}$ by $\partial$ in all terms of $Q^{\prime\prime}$, we get again the operator $Q$).
Lemma 2.2.
For any $i=0,\ldots,k-1$ we have $\mathop{\rm\bf ord}(c_{i})\leq\mathop{\rm\bf ord}(Q)$ in formula (2.2), where the order $\mathop{\rm\bf ord}$ is defined on $\tilde{R}$ in the same way as on ^$\hat{{{\cal M}}}_{1}$.
Proof.
Since the elements $c_{i}$ are polynomials in ~$\tilde{\partial}^{-1}$, the expression in (2.2) can be written in the form
$(\tilde{c}_{0}+\tilde{A}_{1}\tilde{c}_{1}+\ldots+\tilde{A}_{k-1}\tilde{c}_{k-1})\tilde{\partial}^{-m},$
where $\tilde{c}_{i}\in\tilde{K}[[\tilde{\partial}]]$ and $m\geq 0$, i.e. the series in brackets is divisible by ~$\tilde{\partial}^{m}$. We’ll additionally assume that $m$ is minimal, i.e. $GCD(\tilde{c}_{0},\ldots,\tilde{c}_{k-1})=1$ in the ring $\tilde{K}[[\tilde{\partial}]]$.
Obviously, the homogeneous decomposition is unique also in the space $\tilde{R}$, and therefore in (2.2) we have the unique homogeneous decomposition
$c_{0}+\tilde{A}_{1}c_{1}+\ldots+\tilde{A}_{k-1}c_{k-1}=\sum_{l\geq-m}(c_{0,l}+c_{1,l}\tilde{A}_{1}+\ldots+c_{k-1,l}\tilde{A}_{k-1})\tilde{\partial}^{l},$
where $c_{i,j}\in\tilde{K}$. Since $\mathop{\rm\bf ord}(Q)<\infty$ and $\mathop{\rm\bf ord}(A_{k;i})=0$, we should have
$\mathop{\rm\bf ord}(c_{0}+\tilde{A}_{1}c_{1}+\ldots+\tilde{A}_{k-1}c_{k-1})\leq\mathop{\rm\bf ord}(Q)$
and therefore by lemma 2.1 $c_{i,l}=0$ for all $l>\mathop{\rm\bf ord}(Q)$ and all $i=0,\ldots,k-1$. This means that $\mathop{\rm\bf ord}(c_{i})\leq\mathop{\rm\bf ord}(Q)$ for any $i$. ∎
From lemma 2.2 it follows that the series $\tilde{c}_{i}$ will belong to $\hat{D}_{1}^{sym}$ after replacing $\tilde{\partial}$ by $\partial$ in all terms. Now note that in $\hat{D}_{1}^{sym}$ we have
$(\tilde{c}_{0}+\tilde{A}_{1}\tilde{c}_{1}+\ldots+\tilde{A}_{k-1}\tilde{c}_{k-1})\tilde{\partial}^{-m}|_{\tilde{\partial}\mapsto\partial}=(\tilde{c}_{0}+\tilde{A}_{1}\tilde{c}_{1}+\ldots+\tilde{A}_{k-1}\tilde{c}_{k-1})|_{\tilde{\partial}\mapsto\partial}\int^{m}$
i.e. $Q$ can be written in the form:
$Q=(\tilde{c}_{0}+\tilde{A}_{1}\tilde{c}_{1}+\ldots+\tilde{A}_{k-1}\tilde{c}_{k-1})|_{\tilde{\partial}\mapsto\partial}\int^{m}.$ (2.3)
Besides, all summands in the sum (2.3) are well defined elements of $\hat{D}_{1}^{sym}$ of order $\leq\mathop{\rm\bf ord}(Q)$ and their sum is also well defined in $\hat{D}_{1}^{sym}$.
Lemma 2.3.
In the formula (2.3) we have $m\leq k-1$.
Proof.
Assume the converse: $m\geq k$. By our assumption, the homogeneous component $Q_{-m}\neq 0$, and
$Q_{-m}=(c_{0,-m}+c_{1,-m}{A}_{1}+\ldots+c_{k-1,-m}{A}_{k-1})\int^{m}$
Since $[\partial^{k},Q]=0$, we have also
$0=[\partial^{k},Q_{-m}]=\partial^{k}(Q_{-m})+\sum_{i=1}^{k-1}\binom{k}{i}\partial^{k-i}(Q_{-m})\partial^{i}.$
Let the canonical form of $Q_{-m}$ be $Q_{-m}=\sum_{p\geq 0}a_{m+p}x^{m+p}\partial^{p}$. Let $a_{m+z}$ be the first coefficient not equal to zero. Then we have
$\partial^{k}(Q_{-m})+\sum_{i=1}^{k-1}\binom{k}{i}\partial^{k-i}(Q_{-m})\partial^{i}=\\ a_{m+z}\frac{(m+z)!}{(m+z-k)!}x^{m+z-k}\partial^{z}+\sum_{p>0}a_{m+z+p}^{\prime}x^{m+z-k+p}\partial^{z+p}\neq 0$
(here $a_{j}^{\prime}\in\tilde{K}$ are appropriate coefficients) – a contradiction. ∎
Now, expanding all brackets in (2.3) and using the identities from item 1, we can rewrite $Q$ in the form stated in item 2. The uniqueness of coefficients follows immediately from lemma 2.1.
Let’s recall one notation and definition from [42]. For any $P\in\hat{D}_{1}^{sym}$ we put
$P_{[q]}:=\frac{x^{q}}{q!}P_{(q)},\mbox{ where}$
[…]
The expression $P_{(q)}$ is called a slice and the sum $P=\sum_{q\geq 0}P_{[q]}$ is called (a partial slice decomposition).
Consider the space $F=K[\partial]$. It has a natural structure of a right $\hat{D}_{1}^{sym}$-module via the isomorphism of vector spaces $F\simeq\hat{D}_{1}^{sym}/\mathfrak{m}\hat{D}_{1}^{sym}$.
Definition 2.2.
(cf. [42], Def. 6.4) An element $P\in\hat{D}_{1}^{sym}$ is called regular if the $K$–linear map $F\xrightarrow{(-\circ\sigma(P))}F$ is injective, where $\circ$ means the action on the module $F$. In particular, $P$ is regular if and only if its symbol $\sigma(P)$ is regular.
Proposition 2.2.
([42],Proposition 7.2) Let $P\in\hat{D}^{sym}_{1}$, $\mathop{\rm\bf ord}(P)=k>0$ be a regular operator. Then there exists an invertible operator $S\in\hat{D}^{sym}_{1}$ with $\mathop{\rm\bf ord}(S)=0$ such that
$P=S^{-1}\partial^{k}S$
and $S_{[0]}=1$, $S_{[i]}=0$ for $0<i<k$.
Example 2.1.
It’s easy to see that an operator $P\in D_{1}$ with an invertible highest coefficient (cf.remark 2.1) is an example of a regular operator.
Recall that any such operator can be normalised, i.e. reduced to the form $P=\partial^{k}+c_{k-2}\partial^{k-2}+\ldots+c_{0}$, with the help of some change of variables and conjugation by invertible function, see e.g. [4, Prop. 1.3 and Rem. 1.6].
Proposition 2.3.
Let $P\in D_{1}$ be a normalized operator of positive degree, i.e. $P=\partial^{k}+c_{k-2}\partial^{k-2}+\ldots+c_{0}$, $k>0$. Then there exists an operator $S$ from proposition 2.2 such that $S_{0}=1$, $S_{-1}=0$.
Proof.
The proof will follow the proof of Prop. 7.2 in [42]. $P$ is regular, since $\sigma(P)=\partial^{k}$. Recall that there exist $S$ such that $S_{[0]}=1$, $S_{[i]}=0$ for ($0<i<k$) and each slice can be found from the system
$(\partial^{k}S)_{[p]}=(SP)_{[p]},\quad p\geq 0.$
Note that
$(\partial^{k}S)_{[p]}=\partial^{k}(S_{[p+k]})+\sum_{j=0}^{p+k-1}(\partial^{k}S_{[j]})_{[p]}$
and $(SP)_{[p]}=((S_{[0]}+\ldots+S_{[p]})P)_{[p]}$. Then for $p=0$ we have
$(\partial^{k}S)_{[0]}=\partial^{k}(S_{[k]})+\partial^{k}=P_{[0]},$
and therefore the slice $S_{(k)}$ is uniquely determined; besides, $\mathop{\rm\bf ord}(S_{[k]})=\mathop{\rm\bf ord}(P-\partial^{k})-k\leq-2$.
Now we can use induction on $p$. By induction, we can assume $\mathop{\rm\bf ord}(S_{[k+j]})\leq-2$ for all $j<p$. Note that for $p>0$ we have $\mathop{\rm\bf ord}(P_{[p]})<k-2$ because $P$ is normalized. Then $\mathop{\rm\bf ord}((SP)_{[p]})\leq k-2$. On the other hand, we have $\mathop{\rm\bf ord}((\partial^{k}S_{[j]})_{[p]})\leq k-2$ for all $j<p+k$ (by induction). From the equation above the slice $S_{(k+p)}$ is uniquely determined as
$\partial^{k}(S_{[k+p]})=(SP)_{[p]}-\sum_{j=0}^{p+k-1}(\partial^{k}S_{[j]})_{[p]},$
and therefore $\mathop{\rm\bf ord}(S_{[k+p]})\leq-2$.
Therefore, the homogeneous decomposition of $S$ has no homogeneous terms of order $-1$, and $S_{1}=0$.
Lemma 2.4.
Let $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ be a homogeneous operator commuting with $\partial^{k}$ and $-k<\mathop{\rm\bf ord}(P)=l<0$ (if $\mathop{\rm\bf ord}(P)\leq-k$ then $P=0$ by proposition 2.1, item 2). Then
$P=(\sum_{j=0}^{k-1}c_{j}A_{j})\int^{-l},$
where $c_{j}\in\tilde{K}$ satisfy the following conditions:
$\sum_{j=0}^{k-1}c_{j}=0,\quad\sum_{j=1}^{k-1}c_{j}(\xi^{j}-1)^{q}=0,\quad 1\leq q\leq-l-1,$
or, equivalently,
$\sum_{j=0}^{k-1}c_{j}\xi^{j(q-1)}=0\quad\mbox{for\quad}q=1,\ldots,-l.$
Vice versa, any such operator $P$ commutes with $\partial^{k}$.
Proof.
All assertions of lemma can be deduced from the proof of [42, Prop.7.1]. Another proof is as follows.
By proposition 2.1, item 2 $P$ has the form as it is claimed and we only need to prove the relations between coefficients $c_{j}$. Since the order of $P$ is negative, we have $P_{[q]}=0$ for $q=0,\ldots,-l-1$. On the other hand, using lemma 2.5, we have
$0=[P,\partial^{k}]=(\sum_{j=0}^{k-1}c_{j}A_{j})(1-\sum_{q=1}^{-l}B_{q})\partial^{k+l}-(\sum_{j=0}^{k-1}c_{j}A_{j})\partial^{k+l}=-\sum_{j=0}^{k-1}\sum_{q=1}^{-l}c_{j}\xi^{j(q-1)}B_{q}\partial^{k+l},$
hence $\sum_{j=0}^{k-1}c_{j}\xi^{j(q-1)}=0$ for $q=1,\ldots,-l$.
The same calculation proves the last assertion. ∎
2.2 Basic formulae in $\hat{D}_{1}^{sym}$
In this section we collect useful commutation relations between operators in the ring $\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ which will be used later.
First we define a series of operators $B_{i}$: $B_{1}:=\delta,B_{2}:=x\delta\partial,\ldots,B_{n}:=\frac{1}{(n-1)!}x^{n-1}\delta\partial^{n-1}$, and $B_{j}:=0$ for any $j\leq 0$. Define $\Gamma_{i}=(x\partial)^{i}$ for $i\geq 0$, and for $i<0,\Gamma_{i}=0$. For convenience, we introduce also a new notation: for any integer $n$ we set $D^{n}=\partial^{n}$ if $n\geq 0$ and $D^{n}=\int^{-n}$ otherwise. Obviously, we have $\partial\delta=\delta x=0$ and $\mathop{\rm\bf ord}(B_{j})=0$ for any $j\in\mathbb{N}$.
Lemma 2.5.
[…]
  1. 1.
    $A_{i}\cdot B_{j}=B_{j}\cdot A_{i}=\xi^{i(j-1)}B_{j}$ for any $i,j\in\mathbb{N}$;
  2. 2.
    $\int x^{m}=\frac{m!}{(m+1)!}x^{m+1}+\sum_{i=1}^{\infty}(-1)^{i}\frac{m!}{(m+i+1)!}x^{m+i+1}\partial^{i};$
    In particular, $\int x^{m}\delta=\frac{1}{m+1}x^{m+1}\delta$;
  3. 3.
    $\int^{m}\cdot\partial^{m}=1-\sum_{k=1}^{m}B_{k}$ for any $m\in\mathbb{N}$;
  4. 4.
    $\int^{u}f(x)=f(x)\int^{u}+\sum_{l=1}^{\infty}\binom{-u}{l}f(x)^{(l)}\int^{u+l}$
[…]
  5. 5.
    $B_{i}B_{j}=\delta_{i}^{j}B_{j}$, where $\delta_{i}^{j}$ is the Kronecker delta;
  6. 6.
    $A_{i}\Gamma_{j}=\Gamma_{j}A_{i}$;
  7. 7.
    $D^{i}\Gamma_{j}=\sum_{l=0}^{j}\binom{j}{l}i^{j-l}\Gamma_{l}D^{i},\quad\Gamma_{j}x^{i}=x^{i}(\sum_{l=0}^{j}\binom{j}{l}i^{j-l}\Gamma_{l})$
  8. 8.
    $\Gamma_{i}B_{j}=B_{j}\Gamma_{i}=(j-1)^{i}B_{j}$;
  9. 9.
    $D^{u}B_{j}=B_{j-u}D^{u}$. In particular, $D^{u}B_{j}=0$ if $u>0$ and $j-u\leq 0$ or $u<0$ and $j-u\leq-u$,
where we assume in all these formulae that $10^{0}:=$ 1.
Proof.
1 can be directly calculated:
$A_{i}B_{j}=\frac{1}{(j-1)!}A_{i}x^{j-1}\delta\partial^{j-1}=\frac{1}{(j-1)!}\xi^{i(j-1)}x^{j-1}A_{i}\delta\partial^{j-1}=\frac{1}{(j-1)!}\xi^{i(j-1)}x^{j-1}\delta\partial^{j-1}=\xi^{i(j-1)}B_{j},$
$B_{j}A_{i}=\frac{1}{(j-1)!}x^{j-1}\delta\partial^{j-1}A_{i}=\frac{1}{(j-1)!}\xi^{i(j-1)}x^{j-1}\delta A_{i}\partial^{j-1}=\frac{1}{(j-1)!}\xi^{i(j-1)}x^{j-1}\delta\partial^{j-1}=\xi^{i(j-1)}B_{j}.$
2. The proof is by induction on $m$. For $m=0$ it is true by definition of $\int$. For generic $m$ we have
$\int x^{m}=(\int x^{m-1})x=(\frac{(m-1)!}{m!}x^{m+1}+\sum_{i=1}^{\infty}(-1)^{i}\frac{(m-1)!}{(m+i)!}x^{m+i+1}\partial^{i})+\\ (\sum_{i=1}^{\infty}(-1)^{i}\frac{(m-1)!i}{(m+i)!}x^{m+i}\partial^{i-1})=\frac{m!}{(m+1)!}x^{m+1}+\sum_{i=1}^{\infty}(-1)^{i}\frac{m!}{(m+i+1)!}x^{m+i+1}\partial^{i}.$
3. The proof is by induction on $m$. When $m=1$, it is obvious. Suppose it has been done when $i<m$. Then
$\int^{m}\cdot\partial^{m}=\int\cdot\int^{m-1}\cdot\partial^{m-1}\cdot\partial=\int\cdot(1-\sum_{k=1}^{m-1}B_{k})\cdot\partial$
$=\int\cdot(1-\sum_{k=1}^{m-2}B_{k})\cdot\partial+\int\cdot(B_{m-1})\cdot\partial=\int\cdot\int^{m-2}\cdot\partial^{m-2}\cdot\partial+\int\cdot(B_{m-1})\cdot\partial$
$=1-\sum_{k=1}^{m-1}B_{k}-\int(\frac{1}{(m-2)!}x^{m-2}\delta\partial^{m-2})\partial=1-\sum_{k=1}^{m-1}B_{k}-B_{m}.$
4. Note that the equality $\int x=x\int-\int^{2}$ hold iff $\int x\partial^{q}=(x\int-\int^{2})\partial^{q}$ hold for any $q\geq 0$ (as it follows from definition of the ring $\hat{D}_{1}^{sym}$). Take $q=2$. Then we have
$\int x\partial^{2}=\int\partial^{2}x-2\int\partial=(1-B_{1})\partial x-2(1-B_{1})=x\partial-1+B_{1}$
On the other hand,
$(x\int-\int^{2})\partial^{2}=x(1-B_{1})\partial-(1-B_{1}-B_{2})=x\partial-1+B_{1}.$
So, $\int x\partial^{2}=(x\int-\int^{2})\partial^{2}$ and therefore $\int x=x\int-\int^{2}$. The second formula follows immediately by induction (note that each summand is homogeneous of order $(-u+l)$ and therefore the total sum is well defined for any series $f(x)$).
5. Notice that $\partial^{i}x^{j}$ has a constant term only when $i=j$. By this reason $B_{i}B_{j}=0$ if $i\neq j$, and
$B_{i}B_{i}=\frac{1}{(i-1)!}x^{i-1}\delta\partial^{i-1}\frac{1}{(i-1)!}x^{i-1}\delta\partial^{i-1}=\frac{1}{(i-1)!}x^{i-1}\delta\delta\partial^{i-1}=B_{i}.$
6.
$A_{i}(x\partial)^{j}=(x\partial)A_{i}(x\partial)^{j-1}=\cdots=(x\partial)^{j}A_{i}$
7. We have
$\int(x\partial)=(\int x)\partial=(x\int-\int^{2})\partial=x(1-\delta)-\int(1-\delta)\\ =x-\int-x\delta+\int\delta=x-\int=(x\partial-1)\int.$
Hence
$D^{i}(x\partial)=(x\partial+i)D^{i},\quad(x\partial)x^{i}=x^{i}(x\partial+i).$
and we have
$D^{i}\Gamma_{j}=(x\partial+i)^{j}D^{i}=\sum_{l=0}^{j}\binom{j}{l}i^{j-l}(x\partial)^{l}D^{i},\quad\Gamma_{j}x^{i}=x^{i}(x\partial+i)^{j}=x^{i}(\sum_{l=0}^{j}\binom{j}{l}i^{j-l}(x\partial)^{l}).$
8. First note that
$(x\partial)B_{j}=(x\partial)\frac{1}{(j-1)!}x^{j-1}\delta\partial^{j-1}\\ =(j-1)\frac{1}{(j-1)!}x^{j-1}\delta\partial^{j-1}+\frac{1}{(j-1)!}x^{j}\partial\delta\partial^{j-1}\\ =(j-1)B_{j}+0$
Analogously, $B_{j}(x\partial)=(j-1)B_{j}$. Hence
$\Gamma_{i}B_{j}=(x\partial)^{i}B_{j}=(j-1)(x\partial)^{i-1}B_{j}=(j-1)^{i}B_{j}.$
9. Like before, notice that we have $\partial B_{j}=B_{j-1}\partial$ and $\int B_{j}=B_{j+1}\int.$
2.3 Homogeneous canonical polynomials
[…]
The following definition is motivated by proposition 2.1 and some calculations below.
Definition 2.3.
Let $\tilde{K}=K[\xi]$. An element $H\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ is called homogeneous canonical polynomial (in short of HCP) if $H$ can be written in the form
$H=[\sum_{0\leq i<k}f_{i;r}(x,A_{k;i},\partial)+\sum_{0<j\leq N}g_{j;r}B_{j}]D^{r}$ (2.4)
[…]
  1. 1.
    $f_{i;r}(x,A_{k;i},\partial)$ is a polynomial of $x,A_{k;i},\partial$, $\mathop{\rm\bf ord}(f_{i;r})=0$, of the form
    $f_{i;r}(x,A_{k;i},\partial)=\sum_{0\leq l\leq d_{i}}f_{l,i;r}x^{l}A_{k;i}\partial^{l}$
[…]
  2. 2.
    $g_{j;r}\in\tilde{K}$.
  3. 3.
    $g_{j;r}=0$ for $j\leq-r$ if $r<0$.
In particular, $H$ is homogeneous and $\mathop{\rm\bf ord}(H)=r$.
Using the results of previous section, it can be shown that the form (2.4) of HCP from definition is uniquely defined. Namely, this follows from lemma:
Lemma 2.6.
Let $H$ be a HCP. Then $H=0$ iff $g_{j;r}=0$ for all $j$ and $f_{l,i;r}=0$ for all $i,l$.
In particular, any HCP can be uniquely written in the form (2.4).
Proof.
Obviously, of all coefficients are equal to zero, then $H=0$. Now assume the converse:
$H=[\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}x^{l}A_{k;i}\partial^{l}+\sum_{0<j\leq N}g_{j;r}B_{j}]D^{r}=0$
and some coefficients $g_{j;r},f_{l,i;r}$ are not equal to zero. Then necessarily $H^{\prime}:=HD^{-r}=0$.
Note that, using lemma 2.5, we can rewrite the first sum as
$\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}x^{l}A_{k;i}\partial^{l}=\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}\xi^{-il}A_{k;i}x^{l}\partial^{l}=\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}^{\prime}\xi^{-il}A_{k;i}\Gamma_{l},$
where $f_{d_{i},i;r}^{\prime}\in\tilde{K}$ are some new coefficients, but $f_{d_{i},i;r}^{\prime}=f_{d_{i},i;r}$ for all $i$.
Next, note that
$H^{\prime}=[\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}^{\prime}\xi^{-il}A_{k;i}\Gamma_{l}+\sum_{0<j\leq N^{\prime}}g_{j;r}^{\prime}B_{j}]$
for some new $^{\prime}\in{\mbox{\dbl N}}$ ℕ, $g_{j;r}^{\prime}\in\tilde{K}$, but with the same coefficients $f_{l,i;r}^{\prime}$. Indeed, if $r\geq 0$, then $D^{r}D^{-r}=1$ and therefore all coefficients of $H^{\prime}$ are the same. If $r<0$, then by lemma 2.5 $D^{r}D^{-r}=1-\sum_{i=1}^{r}B_{i}$, and by the same lemma any product
$[\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}^{\prime}\xi^{-il}A_{k;i}\Gamma_{l}+\sum_{0<j\leq N}g_{j;r}B_{j}]B_{i}$
is just a linear combination of some $B_{j}$.
Let $d_{I}$ be a maximal $x$-degree, i.e. $f_{d_{I},I;r}^{\prime}$ is the highest non-zero coefficient. Note that for any $0\leq t<k$ and for any $n\gg 0$ we have by lemma 2.5
$0=[\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}^{\prime}\xi^{-il}A_{k;i}\Gamma_{l}+\sum_{0<j\leq N^{\prime}}g_{j;r}^{\prime}B_{j}]B_{kn+t+1}=\\ \sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}^{\prime}\xi^{-il}\xi^{it}(kn+t)^{l}B_{kn+t+1}$
where
$\sum_{0\leq i<k}f_{d_{I},i;r}^{\prime}\xi^{i(t-d_{I})}=0,\quad 0\leq t<k$
(where we assume $f_{d_{I},i;r}^{\prime}=0$ if $d_{I}>d_{i}$). But by the well known property of the Vandermonde matrix this system has the unique solution $f_{d_{I},0;r}^{\prime}=\ldots=f_{d_{I},k-1;r}^{\prime}=0$, a contradiction. So,
$H=[\sum_{0<j\leq N}g_{j;r}B_{j}]D^{r}.$
Assume $g_{j_{0};r}$ is the first non-zero coefficient. Then again by lemma 2.5 $B_{j_{0}}H=g_{j_{0};r}B_{j_{0}}D^{r}\neq 0$, a contradiction. ∎
Definition 2.4.
Let $H$ be a HCP. We define
$Sdeg_{A}(H)=\max\{d_{i}|\quad 0\leq i<k\}\quad\mbox{or $-\infty$, if all $f_{l,i;r}=0$ }$
and
$Sdeg_{B}(H)=\max\{j|\quad g_{j;r}\neq 0\}\quad\mbox{or $-\infty$, if all $g_{j;r}=0$}$
We define a homogeneous canonical polynomial combination (in short HCPC) as a finite sum of HCP (of different orders). We extend the functions $Sdeg_{A}$, $Sdeg_{B}$ in an obvious way to all HCPCs.
We’ll say that a HCPC $H$ doesn’t contain $A_{i}$ if $f_{l,i;r}=0$ for all $i>0$ and all $r$. We’ll say that a HCPC $H$ doesn’t contain $B_{j}$ if $Sdeg_{B}(H)=-\infty$.
Example 2.2.
Suppose $H$ is a HCP and $H\in D_{1}\hat{\otimes}_{K}\tilde{K}$ is a differential operator. Then it is easy to see that it can be uniquely written in the form (2.4), which does not contain $A_{i}$ and $B_{j}$.
In practice it is more convenient to work with another form of HCPs, which we have already used in the proof of lemma 2.6:
Definition 2.5.
Suppose $H$ is a HCP. Then $H$ can be (uniquely) written in another form:
$H=(\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f^{\prime}_{l,i;r}\Gamma_{l}A_{i}+\sum_{0<j\leq N}g_{j;r}B_{j})D^{r}$
which we’ll call the G-form of $H$.
It’s easy to see that two forms of $H$ are one to one correspondence to each other, and therefore the G-form is also uniquely defined. Besides, the definitions of $Sdeg_{A}$, $Sdeg_{B}$ does not depend on the form, i.e. $Sdeg_{A}(H)$ is again the maximal $d_{i}$.
The following lemma is the first obvious property of HCPCs.
Lemma 2.7.
Suppose $H$ and $M$ are two HCPCs, $k_{1},k_{2}\in\tilde{K}$ are two arbitrary constant. Then $T=k_{1}H+k_{2}M$ is also a HCPC, with
$Sdeg_{A}(T)\leq\max\{Sdeg_{A}(H),Sdeg_{A}(M)\}$
$Sdeg_{B}(T)\leq\max\{Sdeg_{B}(H),Sdeg_{B}(M)\}$
Before presenting next result, we need a lemma from standard ODE book:
Lemma 2.8.
([30, Ch.2, item 10 Th.8]) Suppose $L,f\in\tilde{K}[t]$ are non-zero polynomials with $\deg f=r$, $\lambda\in\tilde{K}$, and $q$ is the multiplicity of $t-\lambda$ in the polynomial $L(t)$ (if $L(\lambda)\neq 0$, then $q=0$). Then the ODE
$L(\partial)z=\sum_{i}a_{i}z^{(i)}=f(t)e^{\lambda t}$
has a solution in the form of
$z_{0}=t^{q}g(t)e^{\lambda t},$
where $g(t)$ is a polynomial of degree $r$ (the same as $f$).
Remark 2.3.
[…]
Now we are ready to prove the following claim.
Lemma 2.9.
Suppose $M$ is a HCPC. Suppose $H\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ is a HCPC satisfying the condition
$[\partial^{k},H]=M$
where $k$ is the original $k$ we have fixed. Then we have
  1. 1.
    If $Sdeg_{A}(M)\neq-\infty$ then $Sdeg_{A}(H)=Sdeg_{A}(M)+1$ and $Sdeg_{B}(H)=Sdeg_{B}(M)$.
  2. 2.
    If $Sdeg_{A}(M)=-\infty$ then $Sdeg_{A}(H)$ is either 00 or $\infty$- ∞ and $Sdeg_{B}(H)=Sdeg_{B}(M)$.
Proof.
  1. 1.
    We use the same idea as in the proof of proposition 2.1. Considering $\partial$ as a constant the original equation becomes an ODE in the ring $\tilde{R}$:
    $H^{(k)}+\dots+kH^{\prime}\tilde{\partial}^{k-1}=M$ (2.5)
    The eigen-polynomial of this ODE is $L(t):=(t-\tilde{\partial})^{k}-\tilde{\partial}^{k}$. Any solution $H$ of this equation looks like $H=H_{0}+\sum_{i=0}^{k-1}c_{i}\tilde{A}_{k,i}$, where $c_{i}\in\tilde{K}((\tilde{\partial}))$ and $H_{0}$ is a special solution. Since $M$ is a HCP, it can be represented as a linear combination of quasi-polynomials in the ring $\tilde{R}$, which are homogeneous with respect to the order $\mathop{\rm\bf ord}$. Therefore, there is a partial solution equal to a sum of partial solutions of similar equations with monomial right hand side $M$.
  2. 2.
    We have two possible case of such monomials:
    1. (a)
      $M=f_{i}(x)\tilde{A}_{i}$. Notice that $\tilde{A}_{i}=e^{(\xi^{i}-1)x\tilde{\partial}}$ and $(\xi^{i}-1)\tilde{\partial}$ is the root of $L(t)$ with multiplicity 1. Then according to Lemma 2.8, such ODE has a special solution $H_{0}=xg_{i}(x)\tilde{A}_{i}$ (with $deg_{x}(f_{i})=deg_{x}(g_{i})$).
      Now note that, since $M$ is homogeneous, a special solution can be chosen to be also homogeneous (just take the homogeneous component of any special solution), because the left hand side of our ODE is homogeneous for any homogeneous $H$. Then such a solution will be also an HCP and
      $Sdeg_{A}(H_{0})=Sdeg_{A}(M)+1,\quad Sdeg_{B}(H_{0})=Sdeg_{B}(M)=-\infty.$
    2. (b)
      $M=b_{j}\tilde{B}_{j}$, where $b_{j}\in\tilde{K}$. Notice that $\tilde{B}_{j}=cx^{j-1}\tilde{\partial}^{j-1}\tilde{\delta}=cx^{j-1}\tilde{\partial}^{j-1}e^{-x\tilde{\partial}}$, and $-\tilde{\partial}$ is not a root of $L(t)$. Then according to Lemma 2.8, we know the special solution of this ODE will be in the form
      $H_{0}=h_{j}(x)\tilde{\delta}$
      with $deg_{x}(h_{j})=j-1=deg_{x}(B_{j})$. Again a special solution can be chosen to be homogeneous, i.e. it is a HCP. Thus we can write this $H_{0}$ in the form
      $H_{0}=\sum_{l=1}^{j}h_{l-1,j}\tilde{B}_{l}\tilde{\partial}^{\mathop{\rm\bf ord}(M)-k}$
      with $h_{l-1,j}\in\tilde{K}$, $h_{j-1,j}\neq 0$. Hence $Sdeg_{A}(H_{0})=-\infty$, $Sdeg_{B}(H_{0})=Sdeg_{B}(M)$.
  3. 3.
    Now we have: $H=H_{0}+\sum_{i=0}^{k-1}c_{i}\tilde{A}_{k,i}$, where $c_{i}\in\tilde{K}((\tilde{\partial}))$ and $H_{0}$ is a HCP. Moreover, if $Sdeg_{A}(M)\neq-\infty$, then $Sdeg_{A}(H_{0})=Sdeg_{A}(M)+1$ and $Sdeg_{B}(H_{0})=Sdeg_{B}(M)$. And if $Sdeg_{A}(M)=-\infty$, then $Sdeg_{A}(H_{0})=-\infty$ and $Sdeg_{B}(H_{0})=Sdeg_{B}(M)$. Since our original solution $H\in\hat{D}_{1}^{sym}$, we have $\mathop{\rm\bf ord}(H)<\infty$ and therefore $\mathop{\rm\bf ord}(c_{i})<\infty$ for all $i$, i.e. $c_{i}\in\tilde{K}[\tilde{\partial}^{-1},\tilde{\partial}]$ and $H$ is a HCPC.
    Writing this solution in a canonical form and replacing $\tilde{\partial}$ by $D$, we get, as in the proof of propostion 2.1, that $H$ is a HCPC. Since $Sdeg_{A}(\sum_{i=0}^{k-1}c_{i}{A}_{k,i})=0$, we get our assertion. Thus we complete the proof.
Remark 2.4.
The condition about existence of $H$ from lemma is essential: for example $[\partial^{4},H]=B_{10}\partial^{4}$ doesn’t have any HCPC solution. But we will see if $M$ doesn’t contain $B_{j}$ such $H$ must exists, see lemma 2.11.
The following lemma describes basic properties of functions $Sdeg_{A}$ and $Sdeg_{B}$. Besides, it contains also useful formulae for monomial multiplication.
Lemma 2.10.
Suppose $H,M$ are two HCPCs. Then $T:=HM$ is also a HCPC, what’s more we have
  1. 1.
    $Sdeg_{A}(T)\leq Sdeg_{A}(H)+Sdeg_{A}(M)$ (here we assume $-\infty+n=-\infty$).
  2. 2.
    1. (a)
      If $\mathop{\rm\bf ord}(H)\geq 0$, then $Sdeg_{B}(T)\leq\max\{Sdeg_{B}(H),Sdeg_{B}(M)\}$
    2. (b)
      If $\mathop{\rm\bf ord}(H)<0$, then $Sdeg_{B}(T)\leq\max\{Sdeg_{B}(H),Sdeg_{B}(M)-\mathop{\rm\bf ord}(H),-\mathop{\rm\bf ord}(H)\}$
Proof.
First let’s prove $T$ is a HCPC. Obviously, it is enough to prove this for the case when $H,M$ are monomials written in the G-form. As a byproduct we’ll get convenient multiplication formulae of HCPCs.
Consider the following 4 cases (below we assume in all formulae that $10^{0}:=$ 1):
  1. 1.
    $H=b_{i}B_{i}D^{u},M=c_{j}B_{j}D^{v}$, where $i\geq 1-u$ if $u<0$ and $j\geq 1-v$ if $v<0$, $b_{i},c_{j}\in\tilde{K}$. We have by lemma 2.5
    $HM=b_{i}B_{i}D^{u}c_{j}B_{j}D^{v}=b_{i}c_{j}B_{i}B_{j-u}D^{u}D^{v}=b_{i}c_{j}\delta_{i}^{j-u}B_{i}D^{u+v}$ (2.6)
    (here $\delta_{i}^{j-u}$ is the Kronecker delta), because if $i=j-u$, then $i-1+u+v=j-1+v\geq 0$ and $B_{i}D^{u+v}\neq 0$.
  2. 2.
    $H=b_{i}B_{i}D^{u}$, where $i\geq 1-u$ if $u<0$, $M=a_{l,m}\Gamma_{m}A_{j}D^{v}$, $b_{i},a_{l,m}\in\tilde{K}$. By Lemma 2.5 item 12 we know $B_{i}D^{u}=D^{u}B_{i+u}$, hence $i\geq 1-u$, otherwise $H=0$. Then we have
    $HM=\begin{cases}0&i-1+u+v<0\\ a_{l,m}b_{i}\xi^{j(u+i-1)}(i-1+u)^{m}B_{i}D^{u+v}&\text{Otherwise}\end{cases}$ (2.7)
    This is because
    $B_{i}D^{u}\Gamma_{m}A_{j}D^{v}=B_{i}(x\partial+u)^{m}D^{u}A_{j}D^{v}=\xi^{uj}B_{i}(x\partial+u)^{m}A_{j}D^{u}D^{v}\\ =\xi^{uj}B_{i}A_{j}(x\partial+u)^{m}D^{u}D^{v}=\xi^{(u+i-1)j}B_{i}(x\partial+u)^{m}D^{u}D^{v}\\ =\xi^{(u+i-1)j}(i-1+u)^{m}B_{i}D^{u}D^{v}$
    where the first equality is by Lemma 2.5 item 7; the second equality is by Prop 2.1; the third equality is by Lemma 2.5 item 6; the forth equality is by Lemma 2.5 item 1; the fifth equality is by Lemma 2.5 item 8, since $B_{i}(x\partial)=(i-1)B_{i}$, so $B_{i}(x\partial+u)=(u+i-1)B_{i}$. Notice that when $u<0$, and $v>0$, we have
    $D^{u}D^{v}=(1-\sum_{s=-u-v+1}^{-u}B_{s})D^{u+v}$
    And by item 1 we know $B_{i}B_{j}=\delta^{i}_{j}B_{i}$, hence we know if $-u-v+1\leq i\leq-u$, then
    $B_{i}D^{u}D^{v}=(B_{i}-B_{i})D^{u+v}=0$
    But we already assume $i\geq 1-u$, hence $B_{i}D^{u}D^{v}=B_{i}D^{u+v}$, again by Lemma 2.5 item 9, we know it is 0 when $i-1+u+v<0$.
  3. 3.
    $H=a_{l,m}\Gamma_{m}A_{j}D^{v},M=b_{i}B_{i}D^{u}$, where $i\geq 1-u$ if $u<0$, $b_{i},a_{l,m}\in\tilde{K}$. We have
    $HM=a_{l,m}b_{i}\xi^{j(i-v-1)}(i-v-1)^{m}B_{i-v}D^{v}D^{u}=\begin{cases}0&i-v<1\\ \lambda B_{i-v}D^{u+v}&\text{Otherwise.}\end{cases}$ (2.8)
    where $\lambda=a_{l,m}b_{i}\xi^{j(i-v-1)}(i-v-1)^{m}$. This is because when $i-v<1$ we have $D^{v}B_{i}=B_{i-v}D^{v}=0$ (c.f. Lemma 2.5 item 9 and by assumptions for $r\leq 0$, $B_{r}=0$), so that
    $\Gamma_{m}A_{j}D^{v}B_{i}D^{u}=\Gamma_{m}A_{j}(D^{v}B_{i})D^{u}=0$
    and when $i-v\geq 1$, we have
    $\Gamma_{m}A_{j}D^{v}B_{i}D^{u}=\Gamma_{m}A_{j}B_{i-v}D^{v}D^{u}=\xi^{j(i-v-1)}\Gamma_{m}B_{i-v}D^{v}D^{u}\\ =\xi^{j(i-v-1)}(i-v-1)^{m}B_{i-v}D^{v}D^{u}=\xi^{j(i-v-1)}(i-v-1)^{m}B_{i-v}D^{v+u}$
    The last equality holds obviously when $v\geq 0$ or $u\leq 0$. When $v<0,u>0$, notice that $i>0$, so that $i-v>-v$, hence for any $-u-v+1\leq s\leq-v$, we have $B_{i-v}B_{s}=\delta_{s}^{i-v}=0$ by Lemma 2.5 item 5. Hence
    $B_{i-v}D^{v}D^{u}=B_{i-v}(1-\sum_{s=-u-v+1}^{-v}B_{s})D^{u+v}=B_{i-v}D^{u+v}$
  4. 4.
    $H=a_{i,m}\Gamma_{m}A_{i}D^{u},M=a_{j,n}\Gamma_{n}A_{j}D^{v}$, $a_{i,m},a_{j,n}\in\tilde{K}$. We have
    $HM=a_{i,m}a_{j,n}\xi^{uj}\sum_{l=0}^{n}\binom{n}{l}u^{n-l}\Gamma_{l+m}A_{i+j}D^{u}D^{v}\\ =\begin{cases}a_{i,m}a_{j,n}\xi^{uj}\sum_{l=0}^{n}\binom{n}{l}u^{n-l}\Gamma_{l+m}A_{i+j}D^{u+v}+\sum_{s=-u-v+1}^{-u}\lambda_{s}B_{s}D^{u+v}&u<0,v>0\\ a_{i,m}a_{j,n}\xi^{uj}\sum_{l=0}^{n}\binom{n}{l}u^{n-l}\Gamma_{l+m}A_{i+j}D^{u+v}&\text{Otherwise.}\end{cases}$ (2.9)
    where $\lambda_{s}=-a_{i,m}a_{j,n}\xi^{uj+(i+j)(s-1)}(u+s-1)^{n}(s-1)^{m}$, when
    $\max\{1,-u-v+1\}\leq s\leq-u.$
    This is because
    $\Gamma A_{i}D^{u}\Gamma_{n}A_{j}D^{v}=\Gamma_{m}A_{i}(x\partial+u)^{u}D^{u}A_{j}D^{v}=\xi^{uj}A_{i}\Gamma_{m}(x\partial+u)^{n}A_{j}D^{u}D^{v}\\ =\xi^{uj}A_{i+j}\Gamma_{m}(x\partial+u)^{n}D^{u}D^{v}$
    and for $-u-v+1\leq s\leq-u$ when $u<0,v>0$, we have
    $A_{i+j}\Gamma_{m}(x\partial+u)^{n}B_{s}=(s-1+u)^{n}A_{i+j}\Gamma_{m}Bs\\ =(s-1)^{m}(s-1+u)^{n}A_{i+j}B_{s}=\xi^{(i+j)(s-1)}(s-1)^{m}(s-1+u)^{n}B_{s}$
Hence we know the product is a HCPC. Now we can estimate the $Sdeg$ of $T$:
  1. 1.
    Consider $Sdeg_{A}(T)$. Observe that in cases 1-4 our claim is true. So, it is true for a product of any two HCPCs. Note also that in the case 4, for $H,M$ being monomials, we have the equality $Sdeg_{A}(T)=Sdeg_{A}(H)+Sdeg_{A}(M)$, however, in general, if we take the product of sums of monomials, a strict inequality can appear.
  2. 2.
    Consider $Sdeg_{B}(T)$. Then explicit formulae of cases 1-4 show our assertion in general case.
Corollary 2.1.
For a fixed $k$ all $HCPCs$ form a subring in the ring $\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$.
Remark 2.5.
Denote by
$Hcpc(k):=\{\text{all HCPCs assoicated to }k\}$
the subring from corollary. It’s easy to observe that if $a|b$, $b=ra$ and $\xi,\eta$ are respectively the $a,b$-primitive roots, then $\eta^{r}=\xi$, i.e. $A_{b,ir}=A_{a,i}$. This means
$Hcpc(a)\subseteq Hcpc(b)$
Thus for $p,q$, if we assume $r=gcd(p,q)$, then we have
$Hcpc(p)\bigcap Hcpc(q)\supseteq Hcpc(r).$
Lemma 2.11.
Suppose $M$ is a HCPC which doesn’t contain $B_{j}$ (i.e. $Sdeg_{B}(M)=-\infty$). Then there exists a HCPC $H\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$, which satisfies the condition
$[\partial^{k},H]=M.$
Proof.
Take the HCP decomposition of $M$ $M=\sum_{i=l_{1}}^{l_{2}}M_{i}$ , where $\mathop{\rm\bf ord}(M_{i})=i$.
It’s enough to find HCPs $H_{i}$ such that
$[\partial^{k},H_{i}]=M_{i+k}.$
From now on, we always assume both $H,M$ are HCPs, with $\mathop{\rm\bf ord}(H)=m-k$, $\mathop{\rm\bf ord}(M)=m$.
Written $H,M$ in G-form, we solve the equation into two steps:
  1. 1.
    Assume
    $H=(H_{0;m-k}+H_{1;m-k}A_{1}+\dots+H_{k-1;m-k}A_{k-1})D^{m-k}$
    with
    $M=(M_{0;m}+M_{1;m}A_{1}+\dots+M_{k-1;m}A_{k-1})D^{m}.$
    Since $[\partial^{k},A_{i}]=0$, we know by lemma 2.10 that
    $[\partial^{k},H_{i;m-k}A_{i}D^{m-k}]=M_{i;m}A_{i}D^{m}\qquad[\mod B_{j}],$ (2.10)
    where $[\mod B_{j}]$ denote denote terms containing $B_{j}$s. Now assume
    $H_{i;m-k}=\sum_{l=0}^{t_{i}}h_{l,i;m-k}\Gamma_{l}.$
    Since $[\partial^{k},h_{0,i;m-k}A_{i}D^{m-k}]$ contains only terms with $B_{j}$, we can ignore terms with $h_{0,i;m-k}$ by looking for a solution. Assume
    $M_{i;m}=\sum_{l=0}^{t_{i}-1}m_{l,i;m}\Gamma_{l}.$
    Using Lemma 2.10, we can directly calculate $[\partial^{k},H_{i;m-k}A_{i}D^{m-k}]$. Then we get the linear system for the unknown coefficients ${h_{1,i;m-k},\dots,h_{t_{i},i;m-k}}$:
    $\begin{pmatrix}k&*&\dots&*\\ 0&2k&\dots&*\\ \dots&\dots&\dots&\dots\\ 0&0&\dots&t_{i}k\end{pmatrix}\begin{pmatrix}h_{1,i;m-k}\\ h_{2,i;m-k}\\ \dots\\ h_{t_{i},i;m-k}\end{pmatrix}=\begin{pmatrix}m_{0,i;m}\\ m_{2,i;m}\\ \dots\\ m_{t_{i}-1,i;m}\end{pmatrix}$ (2.11)
    Since this is a lower-triangular matrix, with all diagonal elements non-zero, this equation system always has a solution. Solving it, we get almost all coefficients in $H$ except for $h_{0,0;m-k},\dots,h_{0,k-1;m-k}$.
  2. 2.
    Suppose $\tilde{H}$ is the result we get in step 1. From the discussion in step 1, we know
    $[\partial^{k},\tilde{H}]=M\quad[\mod B_{j}]$
    Now assume with
    $\bar{H}=(h_{0,0;m-k}+\cdots+h_{0,k-1;m-k}A_{k-1})D^{m-k}$
    with $H=\tilde{H}+\bar{H}$. The equation $[\partial^{k},H]=M$ becomes
    $[\partial^{k},\bar{H}]=M-[\partial^{k},\tilde{H}]$ (2.12)
    Notice that terms on the right hand side are already known, and they contains only terms with $B_{j}$ (i.e. $Sdeg_{A}=-\infty$). There are three possible cases:
    1. (a)
      $m\geq k$, in such case no $B_{j}$ can appear on the right hand side of (2.10). So we can simply put $h_{0,0;m-k},\dots,h_{0,k-1;m-k}$ all 0. We will get $[\partial^{k},H]=M$.
    2. (b)
      $m\leq 0$, in such case, notice that
      $D^{m-k}\partial^{k}=\int^{-m+k}\partial^{k}=(1-B_{-m+1}-\cdots-B_{-m+k})$
      So there might be $k$ summands of $B_{j}$ in the left hand side of equation (2.12). On the other hand, we know there might be at most $k$ summands with $B_{j}$ ($B_{-m+1},\dots,B_{-m+k}$) on the right hand side of the equation (according to the uniqueness of HCPC). By the same reason we know the coefficients at $B_{-m+1},\dots,B_{-m+k}$ on both sides must be equal to each other respectively. Hence we get $k$ linear equations for $h_{0,0;m-k},\dots,h_{0,k-1;m-k}$.
      Calculating the coefficients at $B_{-m+j}$ on both sides, we have
      $\sum_{i=0}^{k-1}\xi^{i(j-m-1)}h_{0,i;m-k}=b_{j-m}$
      where $b_{1-m},\dots,b_{k-m}$ are already known on the right hand side. Collecting this linear system, we find we have $k$ variables and $k$ equations, the coefficient matrix is a Vandermonde matrix (hence always of full rank). Thus we can always solve the equation and find the coefficients $h_{0,0;m-k},\dots,h_{0,k-1;m-k}$. After that we get $[\partial^{k},\tilde{H}+\bar{H}]=M$.
    3. (c)
      $0<m<k$, in this case the same arguments as in case b) work, and we omit the details. We only need to notice that here we have $k$ variables but $(k-m)$ equations, and the coefficient matrix becomes the full rank sub-matrix of the Vandermonde matrix. So the solution exists but might be not unique.
2.4 Some necessary conditions on the Schur operator
Let $Q\in D_{1}$ be a normalized differential operator, i.e $Q=\partial^{q}+(\dots)\partial^{q-2}+\dots$. According to Propositions 2.2, 2.3 there exists $S\in\hat{D}_{1}^{sym}$, such that $S^{-1}QS=\partial^{q}$, where $S=S_{0}+S_{-1}+\dots$, with $S_{0}=1,S_{-1}=0$.
In this section we establish several necessary conditions on $S$. Namely, we’ll show that all homogeneous components of $S$, $S^{-1}$ are HCPs with $Sdeg_{B}=-\infty$, and the function $Sdeg_{A}$ has a linear upper and lower bound.
From now on we fix $k=q=\mathop{\rm\bf ord}(Q)$ (recall that in our case $\mathop{\rm\bf ord}(Q)=deg(Q)$). Let $\xi$ be a $q$-th primitive root of $1$, $\tilde{K}=K[\xi]$.
Definition 2.6.
[…]
Example 2.3.
It’s easy to see that an operator $P\in D_{1}$ with an invertible highest coefficient (cf.remark 2.1) written in G-form is totally free of $B_{j}$, i.e. all its homogeneous components have this property.
Lemma 2.12.
The subset of totally free of $B_{j}$ elements from $Hcpc(k)$ for a fixed $k$ form a subring in $\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$.
Proof.
Clearly, this subset form a linear subspace (cf. lemma 2.7). To prove the claim it suffices to prove that the product of two totally free of $B_{j}$ HCPs $H_{1},H_{2}$ is totally free of $B_{j}$.
[…]
By lemma 2.5 we have
$H_{1}(c\Gamma_{n}A_{q}D^{v})=c(x\partial+u)^{n}\xi^{uq}A_{q}(H_{1}D^{v}),$
and $H_{1}D^{v}$ does not contain $B_{j}$, hence $H_{1}(c\Gamma_{n}A_{q}D^{v})$ does not contain $B_{j}$ and we are done. ∎
Proposition 2.4.
Suppose
$0\neq H=\sum_{i=0}^{k}\sum_{m=0}^{d}a_{m,i}A_{i}\Gamma_{m}\int^{u}.$
is a HCP from $Hcpc(k)$, with $\mathop{\rm\bf ord}(H)=-u<0$. Then $H$ is totally free of $B_{j}$ iff the linear system of equations on $a_{m,i}$ holds:
$\sum_{i=0}^{k-1}\sum_{m=0}^{d}\xi^{i(j-1)}(j-1)^{m}a_{m,i}=0$ (2.13)
for any $1\leq j\leq-u$.
Moreover, if $H$ is totally free of $B_{j}$, then $\frac{u}{k}-1<Sdeg_{A}(H).$
Proof.
Suppose $H$ is totally free of $B_{j}$. Since
$H\partial^{u}=\sum_{i=0}^{k-1}\sum_{m=0}^{d}a_{m,i}A_{i}\Gamma_{m}(1-B_{1}-\cdots-B_{u})\\ =\sum_{i=0}^{k-1}\sum_{m=0}^{d}a_{m,i}A_{i}\Gamma_{m}-\sum_{j=1}^{u}\sum_{i=0}^{k-1}\sum_{m=0}^{d}\xi^{i(j-1)}(j-1)^{m}a_{m,i}B_{j}$
is free of $B_{j}$, we get a linear equation system (2.13). Vice versa, this system implies the total freeness of $B_{j}$: indeed, for any $l\geq 0$ $H\partial^{l}=H\partial^{u}D^{l-u}$, hence it is free of $B_{j}$ (for $l<0$ no terms with $B_{j}$ can appear).
The second statement requires more technical arguments.
Lemma 2.13.
Suppose $T$ is the coefficient matrix of system (2.13). If $u=(d+1)k$, $T$ is a square matrix and we have $det(T)\neq 0$.
Proof.
Note that $T$ is in the shape of
$\begin{pmatrix}1&1&1&\cdots&1\\ 1&\xi&\xi^{2}&\cdots&\xi^{u-1}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ 1&\xi^{k-1}&\xi^{2(k-1)}&\cdots&\xi^{(k-1)(u-1)}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ 0&1&2^{d}&\cdots&(u-1)^{d}\\ 0&\xi&\xi^{2}2^{d}&\cdots&\xi^{u-1}(u-1)^{d}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ 0&\xi^{k-1}&\xi^{2(k-1)}2^{d}&\cdots&\xi^{(k-1)(u-1)}(u-1)^{d}\end{pmatrix}$
suppose there exist $\overrightarrow{\alpha}=(\alpha_{1},\dots,\alpha_{(d+1)k}),\alpha_{l}\in\tilde{K},1\leq l\leq k(d+1)$, such that
$\overrightarrow{\alpha}T=0$
Then we have
$(\alpha_{(d+1)k}\xi^{(k-1)(u-1)}+\cdots+\alpha_{(d+1)k-k+1})(u-1)^{d}+(\cdots)(u-1)^{d-1}\\ +\cdots+(\alpha_{k}\xi^{(k-1)(u-1)}+\cdots+\alpha_{1})=0$
and
$(\alpha_{(d+1)k}\xi^{(k-1)(u-2)}+\cdots+\alpha_{(d+1)k-k+1})(u-2)^{d}+(\cdots)(u-2)^{d-1}\\ +\cdots+(\alpha_{k}\xi^{(k-1)(u-2)}+\cdots+\alpha_{1})=0$
$(\alpha_{(d+1)k}\xi^{k-1}+\cdots+\alpha_{(d+1)k-k+1})+(\cdots)\\ +\cdots+(\alpha_{k}\xi^{k-1}+\cdots+\alpha_{1})=0$
Now denote
$\begin{cases}\eta_{0}=\alpha_{(d+1)k}\xi^{(k-1)(u-1)}+\cdots+\alpha_{(d+1)k-k+1}\\ \eta_{1}=\alpha_{(d+1)k-k}\xi^{(k-1)(u-1)}+\cdots+\alpha_{(d+1)k-2k+1}\\ \cdots\\ \eta_{d}=\alpha_{k}\xi^{(k-1)(u-1)}+\cdots+\alpha_{1}\end{cases}$
Notice that $\xi^{k}=1$, we have
$\alpha_{(d+1)k}\xi^{(k-1)(u-k-1)}=\alpha_{(d+1)k}\xi^{(k-1)(u-1)}$
Hence we know
$\begin{cases}(u-1)^{d}\eta_{0}+(u-1)^{d-1}+\cdots+\eta_{d}=0\\ (u-k-1)^{d}\eta_{0}+(u-k-1)^{d-1}+\cdots+\eta_{d}=0\\ \cdots\\ (u-dk-1)^{d}\eta_{0}+(u-dk-1)^{d-2}\eta_{1}+\cdots+\eta_{d}=0\end{cases}$
Considering $\eta_{0},\eta_{1},\dots,\eta_{d}$ as variables, we get a linear equation systems equipped with Vandermonde’s coefficient matrix. Hence we know $\eta_{0}=\eta_{1}=\cdots=\eta_{d}=0$. In the same way, if we denote
$\begin{cases}\tilde{\eta}_{0}=\alpha_{(d+1)k}\xi^{(k-1)(u-2)}+\cdots+\alpha_{(d+1)k-k+1}\\ \tilde{\eta}_{1}=\alpha_{(d+1)k-k}\xi^{(k-1)(u-2)}+\cdots+\alpha_{(d+1)k-2k+1}\\ \cdots\\ \tilde{\eta}_{d}=\alpha_{k}\xi^{(k-1)(u-2)}+\cdots+\alpha_{1}\end{cases}$
we will also get $\tilde{\eta}_{0}=\tilde{\eta}_{1}=\cdots=\tilde{\eta}_{d}=0$. We play with this game for $k$ times, there will be $k$ numbers of equation for $(\alpha_{1+mk},\alpha_{2+mk},\dots,\alpha_{k+mk})$ for any $0\leq m\leq d$, in the form of
$(\alpha_{1+mk},\alpha_{2+mk},\dots,\alpha_{k+mk})\begin{pmatrix}1&1&\cdots&1\\ 1&\xi^{1}&\cdots&\xi^{k-1}\\ \cdots&\cdots&\cdots&\cdots\\ 1&\xi^{k-1}&\cdots&\xi^{(k-1)(k-1)}\end{pmatrix}=0$
The coefficient matrix is again in Vandermonde’s shape, so there are only 00 solutions, this means $\overrightarrow{\alpha}=0$. Hence $det(T)\neq 0$. ∎
Summing up, $T$ induces a linear map:
$T:\tilde{K}^{k(d+1)}\rightarrow\tilde{K}^{u}.$
If $d=Sdeg_{A}(H)\leq\frac{u}{k}-1$, i.e. $(d+1)k\leq u$, then $T$ is injective (as its matrix consists of linearly independent raws by lemma 2.13), this means the linear equation system has only zero solution, this contradicts to $H\neq 0$. ∎
Corollary 2.2.
Let $C(\partial^{k})\subset\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ be the centraliser of $\partial^{k}$, and $H\in C(\partial^{k})$ be a HCP. Then $H$ is totally free of $B_{j}$.
Proof.
The proof follows immediately from lemma 2.4 and proposition 2.4. ∎
Proposition 2.5.
Let $Q\in D_{1}$ be a normalized operator, assume $S\in\hat{D}_{1}^{sym}$ is a Schur operator for $Q$, i.e. $S^{-1}QS=\partial^{q}$, such that $S_{0}=1,S_{-1}=0$. Then we have
  1. 1.
    $S_{-t}$ is a HCP from $Hcpc(q)$ for any $t\geq 0$ (i.e. it can be written as a HCP in $\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$.
  2. 2.
    If $S_{-t}\neq 0$ then $\frac{t}{q}-1<Sdeg_{A}(S_{-t})<t$ for any $t>0$.
  3. 3.
    $S_{-t}$ is totally free of $B_{j}$ for any $t\geq 0$.
Proof.
The proof is by induction on $t$. All claims are obvious for $S_{-1}$. Suppose they are true from $S_{-1}$ to $S_{-i+1}$. Consider the homogeneous decomposition of $Q$:
$Q=Q_{q}+Q_{q-2}+\ldots=\partial^{q}+\eta_{0;q-2}\partial^{q-2}+(\eta_{0;q-3}+\eta_{1;q-2}\Gamma_{1})\partial^{q-3}+\cdots.$
Note that $Q_{j}$ is a HCP for any $j$, $Sdeg_{B}(Q_{j})=-\infty$ for any $j$ and $Sdeg_{A}(Q_{q-i})\leq i-2$ for any $i>1$.
Since $QS=S\partial^{q}$, we have the equality of their $(q-i)$ homogeneous components: $(QS)_{q-i}=(S\partial^{q})_{q-i}=S_{-i}\partial^{q}$, i.e.
$\partial^{q}\cdot S_{-i}+Q_{q-2}\cdot S_{-i+2}+\dots+Q_{q-i}-S_{-i}\partial^{q}=0.$ (2.14)
Put $M_{i}:=-(Q_{q-2}\cdot S_{-i+2}+\dots+Q_{q-i})$, so that this equation becomes
$[\partial^{q},S_{-i}]=M_{i}.$
Then by lemma 2.10 we get that $M_{i}$ is a HCP and $Sdeg_{A}(M_{i})\leq i-2$ (the only possible term of degree $i-2$ is $Q_{q-i}$:
$Q_{q-i}=(\eta_{0;q-i}+\eta_{1;q-i+1}\Gamma_{1}+\dots+\eta_{i-2;q-2}\Gamma_{i-2})\partial^{q-i},$
whose degree depends on vanishing the coefficient $\eta_{i-2;q-2}$). From formulae (2.9) we get also $Sdeg_{B}(M_{i})=-\infty$. Now applying Lemma 2.9, we get $Sdeg_{A}(S_{-i})\leq i-1$ and $Sdeg_{B}(S_{-i})=-\infty$ (note that since $S_{-i}$ is homogeneous, $S_{-i}$ is a HCP).
In view of formulae (2.6)-(2.9) we have $Sdeg_{B}(S_{-t}D^{\lambda})=-\infty$ if $\lambda\leq 0$. Assume $\lambda>0$.
[…]
$\partial^{pq}\cdot S_{-t}+(Q^{p})_{pq-2}\cdot S_{-t+2}+\dots+(Q^{p})_{pq-t}=S_{-t}\partial^{pq}.$
[…]
The second inequality of item 2 follows from proposition 2.4. ∎
Corollary 2.3.
In the notation of proposition 2.5 set $\tilde{S}=S^{-1}$. Then we have
  1. 1.
    $\tilde{S}_{0}=1$, $\tilde{S}_{1}=0$.
  2. 2.
    $\tilde{S}_{-t}$ is a HCP from $Hcpc(q)$ for any $t\geq 0$.
  3. 3.
    If $S_{-t}\neq 0$ then $\frac{t}{u}-1<Sdeg_{A}(\tilde{S}_{-t})<t$ for any $t>0$.
  4. 4.
    $\tilde{S}_{-t}$ is totally free of $B_{j}$ for any $t\geq 0$.
Proof.
We can present $S$ as $S=1-S_{-}$, where $\mathop{\rm\bf ord}(S_{-})<-1$. Then $S^{-1}=1+\sum_{i=1}^{\infty}S_{-}^{i}$, since this series is well defined in the ring $\hat{D}_{1}^{sym}$. Hence $\tilde{S}_{0}=1$, $\tilde{S}_{1}=0$.
Note that any homogeneous component $\tilde{S}_{-t}$, $t>0$ is a finite sum of products of homogeneous components $S_{q}$:
$\tilde{S}_{-t}=\sum_{i=0}^{[t/2]}(S_{-}^{i})_{-t}=\sum_{i=1}^{[t/2]}\sum_{q_{1}+\ldots+q_{i}=t,q_{i}>0}(S_{-})_{-q_{1}}\ldots(S_{-})_{-q_{i}},$
and $(S_{-})_{j}=\pm S_{j}$, thus $\tilde{S}_{-t}$ is a HCP for any $t\geq 0$. Since all $S_{q}$ are totally free of $B_{j}$ by proposition 2.5, $\tilde{S}_{-t}$ is totally free of $B_{j}$ for any $t$ by lemma 2.12. By lemmas 2.10 and 2.5 we have $Sdeg_{A}((S_{-})_{-q_{1}}\ldots(S_{-})_{-q_{t}})<q_{1}+\ldots+q_{i}=t$, and the second inequality of item 3 follows from proposition 2.4.
Remark 2.6.
Note that the statement and proof of proposition 2.5 remain valid also for regular operators $Q\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ whose homogeneous components are HCP from $Hcpc(q)$ totally free of $B_{j}$ and with $Sdeg_{A}(Q_{q-i})\leq i-2$ for any $i>1$, cf. theorem 2.2 below.
Note also that operators $S$, $\tilde{S}$ from these statements are defined over the same field $K$ as the operator $Q$, though their homogeneous components written as HCPs need a formal extension of scalars to be presented in the G-form or standard form.
2.5 Some necessary conditions on normal forms
Let $Q\in D_{1}$ be a normalized operator as in the previous section. Let $P\in D_{1}$ be a monic operator of positive order.
Definition 2.7.
For a given pair of monic operators $Q,P\in D_{1}$ with $\mathop{\rm\bf ord}(Q)=\deg(Q)=q\geq 0$, $\mathop{\rm\bf ord}(P)=\deg(P)=p\geq 0$ we define a normal form of $P$ with respect to $Q$ as the operator $P^{\prime}:=S^{-1}PS$, where $S$ is a Schur operator for $Q$, i.e. $S^{-1}QS=\partial^{q}$.
Remark 2.7.
The Schur operator is not uniquely defined, but up to multiplication by the elements of the centralizer $C(\partial^{q})\subset\hat{D}_{1}^{sym}$ from proposition 2.1. By this reason the normal form of the operator $P$ is not uniquely defined, but up to conjugation by elements from this centralizer.
In the same way we can define normal forms for any regular operators $Q,P\in\hat{D}_{1}^{sym}$. However, by technical reasons we restrict ourself to the case of differential operators (because Schur operators of differential operators satisfy specific properties).
In this section we establish several necessary conditions on homogeneous components of a normal form $P^{\prime}$.
First note that all homogeneous components of $P$ are HCP, and the homogeneous decomposition of $P$ written in the G-form for all homogeneous components looks like
$P=P_{p}+P_{p-1}+\ldots=\partial^{p}+\theta_{0;p-1}\partial^{p-1}+\ldots+(\theta_{0;p-i}+\ldots+\theta_{i-1;p-i}(\Gamma)_{i-1})\partial^{p-i}+\ldots,$
i.e. $Sdeg_{A}(P_{p-i})<i$ and $Sdeg_{B}(P_{p-i})=-\infty$ for all $i>0$. Besides, all homogeneous components of $P$ are totally free of $B_{j}$.
Definition 2.8.
[…]
  1. 1.
    $P_{t}$ is a HCP from $Hcpc(q)$ for all $t$;
  2. 2.
    $P_{t}$ is totally free of $B_{j}$ for all $t$;
  3. 3.
    $Sdeg_{A}(P_{\mathop{\rm\bf ord}(P)-i})<i+k$ for all $i>0$;
  4. 4.
    $\sigma(P)$ does not contain $A_{q;i}$, $Sdeg_{A}(\sigma(P))=k$.
Example 2.4.
From previous section we know that $S,S^{-1}$ satisfy condition $A_{q}(0)$, where $S$ is any monic Schur operator for normalised $Q\in D_{1}$.
With the help of this definition we can prove the following criterion.
Theorem 2.2.
The following statements are equivalent:
  1. 1.
    $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ is a differential operator (i.e. $P\in D_{1}\otimes_{K}\tilde{K}$) with constant highest symbol.
  2. 2.
    $\forall p>1$, $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ satisfies condition $A_{p}(0)$ with an extra property: all homogeneous components $P_{j}$ don’t contain $A_{i}$.
  3. 3.
    $\exists p>1$, $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ satisfies condition $A_{p}(0)$ with an extra property: all homogeneous components $P_{j}$ don’t contain $A_{i}$.
Proof.
To simplify notations, we’ll assume in the course of proof that $D_{1}$ is defined over $\tilde{K}$ (i.e. we assume $K=\tilde{K}$).
1$\Rightarrow$2: It is easy to see that any differential operator $P\in D_{1}$ with constant highest symbol, i.e. $P=a_{q}\partial^{q}+\sum_{i=1}^{q}a_{q-i}\partial^{q-i}$ with $a_{q}\in K$, satisfies condition $A_{p}(0)$ for any $p>0$ with an extra property: all homogeneous components $P_{i}$ don’t contain $A_{i}$, cf. lemma 2.6.
2$\Rightarrow$3 is obvious.
[…]
In the case when $r<0$ the proof is by induction on $r$. Consider first $P_{-1}$; let’s write it in the G-form. If $P_{-1}\neq 0$, according to the assumptions, suppose
$P_{-1}=\sum_{0\leq m\leq d_{-1}}p_{m,0;-1}\Gamma_{m}D^{-1}.$
Since it’s totally free of $B_{j}$, there won’t be $B_{j}$ in $P_{-1}D$. Since
$P_{-1}D=-p_{0,0;-1}B_{1}+\sum_{0\leq m\leq d_{-1}}p_{m,0;-1}\Gamma_{m}$
according to Lemma 2.10 (see the calculations in the proof of item 3), hence by the uniqueness of HCPC (c.f. Lemma 2.6) we know $p_{0,0;-1}=0$. Thus
$P_{-1}=\sum_{1\leq m\leq d_{-1}}p_{m,0;-1}\Gamma_{m}D^{-1}\\ =\sum_{1\leq m\leq d_{-1}}p_{m,0;-1}(x\partial)^{m}D^{-1}\in D_{1}=\sum_{1\leq m\leq d_{-1}}p_{m,0;-1}(x\partial)^{m-1}x\in D_{1}$
For $r>1$, consider $P_{-r}=\sum_{0\leq m\leq d_{-r}}p_{m,0;-r}\Gamma_{m}D^{-r}$. By the same reason $P_{-r}D^{r}$ doesn’t contain $B_{1}$, so we get $p_{0,0,-r}=0$. Hence
$P_{-r}=\sum_{1\leq m\leq d_{-r}}p_{m,0;-r}\Gamma_{m}D^{-r}=\sum_{1\leq m\leq d_{-r}}p_{m,0;-r}\Gamma_{m-1}xD^{-r+1}\\ =\sum_{1\leq m\leq d_{-r}}p_{m,0;-r}x(\Gamma_{1}+1)^{m-1}D^{-r+1}=x\sum_{0\leq n\leq d_{-r}-1}\tilde{p}_{n,0;-r}\Gamma_{n}D^{-r+1}$
where
$\tilde{p}_{n,0;-r}=\sum_{m=1}^{d_{-r}}\binom{m-1}{n}p_{m,0;-r}$
Denote $H=\sum_{0\leq n\leq d_{-r}-1}\tilde{p}_{n,0;-r}\Gamma_{n}D^{-r+1}$, then $P_{-r}=xH$, it is written in G-form and totally free of $B_{j}$. Then by Lemma 2.14 (see below) $H$ is also totally free of $B_{j}$. By induction, $H\in D_{1}\Rightarrow P_{-r}=xH\in D_{1}$.
Since $P$ satisfies $A_{p}(0)$, we have for all $i>0$ $Sdeg_{A}(P_{\mathop{\rm\bf ord}(P)-i})<i\Rightarrow deg(P_{\mathop{\rm\bf ord}(P)-i})<\mathop{\rm\bf ord}(P)$. Also we have $P_{\mathop{\rm\bf ord}(P)}=\sigma(P)\in D_{1}$ with $deg(\sigma(P))=\mathop{\rm\bf ord}(\sigma(P))$. Thus
$P=\sum_{i=0}^{\infty}P_{\mathop{\rm\bf ord}(P)-i}\in D_{1},$
$deg(P)=\mathop{\rm\bf ord}(P)$ and $P$ has a highest constant symbol. ∎
Lemma 2.14.
Suppose $H$ is a HCP from $Hcpc(p)$ and doesn’t contain $A_{i}$, with $Sdeg_{A}(H)<\infty$. Then $H$ is totally free of $B_{j}$ if and only if $xH$ is totally free of $B_{j}$.
Proof.
Suppose $H$ is totally free of $B_{j}$, thus $H$ should be in the form of
$H=\sum_{0\leq m\leq d}h_{m,0;r}\Gamma_{m}D^{r}$
we want to show $xH$ is totally free of $B_{j}$, i.e to show for any $k\in\mathbb{Z}$, $xHD^{k}$ doesn’t contain $B_{j}$. Since $H$ is totally free of $B_{j}$, we know $Sdeg_{B}(HD^{k})=-\infty$. According to Lemma 2.10 (notice $-\mathop{\rm\bf ord}(\int)=1$), we know
$Sdeg_{B}(\int HD^{k})\leq 1$
Hence if we write $\int HD^{k}$ into G-form, it should be like
$\int HD^{k}=(\cdots)D^{r+k-1}+\lambda B_{1}D^{r+k-1}$
Where $\cdots$ means the polynomial of $\Gamma_{1}=(x\partial)$ and $\lambda\in\tilde{K}$. But notice that $\partial B_{1}=0$, so the terms with $B_{1}$ will be eliminate in $xHD^{k}$, i.e.
$xHD^{k}=(x\partial)(\int HD^{k})=(\cdots)D^{r+k-1}$
Hence $xH$ is totally free of $B_{j}$.
On the other hand, if we know $xH$ is totally free of $B_{j}$, suppose $H$ is not totally free of $B_{j}$, hence there exist $k$, such that $HD^{k}$ contains $B_{j}$, suppose $Sdeg_{B}(HD^{k})=j>0$, with
$HD^{k}=\sum_{m=0}^{d_{r}}\tilde{h}_{m,0;r}\Gamma_{m}D^{r+k}+\sum_{t=1}^{j-1}g_{t;r+k}B_{t}D^{r+k}+\lambda B_{j}D^{r+k}$
Since
$xB_{j}D^{r+k}=\Gamma_{1}\int B_{j}D^{r+k}=\Gamma_{1}B_{j+1}\int D^{r+k}=jB_{j+1}\int D^{r+k}$
Notice that $\int D^{r+k}=D^{r+k-1}$ when $r+k\leq 0$ and $(1-B_{1})D^{r+k-1}$ when $r+k>0$, but $j>1$ and we know $B_{j+1}B_{1}=\delta_{1}^{j+1}B_{1}=0$, hence $B_{j+1}\int D^{r+k}=B_{j+1}D^{r+k-1}$, thus
$xHD^{k}=\sum_{m=0}^{d_{r}+1}\tilde{\tilde{h}}_{m,0;r}\Gamma_{m}D^{r+k-1}+\sum_{t=1}^{j}\tilde{g}_{t;r+k}B_{t}D^{r+k-1}+j\lambda B_{j+1}D^{r+k-1}$
This is a contradiction with $xH$ totally free of $B_{j}$. ∎
Remark 2.8.
Note that the criterion 2.2 holds also if $P$ is defined over $K$. In this case in items 2 and 3 we need to add that $P_{j}$ are defined over $K$. The proof remains the same.
Lemma 2.15.
Suppose $P,Q\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ satisfy conditions $A_{q}(k_{1})$, $A_{q}(k_{2})$ correspondingly. Then $PQ$ satisfies condition $A_{q}(k_{1}+k_{2})$.
Proof.
Let $p=\mathop{\rm\bf ord}(P)$, $q=\mathop{\rm\bf ord}(Q)$. Since $\sigma(P)$, $\sigma(Q)$ does not contain $A_{q;i}$ and $B_{j}$, they are differential operators, i.e. belong to $D_{1}$, and therefore $0\neq\sigma(P)\sigma(Q)=\sigma(PQ)$, cf. remark 2.2, and thus $\sigma(PQ)$ does not contain $A_{q;i}$ and $B_{j}$ and $\mathop{\rm\bf ord}(PQ)=p+q$. Moreover, in this case $Sdeg_{A}(\sigma(PQ))=Sdeg_{A}(\sigma(P))+Sdeg_{A}(\sigma(Q))=k_{1}+k_{2}$, cf. formulae (2.9). For other homogeneous components of $PQ$ we have
$(PQ)_{p+q-i}=\sum_{i_{1}+i_{2}=i}P_{p-i_{1}}Q_{q-i_{2}},$
is a HCP from $Hcpc(q)$ and
$Sdeg_{A}(PQ)_{p+q-i}\leq\max\{Sdeg_{A}(P_{p-i_{1}}Q_{q-i_{2}})\}<i_{1}+k_{1}+i_{2}+k_{2}=i+k_{1}+k_{2}$
for all $i>0$ by lemmas 2.7, 2.10. Besides, $(PQ)_{p+q-i}$ is totally free of $B_{j}$ by lemma 2.12 for all $i>0$. ∎
Corollary 2.4.
Suppose $Q\in D_{1}$ is a monic operator with $\mathop{\rm\bf ord}(Q)=\deg(Q)=q>0$. Suppose $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ satisfies condition $A_{q}(0)$, $\mathop{\rm\bf ord}(P)=p$. Put $P^{\prime}=S^{-1}PS$, where $S$ is a Schur operator for $Q$ from proposition 2.5.
Then $P^{\prime}$ satisfies condition $A_{q}(0)$.
Proof.
By proposition 2.5 and corollaries 2.1, 2.3 the operators $S,S^{-1}$ satisfy condition $A_{q}(0)$. So, our claim immediately follows from lemma 2.15.
Remark 2.9.
Note that if $P$ is defined over $K$, then $P^{\prime}$ will be defined over $K$ too.
3 Normal forms for commuting operators
Let $Q\in D_{1}$ be a normalized operator as in section 2.4. Let $P\in D_{1}$ be a monic operator of positive order $p$ and $[P,Q]=0$. Fix $k=q=\mathop{\rm\bf ord}(Q)$, suppose $\tilde{K}$ is an algebraic closure of $K$, $A_{i}:=A_{k;i}$ as above.
In this section we give a convenient description of the centraliser $C(\partial^{q})\subset\hat{D}_{1}^{sym}\hat{\otimes}\tilde{K}$ and of normal forms of the operator $P$ with respect to $Q$. With the help of this description we give a new parametrisation of torsion free sheaves of rank one with vanishing cohomology groups on a projective curve (according to the well known classification theory of commuting ordinary differential operators such sheaves describe rank one subrings of commuting operators with a given spectral curve, cf. [43, Th. 10.26]).
Consider the subring $\tilde{K}[A_{1},\ldots,A_{k-1}]\subset\hat{D}_{1}^{sym}$. Clearly, $\tilde{K}[A_{1},\ldots,A_{k-1}]\cong\tilde{K}[x]/(x^{k}-1)$. Note that we have an isomorphism of $\tilde{K}$-algebras
$\Phi:\tilde{K}[A_{1},\ldots,A_{k-1}]\rightarrow\tilde{K}^{\oplus k},\quad P\mapsto(1\circ P,\ldots,\partial^{k-1}\circ P)$
(here $\partial^{l}\circ P$ is the notation from definition 2.2, i.e. if $P=\sum_{i=0}^{k-1}p_{i}A_{i}$, then $\partial^{l}\circ P=\sum p_{i}\xi^{il}$, and $\tilde{K}^{\oplus k}$ denotes the semisimple algebra - the direct sum of algebras $\tilde{K}$). Indeed, $\Phi$ is obviously linear, and, since
$\Phi(h_{l}A_{1}^{l}\cdot m_{t}A_{1}^{t})=(h_{l}m_{t},h_{l}m_{t}\xi^{l+t},\ldots,h_{l}m_{t}\xi^{(l+t)(k-1)})=\Phi(h_{l}A_{1}^{l})\cdot\Phi(m_{t}A_{1}^{t}),$
$\Phi$ is a $\tilde{K}$-algebra homomorphism. It’s easy to see that it is surjective and injective.
Remark 3.1.
Note that, by definition of $\Phi$, if $P\in\tilde{K}[A_{1}]$ belong to $\hat{D}_{1}^{sym}$ (i.e. all coefficients of the series representing $P$ in $\hat{D}_{1}^{sym}$ belong to $K$), then $\Phi(P)\in K^{\oplus k}$, i.e. $\Phi$ is compatible with the extension of scalars of $\hat{D}_{1}^{sym}$.
Now consider the skew polynomial ring ${{\cal R}}=\tilde{K}^{\oplus k}[D,\sigma]$, where $\sigma(a_{0},\ldots,a_{k-1})=(a_{k-1},a_{0},\ldots,a_{k-2})$  [^5] We use a standard notations and constructions from the books [9] and [23]. A short self contained exposition of all necessary constructions and facts see e.g. in [43], Ch. 2,3 and 13.1. Recall that this notation means that we have the following commutation relation between $D$ and $(a_{0},\ldots,a_{k-1})$: $(a_{0},\ldots,a_{k-1})D=D\sigma(a_{0},\ldots,a_{k-1})$. . It is a right and left noetherian ring. The multiplicatively closed subset $S=\{D^{k},k\geq 0\}$ obviously satisfies the right Ore condition, consists of regular elements, and ${\bf ass}(S)=0$. So, the right quotient ring $\mathfrak{B}={{\cal R}}_{S}$ exists. Clearly,
$\mathfrak{B}\simeq\tilde{K}^{\oplus k}[D,D^{-1}]=\{\sum_{l=M}^{N}P_{l}D^{l}|\quad P_{l}\in\tilde{K}^{\oplus k}\}\simeq\tilde{K}[A_{1}][D,D^{-1}],$
i.e. any element from $\mathfrak{B}$ can be written as a Laurent polynomial, and the commutativity relations of polynomials are given above: $D^{-1}a=\sigma(a)D^{-1}$, $a\in\tilde{K}^{\oplus k}$.
Let $C(\mathfrak{B})$ be the center of $\mathfrak{B}$, Obviously, $D^{k},D^{-k}\in C(\mathfrak{B})$. Since an element $(h_{0},h_{1},\dots,h_{k-1})\in\tilde{K}^{\oplus k}$ doesn’t commute with any $D^{l}$, if not all of $h_{i}$ equal to each other, we have
$C(\mathfrak{B})\cong\tilde{K}[D^{k},D^{-k}],$
where $\tilde{K}$ is diagonally embedded into $\tilde{K}^{\oplus k}$. So, $\mathfrak{B}$ is a finite dimensional algebra over its center.
Lemma 3.1.
There is an isomorphism of $\tilde{K}$-algebras
$\mathfrak{B}\cong M_{k}(C(\mathfrak{B})).$
Proof.
Consider $\psi:\mathfrak{B}\rightarrow M_{k}(C(\mathfrak{B}))$, with
$\psi\begin{pmatrix}h_{0}\\ h_{1}\\ \cdots\\ h_{k-1}\end{pmatrix}=\begin{pmatrix}h_{0}&&&\\ &h_{1}&&\\ &&\cdots&\\ &&&h_{k-1}\end{pmatrix}\quad\psi(D)=T:=\begin{pmatrix}&1&&\cdots&\\ &&1&\cdots&\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ &&&\cdots&1\\ D^{k}&&&\cdots&\end{pmatrix}$
with $\psi(D^{l})=T^{l}$, and extend $\psi$ by linearity. Direct calculations show that $\psi$ is a homomorphism of $\tilde{K}$-algebras. Now consider
$H_{ij}=\begin{cases}(0,\cdots,1,\cdots,0)D^{j-i}&i\leq j\\ (0,\cdots,1,\cdots,0)D^{j+k-i}&i>j\end{cases}$
where $1$ is located at the $i$th entry, so that $\psi(H_{ij})=E_{ij}D^{k}$, $i>j$ or $E_{ij}$, $i\leq j$. This means $\psi$ is a surjective. Obviously, $\mathfrak{B}$ has dimension $k^{2}$ over $C(\mathfrak{B})$ and $\dim_{C(\mathfrak{B})}(M_{k}(C(\mathfrak{B})))=k^{2}$, too. Besides, $\psi(C(\mathfrak{B}))=C(M_{k}(C(\mathfrak{B})))=C(\mathfrak{B})\cdot Id$. So $\psi$ is an isomorphism of $\tilde{K}$-algebras. ∎
Now consider the ring of skew pseudo-differential operators
$E_{k}:=\tilde{K}[\Gamma_{1},A_{1}]((\tilde{D}^{-1}))=\{\sum_{l=M}^{\infty}P_{l}\tilde{D}^{-l}|\quad P_{l}\in\tilde{K}[\Gamma_{1},A_{1}]\}\simeq\tilde{K}^{\oplus k}[\Gamma_{1}]((\tilde{D}^{-1}))$
with the commutation relation as above (here $\tilde{K}[\Gamma_{1},A_{1}]$ is a commutative subring in $\hat{D}_{1}^{sym}$): [^6] The ring $E_{k}$ is constructed in the same way as splittable local skew fields, cf. [39]
$\tilde{D}^{-1}a=\sigma(a)\tilde{D}^{-1},\quad a\in\tilde{K}[\Gamma_{1},A_{1}]\quad\mbox{where \quad}\sigma(A_{1})=\xi^{-1}A_{1},\quad\sigma(\Gamma_{1})=\Gamma_{1}+1.$
The ring $E_{k}$ is endowed with a natural discrete pseudo-valuation, which we will denote as $-\mathop{\rm ord}\nolimits_{\tilde{D}}$ (i.e. $\mathop{\rm ord}\nolimits_{\tilde{D}}(\sum_{l=M}^{\infty}P_{l}\tilde{D}^{-l})=M$). We extend the usual terminology used in this paper also for operators from $E_{q}$ (such as the notion of the highest coefficient, monic operators, etc.)
Denote $\widehat{Hcpc}_{B}(k)$ as the $\tilde{K}$-subalgebra in $\hat{D}_{1}^{sym}\hat{\otimes}\tilde{K}$ consisting of operators whose homogeneous components are HCPs totally free of $B_{j}$ (cf. lemma 2.12).
Lemma 3.2.
The map
$\hat{\Phi}:\widehat{Hcpc}_{B}(k)\longrightarrow E_{k},$
defined on monomial HCPs from $\widehat{Hcpc}_{B}(k)$ as $\hat{\Phi}(aA_{j}\Gamma_{i}D^{l}):=a\Phi(A_{j})\Gamma_{i}\tilde{D}^{l}$ and extended by linearity on the whole $\tilde{K}$-algebra $\widehat{Hcpc}_{B}(k)$, is an embedding of $\tilde{K}$-algebras.
Proof.
Since all operators in $\widehat{Hcpc}_{B}(k)$ are totally free of $B_{j}$, the proof is almost obvious in view of formulae (2.6)-(2.9). ∎
Remark 3.2.
Again as in remark 3.2, if $P\in\hat{D}_{1}^{sym}\cap\widehat{Hcpc}_{B}(K)$, then $\hat{\Phi}(P)$ will be an operator with coefficients from $K$.
Remark 3.3.
Note that the ring $\mathfrak{B}$ is naturally embedded into the ring $E_{k}$. In particular, all normal forms of a monic differential operator $P$ with respect to the commuting with $P$ differential operator $Q$, or more generally any operator from the centralizer $C(\partial^{k})$, can be embedded in $E_{k}$ via $\hat{\Phi}$. Recall that any such normal form is defined up to conjugation with an operator from the centralizer $C(\partial^{k})$, and $C(\partial^{k})\subset\widehat{Hcpc}_{B}(k)$ by corollary 2.2.
Next, note that any operator from the centralizer $C(\partial^{k})$ embedded into $\mathfrak{B}$ goes under the isomorphism $\psi:\mathfrak{B}\cong M_{k}(C(\mathfrak{B}))$ to a matrix with entries belonging to the subring $\tilde{K}[D^{k}]$. Indeed, we know from lemma 2.4 that all coefficients of homogeneous terms of negative order satisfy the following conditions: if $P\in\hat{\Phi}(C(\partial^{k}))\subset\mathfrak{B}$, and $p_{l}\in\tilde{K}^{\oplus k}$ denote its coefficients, then always $\tilde{p}_{l,j}=$ for $j=0,\ldots,-l-1$ if $l<0$. Therefore all homogeneous terms of negative order will go to matrices with constant coefficients (see the proof of lemma 3.1), and terms of non-negative order go to matrices with entries belonging to the subring $\tilde{K}[D^{k}]$. Finally, we can observe that in fact we get an isomorphism $C(\partial^{k})\simeq M_{k}(\tilde{K}[D^{k}])$.
Therefore, for any $P\in C(\partial^{k})$ the characteristic polynomial $\det(\psi(P)-\lambda)\in\tilde{K}[\lambda,D^{k}]$ defines an algebraic relation between $P$ and $\partial^{k}$ or, in degenerate cases, it defines an algebraic dependence over $\tilde{K}$. Note that if $P^{\prime}$ is a normal form of a monic differential operator $P$ with respect to the commuting with $P$ differential operator $Q$, and the order of $P$ is coprime with the order of $Q$, then $P^{\prime}$ is monic and $\det(\psi(P)-\lambda)\in\tilde{K}[\lambda,D^{k}]$ is a polynomial of the form similar to the Burchnall-Chaundy polynomial [^7] Recall the famous Burchnall-Chaundy lemma: if $P,Q\in D_{1}$ are differential operators of coprime orders $p$ and $q$, then they are algebraically dependent, and satisfy a polynomial relation of the form $f(X,Y)=\alpha X^{q}\pm Y^{p}+\ldots=0$, where $GCD(p,q)=1$ and $\lambda\neq 0$ (here the weighted degree of $f$ is $pq$, and $\ldots$ mean terms of lower weighted degree, where the weight of $X$ is $p$, and the weight of $Y$ is $q$; in particular, for coprime $p$ and $q$ it is automatically irreducible). , i.e. $\det(\psi(P)-\lambda)=\pm\lambda^{q}\pm D^{kp}+\ldots$, in particular it is irreducible and therefore coincides with the Burchnall-Chaundy polynomial for $P$ and $Q$ up to a multiplicative constant.
Embedding normal forms to the ring $E_{q}$, we can transform them to a simpler form by conjugation:
Lemma 3.3.
Assume $\tilde{P}\in\tilde{K}^{\oplus k}((\tilde{D}^{-1}))\subset E_{q}$ is a monic operator with $\mathop{\rm ord}\nolimits_{\tilde{D}}(\tilde{P})=p$. Let $p=dn$, $q=dm$, where $d=GCD(p,q)$.
Then there exists a monic invertible operator $S\in\tilde{K}^{\oplus k}((\tilde{D}^{-1}))\subset E_{q}$ with $\mathop{\rm ord}\nolimits_{\tilde{D}}(S)=0$ such that all coefficients of the operator $S^{-1}\tilde{P}S$ commute with $\tilde{D}^{d}$.
Proof.
We will find the operator $S$ as the limit of a Cauchy sequence888with respect to the topology defined by the pseudo-valuation $-\mathop{\rm ord}\nolimits_{\tilde{D}}$. Assume
$\tilde{P}=\tilde{D}^{dn}+\sum_{i=1}^{\infty}p_{i}\tilde{D}^{dn-i}\quad p_{i}\in\tilde{K}^{\oplus k},$
and let $p_{l}=(p_{l,0},\ldots,p_{l,q-1})$ be the first coefficient not commuting with $\tilde{D}^{d}$. Consider the operator $S_{l}:=1+s_{l}\tilde{D}^{-l}$, where $s_{l}=(s_{l,0},\ldots,s_{l,q-1})$. Then an easy direct calculation shows that
$S_{l}^{-1}\tilde{P}S_{l}=\tilde{D}^{dn}+\ldots+(p_{l}-s_{l}+\sigma^{p}(s_{l}))\tilde{D}^{dn-l}+\mbox{terms of lower order},$
where $\ldots$ are the same terms of $\tilde{P}$. Consider the following $d$ systems of linear equations: for $i=0,\ldots,d-1$ [mod $d$] set $b_{i}:=\sum_{k=0}^{m-1}p_{l,[(i+pk)\mbox{mod $q$}]}$, then the $i$-th system is (below all indices are considered modulo $q$)
$\begin{cases}p_{l,i}-s_{l,i}+s_{l,i+p}=b_{i}/m\\ p_{l,i+p}-s_{l,i+p}+s_{l,i+2p}=b_{i}/m\\ \ldots\\ p_{l,i+(m-1)p}-s_{l,i+(m-1)p}+s_{l,i}=b_{i}/m\end{cases}$
This system is solvable and give explicit formulae for unknown variables $s_{l,j}$: for $r=1,\ldots,(m-1)$ from the first $(m-1)$ equations we get
$s_{l,i+rp}=rb_{i}/m+s_{l,i}-(\sum_{j=1}^{r}p_{l,i+jp})$
($s_{l,i}$ is a free parameter), and the last equation becomes an identity under substitution of these values. Taking such $s_{l}$ we get
$\tilde{p}_{l}:=p_{l}-s_{l}+\sigma^{p}(s_{l})=(b_{0}/m,\ldots,b_{d-1}/m,b_{0}/m,\ldots,b_{d-1}/m,\ldots)$
which obviously commutes with $\tilde{D}^{d}$.
It’s easy to see that the system $\{\prod_{i=1}^{j}S_{i}\}$, $j\geq 1$ is a Cauchy system in $E_{q}$, and its limit $S$ is the needed operator. ∎
Corollary 3.1.
Let $B^{\prime}\subset\tilde{K}^{\oplus k}((\tilde{D}^{-1}))\subset E_{q}$ be a commutative subring containing $\tilde{D}^{q}$ and a monic operator $P^{\prime}$ of order $p$. Suppose $GCD(p,q)=1$.
Then $S^{-1}B^{\prime}S\subset\tilde{K}((\tilde{D}^{-1}))$ for an operator $S$ from lemma 3.3.
The proof is obvious.
Normal forms of commuting differential operators can be normalised (s.t. the normalised normal form will be uniquely defined up to conjugation). If the orders of these operators are coprime, such a normalisation can be easily presented:
Lemma 3.4.
[…]
Then there is a normalised normal form $P^{\prime}$ of $P$ with respect to $Q$ uniquely defined up to conjugation by an element from the centraliser $C(\partial^{q})$. Namely, we can explicitly describe its image under the embedding $\hat{\Phi}$ as follows:
$\hat{\Phi}(P^{\prime})=\tilde{D}^{p}+\sum_{l=-q+1}^{p-1}p_{l}\tilde{D}^{l},\quad p_{l}=(p_{l,0},\ldots,p_{l,q-1})\in\tilde{K}^{\oplus q},$
where
$p_{l,j}=\begin{cases}\mbox{If $p\geq q$ then}\begin{cases}\mbox{$0$ for $j\in N_{p-l}$}\quad l>p-q\\ \mbox{$0$ for $j=0,\ldots,-l-1$}\quad l<0\end{cases}\\ \mbox{If $p<q$ then}\begin{cases}\mbox{$0$ for $j\in N_{p-l}$}\quad l\geq 0\\ \mbox{$0$ for $j\in N_{p-l}$ or $j=0,\ldots,-l-1$}\quad p-q<l<0\\ \mbox{$0$ for $j=0,\ldots,-l-1$}\quad l\leq p-q\end{cases}\end{cases}$
Let’s call coefficients $p_{l,j}$ supplementary to the list above as coordinates of $P^{\prime}$. If $P^{\prime}$ and $P^{\prime\prime}$ are two operators with different values of coordinates, then there are no invertible operators $S\in C(\partial^{q})$ such that $P^{\prime}=S^{-1}P^{\prime\prime}S$.
Proof.
To show that such normalised form exist, we can follow the arguments in lemma 3.3. If $P^{\prime}$ is any given normal form of $P$ with respect to $Q$, we will find a monic operator of order zero $S\in C(\partial^{q})$ (thus it will be automatically invertible) step by step, as a product of operators $S_{l}=1+s_{l}\int^{l}\in C(\partial^{q})$, $l=1,\ldots,q-1$. Since $\hat{\Phi}$ is an embedding of rings, we can provide all calculations in the ring $E_{q}$. Note that by definition of $\hat{\Phi}$ and from lemma 2.4 we get $\hat{\Phi}(S_{l})=1+\tilde{s}_{l}\tilde{D}^{-l}$, where $\tilde{s}_{l,j}=0$ for $j=0,\ldots,l-1$. Using calculations from lemma 3.3, we get for $\tilde{P}^{\prime\prime}:=\hat{\Phi}(S_{l}^{-1}P^{\prime}S_{l})=\tilde{D}^{p}+\ldots+\tilde{p}_{l}\tilde{D}^{p-l}+\ldots$ that
$\begin{cases}p_{l,0}-\tilde{s}_{l,0}+\tilde{s}_{l,p}=\tilde{p}_{l,0}\\ p_{l,p}-\tilde{s}_{l,p}+\tilde{s}_{l,2p}=\tilde{p}_{l,1}\\ \ldots\\ p_{l,(q-1)p}-\tilde{s}_{l,(m-1)p}+\tilde{s}_{l,0}=\tilde{p}_{l,q-1}\end{cases}$
because $GCD(p,q)=1$ (here again all indices are taken modulo $q$). Now we have $\tilde{s}_{l,0}=0$ for all $l=1,\ldots,q-1$, and, starting with the first equation, we can see that for $j$-th equation, where $j\in N_{l}$, we can find $\tilde{s}_{l,jp}$ such that $\tilde{p}_{l,j-1}=0$. On the other hand, $\tilde{s}_{l,jp}=0$ for all $j\notin N_{l}$. Thus, we uniquely determine $\tilde{s}_{l}$ such that conditions of lemma satisfied for $l$-th homogeneous component of $\tilde{P}^{\prime\prime}$.
On the other hand, since $\tilde{P}^{\prime\prime}\in C(\partial^{q})$, we know from lemma 2.4 that always $\tilde{p}_{l,j}=$ for $j=0,\ldots,-l-1$ if $l<0$ (in particular, the number of zero coefficients for $p-q<l<0$ in case $p<q$ is constant). Taking $S=\prod_{j=1}^{q-1}S_{j}$ we get the needed operator: the normal form $S^{-1}P^{\prime}S$ will satisfy all conditions as stated. This completes the proof of the first statement.
To prove the second statement (the uniqueness of such normalised normal form up to conjugation), first note that if $S\in C(\partial^{q})$ is invertible and $\mathop{\rm\bf ord}(S)=t>0$, then the coefficient of the HCP $S_{t}$ must be a zero divisor. Indeed, assume the converse. Suppose $\mathop{\rm\bf ord}(S^{-1})=r$. Then $(SS^{-1})_{t+r}\neq 0$ hence $r=-t<0$. From lemma 2.4 we know then the coefficient of the HCP $(S^{-1})_{r}$ is a zero divisor, hence the coefficient of the HCP $(S^{-1}S)_{t+r}=1$ must be a zero divisor, a contradiction. By the same reason there are no invertible $S$ with $\mathop{\rm\bf ord}(S)<0$ (because all its homogeneous components will be zero divisors by lemma 2.4).
Now suppose we have two normalised normal forms $P^{\prime}$ and $P^{\prime\prime}$ with different values of coordinates, and $P^{\prime}=S^{-1}P^{\prime\prime}S$ for some $S\in C(\partial^{q})$. Suppose $\mathop{\rm\bf ord}(S)=t\geq 0$. From the equality we have $S_{t}\partial^{p}=\partial^{p}S_{t}$. Since $p$ and $q$ are coprime, the coefficient of $S_{t}$ must be a constant. Thus, it is not a zero divisor and therefore $t=0$. Without loss of generality we can assume $S_{0}=1$. But then the same equations as above show that all other homogeneous components of $S$ must be zero (as all coefficients of $P^{\prime}$ and $P^{\prime\prime}$ from the list in the formulation are zero) and therefore the equality $P^{\prime}=S^{-1}P^{\prime\prime}S$ is impossible. ∎
Now we are ready to give a description of a new parametrisation of torsion free sheaves of rank one with vanishing cohomology groups on a projective curve. First let’s recall the classification theorem of rank one commutative subrings of differential operators.
Theorem 3.1.
([43, Th. 10.26]) There is a one-to-one correspondence
$[B\subset{D}\mbox{\quad of rank $1$}]/\thicksim\longleftrightarrow[(C,p,{{\cal F}})\mbox{\quad of rank $1$}]/\simeq$
where
  • $[B]$ means a class of equivalent commutative elliptic subrings (i.e. $B$ containing a monic differential operator), where $B\sim B^{\prime}$ iff $B=f^{-1}B^{\prime}f$, $f\in D^{*}$.
[…]
  • $C$ is a (irreducible and reduced) projective curve over $K$, $p$ is a regular $K$-point and ${{\cal F}}$ is a torsion free sheaf of rank one with $H^{0}(C,{{\cal F}})=H^{1}(C,{{\cal F}})=0$ (a spectral sheaf).
  • $\simeq$ means a natural isomorphism of triples.
Remark 3.4.
The self contained proof of this theorem in such a form is given in [43, Ch. 9, 10], and it uses two other correspondences between these data and equivalence classes of Schur pairs (we recall the definition below, as they will play a key role in our statements). The proof given in [43] is an elaborated version of Mulase’s proof from [26] in a spirit of works [29], [28] and their higher dimensional generalisations in [40], [17].
Recall that this classification has a long history: the first classification results of commuting pairs of ODOs with coprime orders appeared in works [6]-[8]. After that Krichever in works [14], [15] gave classification of commutative subrings of ODOs of any rank in general position in terms of geometric data. His version of classification theorems had a more analytical nature. The other versions (have more algebraic nature) are due to Mumford [27], Drinfeld [11], Verdier [36] and Mulase [26] (cf. also an important paper by Segal and Wilson [34]).
Let $B\subset D_{1}$ be an elliptic subring. Then by Schur theory from [33], cf. [43, T. 4.6, C. 4.7], there exists an invertible operator $S=s_{0}+s_{1}\partial^{-1}+\ldots$ in the usual (Schur’s) ring of pseudo-differential operators $E=K[[x]]((\partial^{-1}))$ such that $A:=S^{-1}BS\subset K((\partial^{-1}))$. Consider the homomorphism (of vector spaces)
$E\rightarrow E/xE\simeq K((\partial^{-1}))$ (3.1)
(sometimes it is called the Sato homomorphism). It defines a structure of an $E$-module on the space $K((\partial^{-1}))$: for any $P\in K((\partial^{-1}))$, $Q\in E$ we put $P\cdot Q=PQ$ (mod $xE$).
Analogously, the homomorphism
$1\circ:\tilde{K}[A_{1}]((\tilde{D}^{-1}))\subset E_{q}\rightarrow\tilde{K}((\tilde{D}^{-1})),\quad\sum_{l}p_{l}\tilde{D}^{l}\mapsto\sum_{l}p_{l,0}\tilde{D}^{l}$
(cf. lemma 3.3) defines a structure of a $\tilde{K}[A_{1}]((\tilde{D}^{-1}))$-module on the space $\tilde{K}((\tilde{D}^{-1}))$: for any $P\in\tilde{K}((\tilde{D}^{-1}))$ and $Q\in\tilde{K}[A_{1}]((\tilde{D}^{-1}))$ we put $P\cdot Q=1\circ(PQ)$.
Now define the space $W:=F\cdot S\subset K((\partial^{-1}))$ (here $F$ is the same as in theorem 2.1). Note that $W$ is an $A$-module, where the module structure is defined via the multiplication in the field $K((\partial^{-1}))$ and this module structure is induced by the $E$-module structure on $K((\partial^{-1}))$, because $K((\partial^{-1}))\subset E$ and $W\cdot A=(F\cdot S)\cdot(S^{-1}BS)=F\cdot(BS)=(F\cdot B)\cdot S=F\cdot S=W$. Note also that the modules $W$ and $F$ are isomorphic ($W$ is an $A$-module, $F$ as a $B$-module, and clearly $A\simeq B$). For convenience of notation, we will replace $\partial^{-1}$ by $z$ in the field $K((\partial^{-1}))$, i.e. $A,W\subset K((z))\simeq K((\partial^{-1}))$.
Analogously, we can define the space $W^{\prime}:=F^{\prime}\cdot S\subset\tilde{K}((\tilde{D}^{-1}))$, where $F^{\prime}=\tilde{K}[\tilde{D}]$ and $S\in\tilde{K}[A_{1}]((\tilde{D}^{-1}))$ is an operator from lemma 3.3. If $B^{\prime}=\hat{\Phi}(S^{-1}BS)$, where $S$ is Schur operator from proposition 2.2, and $A^{\prime}:=S^{-1}B^{\prime}S$, where $S$ is an operator from corollary 3.1, then, clearly, the modules $F^{\prime}$ and $F$ are isomorphic ($F^{\prime}$ as a $B^{\prime}$-module, $F$ as a $B$-module), and $W^{\prime}$ and $F$ are isomorphic ($W^{\prime}$ is a $B^{\prime}$-module, $F$ as a $B$-module), so also $W^{\prime}\simeq W$. Moreover, all these modules are isomorphic as filtered modules (with respect to the order filtration).
For subrings in $K((z))$ we can introduce the same notion of rank as for subrings in $D_{1}$999here $z$ will play a role of $\partial^{-1}$ or $\tilde{D}^{-1}$:
Definition 3.1.
[…]
Definition 3.2.
Let $W$ be a $K$-subspace in $K((z))$. The support of an element $w\in W$ is its highest symbol, i.e. $\sup(w):=HT(w)z^{-\mathop{\rm ord}\nolimits(w)}$. The support of the space is
$\mathop{\rm Supp}W:=\langle\sup(w)|\quad w\in W\rangle$.
Definition 3.3.
An embedded Schur pair of rank $r$ is a pair $(A,W)$ consisting of
  • $A\subset K((z))$ a $K$-subalgebra of rank $r$ satisfying $A\cap K[[z]]=K$;
  • $W\subset K((z))$ a $K$-subspace with $\mathop{\rm Supp}W=K[z^{-1}]$
such that $W\cdot A\subseteq W$.
So, to any elliptic subring $B\subset D_{1}$ we can associate an embedded Schur pair.
Definition 3.4.
Two embedded Schur pairs $(A_{i},W_{i})$, $i=1,2$ of rank $r$ are equivalent if there exists an admissible operator $T$ such that $A_{1}=T^{-1}A_{2}T$, $W_{1}=W_{2}\cdot T$. An operator $T=t_{0}+t_{1}\partial^{-1}+\ldots$ is called admissible if $T^{-1}\partial T\in K((\partial^{-1}))$.
Theorem 3.2.
[…]
A self contained proof of this theorem see e.g. in [43, 10.3].
Let $C$ be a projective curve over $K$ and $p\in C$ be a regular $K$-point. Then $C_{0}:=C\backslash p$ is an affine curve. According to theorem 3.1 for any torsion free rank one sheaf ${{\cal F}}$ on $C$ there exists a normalised elliptic ring of differential operators $B\subset D_{1}$ defined uniquely up to a scale transform, such that $B$ is isomorphic to the ring of regular functions on $C_{0}$, $B\simeq{{\cal O}}_{C_{0}}(C_{0})$. Vice versa, any such ring is isomorphic to the ring of regular functions on some affine curve $C_{0}$, which can be compactified with the help of one regular $K$-point.
Definition 3.5.
We’ll call an affine curve $C_{0}$ over $K$ as affine spectral curve if it can be compactified with the help of one regular $K$-point, i.e. if there exists a projective curve $C$ over $K$ and a regular $K$-point $p$ such that $C_{0}\simeq C\backslash p$ (note that such $C$ is uniquely defined up to an isomorphism, see e.g. [12, Ch1., §6]).
Remark 3.5.
There is the Krichever map
$\chi_{0}:(C,p,{{\cal F}},\pi,\hat{\phi})\rightarrow(A,W)$
defined for any coherent torsion free sheaf ${{\cal F}}$ and any trivialisations $\pi,\hat{\phi}$ (for details see [43, Ch.10]). If ${{\cal F}}={{\cal O}}_{C}$, then $W=A$ and the rank of $A$ is $1$. Moreover, elements of $A$ are Laurent series expansions of functions from ${{\cal O}}_{C}(C\backslash p)$, s.t. the orders of elements of $A$ are the pole orders of corresponding functions. Differential operators corresponding to elements of $A$ via the correspondence from theorems 3.1, 3.2 have the same order as the elements.
If $\mathop{\rm rk}(A)=1$, then we can always choose a system of generators $a_{1},\ldots,a_{m}$ of $A$ as a $K$-algebra, such that $\mathop{\rm ord}\nolimits(a_{1})$ is coprime with the orders of other generators. Without loss of generality $a_{i}$ can be assumed to be monic. Conjugating $A$ by a suitable admissible operator (and using the usual Schur theory), we can get $a_{1}=z^{-\mathop{\rm ord}\nolimits(a_{1})}$.
Since any pair of generators $(a_{1},a_{i})$ correspond to some differential operators of coprime orders, they are algebraically dependent and satisfy some equations of Burchnall-Chaundy type $f_{i}(X,Y)=X^{q}\pm Y^{p}+\ldots=0$ (see remark 3.3). Vice versa, it’s easy to see, using standard arguments from Hensel’s lemma, that this equation uniquely determines a monic element $a_{i}$ from $A$ of a given order such that $f_{i}(a_{i},z^{-\mathop{\rm ord}\nolimits(a_{1})})=0$. Thus, the equations $f_{i}$ completely determine the subring $A$ in $K((z))$.
Let $B\subset D_{1}$ be an elliptic commutative subring of ODOs. Let $P_{1},\ldots,P_{m}$ be its monic generators over $K$ (any generators have constant highest coefficients, cf. [43, Ch.3]), such that $\mathop{\rm\bf ord}(P_{1})=q$ is coprime with the orders of $P_{2},\ldots,P_{m}$.
$B\simeq K[P_{1},\ldots,P_{m}]\simeq K[T_{1},\ldots,T_{m}]/I,$
where $I=(f_{1},\ldots,f_{k})$ is a prime ideal, $f_{i}\in K[T_{1},\ldots,T_{m}]$. By proposition 2.3 and lemma 3.4 there exists a uniquely determined Schur operator $S\in\hat{D}_{1}^{sym}$ such that $B^{\prime}:=S^{-1}BS\in C(\partial^{q})$ and $P_{2}^{\prime}=S^{-1}P_{2}S$ is a normalised normal form of $P_{2}$ with respect to $P_{1}$. From lemma 3.4 we immediately get that the coefficients of all other normal forms $P_{i}^{\prime}$ are uniquely determined. Obviously, $f_{i}(P_{1}^{\prime},\ldots,P_{m}^{\prime})=0$ for all $i$, and therefore define a set of equations on the coefficients of $\hat{\Phi}(P_{i}^{\prime})\in\mathfrak{B}$ in an affine space. Note that any point of the affine algebraic set determined by these equations defines a set of coefficients of commuting operators from $C(\partial^{q})$, and these operators generate a ring isomorphic to $B$101010with the isomorphism sending them to $T_{1},\ldots,T_{m}$.
Definition 3.6.
We’ll call a commutative ring $B^{\prime}\in C(\partial^{q})\subset\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ generated over $\tilde{K}$ by monic operators $P_{1}^{\prime}=\partial^{q},P_{2}^{\prime},\ldots,P_{m}^{\prime}$, where $P_{2}^{\prime}$ is normalised and $q$ is coprime with the orders of $P_{i}^{\prime}$, $i\geq 2$, as a normalised normal form with respect to a (ordered) set of generators $P_{1}^{\prime},\ldots,P_{n}^{\prime}$, and we’ll call the coefficients of the operators $P_{i}^{\prime}$ as coordinates of $B^{\prime}$ (cf. lemma 3.4).
We’ll denote by $X_{[B^{\prime}]}$ the corresponding affine algebraic set determined by the relations on coefficients of operators $P_{i}^{\prime}$ (it is defined by an isomorphism class of the ring $B^{\prime}$).
Lemma 3.5.
Let $f(X,Y)=X^{q}\pm Y^{p}+\ldots\in K[X,Y]$ be a Burchnall-Chaundy polynomial with coprime $p,q$. Assume $P^{\prime}\in C(\partial^{q})\subset\hat{D}_{1}^{sym}\hat{\otimes}\tilde{K}$ is a monic operator with $\mathop{\rm\bf ord}(P^{\prime})=p$ such that $f(P^{\prime},\partial^{q})=0$.
Then $a:=S^{-1}P^{\prime}S\in\tilde{K}((\tilde{D}^{-1}))$, where $S$ is an operator from lemma 3.3, and $a$ is the uniquely defined monic element in $\tilde{K}((\tilde{D}^{-1}))$, satisfying the equation $f(a,\tilde{D}^{q})=0$.
The proof is obvious in view of remark 3.5 and corollary 3.1.
Theorem 3.3.
Let $C_{0}$ be an affine spectral curve over $K$ and $C$ its one-point compactification. Assume
${{\cal O}}_{C}(C_{0})\simeq K[w_{1},\ldots,w_{m}]\simeq K[T_{1},\ldots,T_{m}]/I,$
where $I=(f_{1},\ldots,f_{k})$ is a prime ideal and the order of $w_{1}$ is coprime with the order of $w_{i}$, $i\geq 2$, and the images of $w_{i}$ under the Krichever map (after some choice of $\pi$, $\hat{\phi}$) are monic (cf. remark 3.5).
Then there exist normalised normal forms $B^{\prime}\simeq{{\cal O}}_{C}(C_{0})$ with respect to the ordered set of generators $P_{1}^{\prime}=\partial^{q},\ldots,P_{m}^{\prime}$, where $\mathop{\rm\bf ord}(P_{i}^{\prime})=\mathop{\rm ord}\nolimits(w_{1})$ for all $i\geq 1$, and there is a one to one correspondence between closed points of the affine algebraic set $X_{[{{\cal O}}_{C}(C_{0})]}$ and isomorphism classes of torsion free rank one sheaves ${{\cal F}}$ on $C$ with vanishing cohomologies $H^{0}(C,{{\cal F}})=H^{1}(C,{{\cal F}})=0$.
Proof.
Let ${{\cal F}}$ be a torsion free rank one sheaf on $C$ with vanishing cohomologies. By theorem 3.1 the triple $(C,p,{{\cal F}})$ corresponds to uniquely defined normalised commutative elliptic subring $B\subset D_{1}$ up to a scale transform. Besides, the generators $w_{i}$ corresponds to formally elliptic differential operators $P_{1},\ldots,P_{m}$ of the same orders, and there exists a scale transform that makes them monic. By proposition 2.3 and lemma 3.4 there exists a uniquely determined Schur operator $S\in\hat{D}_{1}^{sym}$ such that $B^{\prime}:=S^{-1}BS\in C(\partial^{q})$ and $P_{2}^{\prime}=S^{-1}P_{2}S$ is a normalised normal form of $P_{2}$ with respect to $P_{1}$, i.e. $B^{\prime}$ is a normalised normal form w.r.t. $P_{1}^{\prime},\ldots,P_{m}^{\prime}$ which have the same orders as $w_{i}$ or $P_{i}$. Note that the scale transform is compatible with conjugation by $S$ and that the coefficients of $\hat{\Phi}(P_{i}^{\prime})$ are invariant under any scale transform of $P_{i}^{\prime}$ for all $i$. Thus, ${{\cal F}}$ determines a closed point of $X_{[{{\cal O}}_{C}(C_{0})]}$.
Vice versa, any closed point of $X_{[{{\cal O}}_{C}(C_{0})]}$ determines a normalised normal form $B^{\prime}$ w.r.t. some $P_{1}^{\prime},\ldots,P_{m}^{\prime}$ which have the same orders as $w_{i}$. By corollary 3.1 there exists an operator $S\in E_{q}$ such that $A^{\prime}:=S^{-1}B^{\prime}S\in\tilde{K}((\tilde{D}^{-1}))$. Note that $W^{\prime}:=F^{\prime}\cdot S$ has support equal to $F^{\prime}$, because $S$ is a monic invertible operator. So, $(W^{\prime},A^{\prime})$ form a Schur pair of rank one. By theorems 3.2 and 3.1 this Schur pair determines a normalised commutative subring $B\subset D_{1}$ of rank one and a torsion free sheaf ${{\cal F}}\simeq\mathop{\rm Proj}\tilde{W^{\prime}}$ of rank one with vanishing cohomologies.
As it was noticed above, the modules $F$, $F^{\prime}$ and $W^{\prime}$ are isomorphic as filtered modules, and all correspondences are compatible with the scale transform. By this reason the maps $p\in X_{[{{\cal O}}_{C}(C_{0})]}\mapsto{{\cal F}}$ and ${{\cal F}}\mapsto p\in X_{[{{\cal O}}_{C}(C_{0})]}$ are mutually inverse. ∎
Remark 3.6.
This result indicates that the moduli space of spectral sheaves of rank one, i.e. sheaves with vanishing cohomologies, is an affine open subscheme of the compactified Jacobian (cf. [22], [35], [32]). We hope to cover this issue, also in the higher rank case, in future works.
Example 3.1.
Let $L=\partial^{2}+u$, $P=4\partial^{3}+6u\partial+3u^{\prime}$, where $u(x)=\sum_{k=0}^{\infty}\frac{1}{k!}u_{k}\cdot x^{k}$, be ODOs of orders 2 and 3. Then $[L,P]=0$ iff $06uu^{\prime}+u^{\prime\prime\prime}=$ 0 (see [37]). It’s easy to see that the coefficients $u_{k}$, $k\geq 3$ are uniquely determined by this equation for any choice of free parameters $u_{0},u_{1},u_{2}$. The spectral curve of these operators is given by
$P^{2}=16L^{3}+4\left(-3u_{0}^{2}-u_{2}\right)L-4u_{0}^{3}+u_{1}^{2}-2u_{0}u_{2},$
The normalised normal form of $P$ (written in G-form) with respect to $L$ is [^11] This example was calculated by V.D. Busov in his master thesis at Lomonosov MSU:
$P^{\prime}=4\partial^{3}+2u_{0}A_{2,1}\partial+u_{1}A_{1}+\frac{2u_{0}^{2}+u_{2}}{2}(-1+A_{2,1})\int.$
If we transfer $P^{\prime}$ into the matrix form (see Lemma 3.1), we get
$\tilde{P}=\psi\circ\hat{\Phi}(P)=\begin{pmatrix}&4\\ 4D^{2}&\end{pmatrix}D^{2}+\begin{pmatrix}u_{1}&2u_{0}\\ -2u_{0}D^{2}&-u_{1}\end{pmatrix}+\begin{pmatrix}0&0\\ (-2u_{0}^{2}-u_{2})D^{2}&0\end{pmatrix}D^{-2}$
So, by fixing an equation of the spectral curve, we get one-dimensional affine algebraic set in ${\mbox{\dbl A}}^{3}$ parametrising torsion free sheaves with vanishing cohomology groups.
4 Normal forms for non-commuting operators
4.1 A Newton Region of operators with the property $A_{q}(k)$
Let $P,Q$ be a pair of monic differential operators from $D_{1}$. If $[P,Q]\neq 0$, it is useful to study the normal forms of $P$ with respect to $Q$ more carefully. The well known and useful technical tool – the Newton polygon of a differential operator from the Weyl algebra - can be naturally defined in our situation and applied to such study. In this section we introduce the notion of a Newton region – a generalisation of the Newton Polygon, suitable for operators from $\hat{D}_{1}^{sym}$ satisfying conditions $A_{q}(k)$, and study its basic properties. In this paper they will be used for the proof of a commutativity criterion in section 4.3. Further study of the Newton region and of normal forms will be continued in subsequent works.
[…]
Definition 4.1.
Suppose $H\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ is a HCP from $Hcpc(k)$, $\mathop{\rm\bf ord}(H)=r$, written in the G-form:
$H=(\sum_{0\leq i<k}\sum_{0\leq l\leq d_{i}}f_{l,i;r}\Gamma_{l}A_{i}+\sum_{0<j\leq N}g_{j;r}B_{j})D^{r}$
We define the set $E(H):=\{(l,r)|\quad\exists i,f_{l,i;r}\neq 0\}$ ($E(H)=\emptyset$ if all coefficients $f_{l,i;r}$ are equal to zero).
Suppose now $H\in\hat{D}_{1}^{sym}$ is such that all homogeneous components $H_{i}$ are HCPs from $Hcpc(k)$ (for example, $H$ satisfies condition $A_{k}(q)$). We define the Newton region $NR(H)$ as the convex hull of the union $E(H):=\cup_{i}E(H_{i})$ (i.e. the region can be unbounded).
We’ll say that the point $(a,b)\in E(H_{b})\subseteq E(H)$ does not contain $A_{i}$ if the coefficients $f_{a,i;b}$ of the G-form of $H_{b}$ satisfy the following property: $f_{a,i;b}=0$ for $i>0$.
We’ll call HCP of the form $f_{l,i;r}\Gamma_{l}A_{i}D^{r}$ or $g_{j;r}B_{j}D^{r}$ as monomials (of $H$). We’ll call HCP of the form $f_{l,i;r}\Gamma_{l}A_{i}D^{r}$ as monomials corresponding to the point $(l,r)$.
Remark 4.1.
This definition slightly differs from the well known definition of the Newton polygon of an operator from the Fist Weyl Algebra $A_{1}$, since the points of the Newton region belong to the $XY$-plane where the $X$-axis stand now for powers of $x\partial$ (hence $X$ equals to $Sdeg_{A}$), and the $Y$-axis stand for the homogeneous order $\mathop{\rm\bf ord}$. Notice that the Newton Polygon of a HCP $H$ will belong to the line $Y=\mathop{\rm\bf ord}(H)$.
Definition 4.2.
Suppose $H\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ is such that all homogeneous components $H_{i}$ are HCPs from $Hcpc(k)$ (for example, $H$ satisfies condition $A_{k}(q)$). For a real pair $(\sigma,\rho)$ with $\sigma\geq 0$, $\rho>0$ we define:
$v_{\sigma,\rho}(H)=\sup\{\sigma l+\rho j|(l,j)\in E(H)\},\quad E(H,\sigma,\rho)=\{(l,j)\in E(H)|v_{\sigma,\rho}(H)=\sigma l+\rho j\},$
where we define $v_{\sigma,\rho}(H):=-\infty$ if $E(H)=\emptyset$, and $E(H,\sigma,\rho):=\emptyset$ if $v_{\sigma,\rho}(H)=\infty$ (note that the set $E(H,\sigma,\rho)$ can be empty also if $v_{\sigma,\rho}(H)<\infty$).
If $E(H,\sigma,\rho)\neq\emptyset$, we define the operator
$f_{\sigma,\rho}(H)=\sum_{(l,j)\in E(H,\sigma,\rho)}\sum_{i}f_{l,i;j}\Gamma_{l}A_{k,i}D^{j}$
which is called the homogeneous (highest) term of $H$ associated to $(\sigma,\rho)$, and the line $l_{0}:\sigma X+\rho Y=v_{\sigma,\rho}(H)$ is called the $(\sigma,\rho)$-top line.
If $E(H,\sigma,\rho)=\emptyset$, we define $f_{\sigma,\rho}(H):=0$.
Remark 4.2.
In the following discussion the top line (of a monic operator) will usually go across some vertex $(0,p)$.
Note that immediately from definition it follows that
[…]
In particular, if $H$ satisfies condition $A_{k}(0)$, then there exists $(\sigma,\rho)$ with $\sigma>0$ such that $v_{\sigma,\rho}(H)<\infty$ (e.g. $(1,1)$).
The specific basic properties of the Newton region somewhat similar to analogous properties of the Newton polygons from the paper [10] are collected in the Appendix.
Further we’ll need several statements about the top lines of operators satisfying conditions $A_{k}(0)$.
Definition 4.3.
Suppose $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ satisfies condition $A_{k}(0)$, $\mathop{\rm\bf ord}(P)=p$. A $(\sigma,\rho)$-top line which goes across $(0,p)\in E(P)$ and contains at least two vertices is called a restriction top line of $NR(P)$.
Remark 4.3.
The restriction top line is uniquely defined if it exists. To show this first note that any real pair $(\sigma,\rho)$ with $\sigma\geq 0$, $\rho>0$ is proportional to some pair $(\tilde{\sigma},1)$, and we can consider only such pairs without loss of generality.
If $(\sigma,1)$-top line is a restriction top line, then it contains the vertex $(0,p)$ and another vertex, say $(l,j)$, with $j<p$, and $\sigma l+j=p$. If $\sigma^{\prime}>\sigma$, then $(\sigma^{\prime},1)$-top line can not be a restriction top line, because $\sigma^{\prime}l+j>\sigma l+j=p$, i.e. it can not go across $(0,p)$. Thus, there exists only one pair $(\sigma,1)$ such that $(\sigma,1)$-top line is a restriction top line.
As we have noted before, the restriction top line is no longer a trivial notion. Since $Sdeg_{A}$ might go to infinity, an operator may not have restriction top line at all.
Definition 4.4.
Suppose $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ satisfies condition $A_{k}(0)$, $\mathop{\rm\bf ord}(P)=p$. If $P$ doesn’t have the restriction top line but there exists a top line $l_{0}:\sigma_{0}X+Y=p$, $\sigma_{0}>0$, such that for any $\sigma>\sigma_{0}$ the line $l:\sigma X+Y=p$ is not the top line of $N(P)$, we call this top line $l_{0}$ as the asymptotic top line.
[…]
Lemma 4.1.
Suppose $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ satisfies condition $A_{k}(0)$, $\mathop{\rm\bf ord}(P)=p$. Then only one of the following conditions holds:
  1. 1.
    $Sdeg_{A}(P)=0$.
  2. 2.
    $Sdeg_{A}(P)>0$, and $P$ has the restriction top line.
  3. 3.
    $Sdeg_{A}(P)>0$, and $P$ has the asymptotic top line.
In particular, the asymptotic top line is uniquely defined if it exists.
Proof.
Suppose $Sdeg_{A}(P)=0$. Then for any pair $(\sigma,\rho)$ with $\sigma\geq 0$, $\rho>0$ we have $v_{\sigma,\rho}(P)=\rho p$, and then, clearly, any $(\sigma,\rho)$-top line is not the restriction top line and not an asymptotic top line, because the set $E(P)$ lies on the line $X=0$.
Suppose $Sdeg_{A}(P)>0$. Then, since $P$ satisfies condition $A_{k}(0)$, the line $l:X+Y=p$ is the $(1,1)$-top line of $P$. Put
$\sigma_{0}=\inf\{\sigma|\quad\sigma X+Y=p\quad\text{is the $(\sigma,1)$-top line of $P$}\}\geq 1.$
It is well-defined (finite) since $Sdeg_{A}(P)>0$. Now consider the line $l_{0}:\sigma_{0}X+Y=p$. If there are more than one vertex on this line, then this line is the restriction top line, and if there is only one point $(0,p)$, then it is the asymptotic top line. ∎
Example 4.1.
Suppose $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$, $\mathop{\rm\bf ord}(P)=p$, satisfies condition $A_{k}(0)$ and $Sdeg_{A}(P_{\mathop{\rm\bf ord}(P)-i})=i-1$ for all $i>0$ (such condition holds for an operator $P^{\prime}$ from corollary 2.4, which comes from a generic pair of operators $P,Q\in D_{1}$).
Then it’s easy to see that $P$ doesn’t have the restriction top line, but the top line $l_{0}:X+Y=p$ is the asymptotic line.
Definition 4.5.
Suppose $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ satisfies condition $A_{k}(0)$. We define the up-edge of the Newton region of $P$ as the set
$Edg_{u}(P):=\{(a,b)\in E(P)|\quad a=Sdeg_{A}(P_{b})\mbox{\quad and\quad}\forall b^{\prime}>b\quad Sdeg_{A}(P_{b^{\prime}})<a\}.$
$Sdeg_{A}$$\mathop{\rm\bf ord}$$(0,p)$
Figure 1: The up-edge of $P$ with the asymptotic line across $(0,p)$
Lemma 4.2.
Suppose $Q\in D_{1}$ is a monic operator with $\mathop{\rm\bf ord}(Q)=\deg(Q)=q=k>0$. Suppose $P\in{D}_{1}$ has constant highest symbol (cf. theorem 2.2), $\mathop{\rm\bf ord}(P)=\deg(P)=p$. Put $P^{\prime}=S^{-1}PS$, where $S$ is a Schur operator for $Q$ from proposition 2.5. Suppose $(a,b)\in Edg_{u}(P^{\prime})$.
Then the point $(a,b)$ doesn’t contain $A_{i}$.
Proof.
By corollary 2.4 and theorem 2.2 the operator $P^{\prime}$ satisfies condition $A_{q}(0)$.
Suppose $(a,b)\in Edg_{u}(P^{\prime})$, and the coefficient at $\Gamma_{a}D^{b}$ of the G-form of $P_{b}^{\prime}$ is $t=\sum t_{i}A_{q;i}$, $t_{i}\in\tilde{K}$. Consider the operator
$\tilde{P}:=(ad\partial^{q})^{a}(P^{\prime})$
[…]
On the other hand, we know
$S^{-1}(ad(Q))^{a}(P)S=(ad(S^{-1}QS))^{a}(S^{-1}PS)=\tilde{P},$
hence we know $\bar{P}:=S\tilde{P}S^{-1}\in D_{1}$, and $\mathop{\rm\bf ord}(\bar{P}=\mathop{\rm\bf ord}(\tilde{P})$. Since $S_{0}=(S^{-1})_{0}=1$, we get $\lambda t=\bar{P}_{b+qa}=\tilde{P}_{b+qa}\in D_{1}$. But then by lemma 2.1 $t_{i}=0$ for all $i>0$, i.e. $(a,b)$ does not contain $A_{i}$. ∎
Just noting that the points on the $(\sigma,\rho)$-top line will be in $Edg_{u}(P^{\prime})$ when $\sigma,\rho>0$, we have the following Corollary.
Corollary 4.1.
In the notation of lemma 4.2 suppose $\sigma,\rho>0$. Then the points on the $(\sigma,\rho)$-top line don’t contain $A_{i}$.
In particular, if $P^{\prime}$ has the restriction top line, then the points on it don’t contain $A_{i}$.
4.2 One combinatorial lemma
[…]
$(D+L)^{k}=\sum c_{k;t_{1},\ldots,t_{m},l}L^{(t_{1})}L^{(t_{2})}\cdots L^{(t_{m})}D^{l}$
[…]
Denote by $L^{(t_{1},\ldots,t_{m})}:=L^{(t_{1})}L^{(t_{2})}\cdots L^{(t_{m})}$, and put $L^{(t_{1},\ldots,t_{m})}=1$ if $m=0$. We’ll call the index $m$ as the multiple index, and define the partial degree of $L^{(t_{1},\ldots,t_{m})}$ as
$Pdeg(L^{(t_{1},\ldots,t_{m})})=t_{1}+t_{2}+\ldots+t_{m}.$
It is easy to observe that the coefficient at $D^{k}$ in $(D+L)^{k}$ is $1$ so that it’s multiple index is 0, but except for $D^{k}$, the other terms have multiple index more than 0. Denote by $T_{i,j,k}$ the sum of monomials from the coefficient of $D^{k-i}(i>0)$ in $(D+L)^{k}$ with partial degree $Pdeg(L^{(t_{1},\ldots,t_{m})})=j\geq 0$.
Lemma 4.3.
(Combinatorial) We have
$(D+L)^{k}=D^{k}+\sum_{i=1}^{k}\sum_{j=0}^{i-1}T_{i,j,k}D^{k-i},$ (4.1)
where every monomial in $T_{i,j,k}$ has multiple index $m=i-j$, i.e.
$T_{i,j,k}=\sum_{t_{1}+\ldots+t_{m}=j\atop m=i-j}f_{i,j,k}(t_{1},\dots,t_{m})L^{(t_{1},\dots,t_{m})},$
where
$f_{i,j,k}(t_{1},\dots,t_{m})=\binom{k}{i}g(t_{1},\dots,t_{m}),$
where the function $g$ is defined by recursion:
  1. 1.
    For $m=1$ $g(t_{1})\equiv 1$.
  2. 2.
    For any $m$ with $t_{1}=\ldots=t_{m}=0$ $g(t_{1},\cdots,t_{m})=1$.
  3. 3.
    For $m>1$, when $t_{1}=0$:
    $g(0,t_{2},\ldots,t_{m})=g(t_{2},\ldots,t_{m})+g(0,t_{2}-1,\ldots,t_{m})+\ldots+g(0,t_{2},\ldots,t_{m}-1)$
  4. 4.
    For $m>1$, when $t_{1}\geq 1$:
    $g(t_{1},t_{2},\ldots,t_{m})=g(t_{1}-1,t_{2},\ldots,t_{m})+g(t_{1},t_{2}-1,\ldots,t_{m})+\ldots+g(t_{1},t_{2},\ldots,t_{m}-1),$
and we assume that $g(t_{1},t_{2},\ldots,t_{m})=0$ if $t_{i}<0$ for at least one $i$.
Proof.
The proof is by induction on $k$. When $k=1$, $(D+L)^{k}=D+L$, and it’s easy to see that $T_{1,0,1}$ satisfies all conditions in the lemma. Now suppose it is true for $k-1$, consider the generic case. Note that
$(D+L)^{k}=(D+L)(D+L)^{k-1}=(D+L)^{k-1}D+[D,(D+L)^{k-1}]+L(D+L)^{k-1},$
so that all three summands are written in standard form. By induction we have
$(D+L)^{k-1}D+[D,(D+L)^{k-1}]+L(D+L)^{k-1}=\\ D^{k}+\sum_{i=1}^{k-1}\sum_{j=0}^{i-1}T_{i,j,k-1}D^{k-i}+\sum_{i=1}^{k-1}\sum_{j=0}^{i-1}[D,T_{i,j,k-1}]D^{k-1-i}+LD^{k-1}+\sum_{i=1}^{k-1}\sum_{j=0}^{i-1}LT_{i,j,k-1}D^{k-1-i}.$ (4.2)
Note that for any $t_{1},\ldots,t_{m}$ we have $[D,L^{(t_{1},\dots,t_{m})}]=L^{(t_{1}+1,\dots,t_{m})}+\ldots+L^{(t_{1},\dots,t_{m}+1)}$, where all monomials have multiple index $m$, and $[D,T_{i,j,k-1}]\in T_{i+1,j+1,k}$. Analogously, $T_{i,j,k-1}\in T_{i,j,k}$ and $LT_{i,j,k-1}\in T_{i+1,j,k}$, where the multiple index of $LT_{i,j,k-1}$ is $i-j+1$. So, all monomials of $T_{i,j,k}$ (for arbitrary $i,j,k$) have the multiple index $i-j$ as claimed, and therefore
$T_{i,j,k}=\sum_{t_{1}+\ldots+t_{m}=j\atop m=i-j}f_{i,j,k}(t_{1},\dots,t_{m})L^{(t_{1},\dots,t_{m})},$
for some $f_{i,j,k}(t_{1},\dots,t_{m})\in K$. Let’s calculate $T_{i,j,k}$ explicitly. We can rewrite formula (4.2) as
$D^{k}+\sum_{s=1}^{k-1}\sum_{j=0}^{s-1}T_{s,j,k-1}D^{k-s}+\sum_{s=2}^{k}\sum_{j=0}^{s-2}[D,T_{s-1,j,k-1}]D^{k-s}+LD^{k-1}+\sum_{s=2}^{k}\sum_{j=0}^{s-2}LT_{s-1,j,k-1}D^{k-s}\\ =D^{k}+(T_{1,0,k-1}+L)D^{k-1}+\sum_{s=2}^{k-1}(\sum_{j=0}^{s-2}(T_{s,j,k-1}+[D,T_{s-1,j,k-1}]+LT_{s-1,j,k-1})+T_{s,s-1,k-1})D^{k-s}+\\ \sum_{j=0}^{k-2}([D,T_{k-1,j,k-1}]+LT_{k-1,j,k-1}),$
whence we get
$T_{1,0,k}=T_{1,0,k-1}+L=\binom{k}{1}L,$ (4.3)
for $\mbox{for $1<s<k$}\quad T_{s,j,k}=\begin{cases}T_{s,j,k-1}+LT_{s-1,j,k-1}&j=0\\ T_{s,j,k-1}+[D,T_{s-1,j-1,k-1}]+LT_{s-1,j,k-1}&0<j<s-1,\\ [D,T_{s-1,s-2,k-1}]+T_{s,s-1,k-1}&j=s-1\end{cases}$ (4.4)
$T_{k,j,k}=\begin{cases}LT_{k-1,0,k-1}&j=0\\ LT_{k-1,j,k-1}+[D,T_{k-1,j-1,k-1}]&0<j<k-1.\\ [D,T_{k-1,k-2,k-1}]&j=k-1\end{cases}$ (4.5)
Now for $j=0$ and $1<s<k$ we get
$T_{s,0,k}=\binom{k-1}{s}L^{(0,\ldots,0)}+\binom{k-1}{s-1}L^{(0,\ldots,0)}=\binom{k}{s}L^{(0,\ldots,0)}$
as claimed, and for $s=k$ we also get $T_{k,0,k}=L^{(0,\ldots,0)}$ as claimed.
For generic $s,j$ we have
$[D,T_{s-1,j-1,k-1}]=\sum_{t_{1}+\ldots+t_{m}=j-1\atop m=s-j}\binom{k-1}{s-1}g(t_{1},\ldots,t_{m})(L^{(t_{1}+1,\ldots,t_{m})}+\ldots+L^{(t_{1},\ldots,t_{m}+1)})=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}(g(t_{1}^{\prime}-1,t_{2}^{\prime},\ldots,t_{m}^{\prime})+\ldots+g(t_{1}^{\prime},\ldots,t_{m-1}^{\prime},t_{m}^{\prime}-1))L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=s-j}\binom{k-1}{s-1}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}(g(0,t_{2}^{\prime}-1,t_{3}^{\prime},\ldots,t_{m}^{\prime})+\ldots+g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime}-1))L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}$
and then for $1<s<k$ and $10<j<s-$ 1 we get from (4.4)
$T_{s,j,k}=\sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s}g(t_{1}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=s-j}\binom{k-1}{s-1}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}(g(0,t_{2}^{\prime}-1,t_{3}^{\prime},\ldots,t_{m}^{\prime})+\ldots+g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime}-1))L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}g(t_{2}^{\prime},t_{3}^{\prime},\ldots,t_{m}^{\prime})L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=s-j}\binom{k}{s}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k}{s}g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k}{s}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}$
as claimed. For $j=s-1$ we get $m=1$ and therefore $T_{s,s-1,k}=\binom{k-1}{s-1}L^{(s-1)}+\binom{k-1}{s}L^{(s-1)}=\binom{k}{s}L^{(s-1)}$ as claimed.
For $s=k$ and $j=k-1$ we get $m=1$ and therefore $T_{k,k-1,k}=L^{(k-1)}$ as claimed. For generic $j$ we have
$T_{k,j,k}=\sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=k-j}g(t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}+\sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=k-j}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=k-j}(g(0,t_{2}^{\prime}-1,t_{3}^{\prime},\ldots,t_{m}^{\prime})+\ldots+g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime}-1))L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=k-j}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}$
as claimed and we are done. ∎
4.3 Commutativity criterion for normal forms having the restriction top line
In this section we’ll prove a commutativity criterion for a pair of differential operators whose normal form has the restriction top line.
Before we formulate the theorem, we fix the notation and give several additional definitions. Let $(P,Q)\in D_{1}$ be a monic pair of differential operators, $Q$ is normalized, with $\mathop{\rm\bf ord}(Q)=\deg(Q)=q>0$, $\mathop{\rm\bf ord}(P)=\deg(P)=p$. Put $Q^{\prime}=S^{-1}QS=\partial^{q}$, $P^{\prime}=S^{-1}PS$, where $S$ is a Schur operator for $Q$ from proposition 2.5. By corollary 2.4 $P^{\prime}$ satisfies condition $A_{q}(0)$, i.e. in particular all its homogeneous components are totally free of $B_{j}$.
Assume $F\in K[X,Y]$ is a non-zero polynomial such that $F(P,Q):=\sum_{i,j}c_{i,j}P^{i}Q^{j}=0$. Then $F$ can be presented as a sum of $(p,q)$-homogeneous polynomials: $F=F_{1}+\ldots+F_{N}$, where
$F_{j}(X,Y):=k_{1}^{(j)}X^{u_{1}^{(j)}}Y^{v_{1}^{(j)}}+\dots+k_{m^{(j)}}^{(j)}X^{u_{m^{(j)}}^{(j)}}Y^{v_{m^{(j)}}^{(j)}},\quad k_{i^{(j)}}^{(j)}\in K,$
$N_{F_{j}}:=pu_{1}^{(j)}+qv_{1}^{(j)}=\cdots=pu_{m^{(j)}}^{(j)}+qv_{m^{(j)}}^{(j)}.$
Obviously, we have also the equation $F(P^{\prime},Q^{\prime})=0$, and since $F(P^{\prime},Q^{\prime})\in\hat{D}_{1}^{sym}$ is an operator whose homogeneous terms are HCPs from $Hcpc(q)$, this equation is equivalent to the system of infinite number of equations on coefficients of homogeneous terms of this operator, written if the G-form. Denote by $f_{l,i;r}(H)$ the coefficient of a HCP $H$ from $Hcpc(q)$ (see definition 2.3). So,
[…]
Definition 4.6.
We say the identity of type $i$ for $F_{j}$ holds if
$\sum_{1\leq l\leq m^{(j)}}\binom{u^{(j)}_{l}}{i}k_{l}^{(j)}=0.$ (4.6)
Definition 4.7.
[…]
$H_{d;(\sigma,\rho)}(L):=\sum_{\sigma l+\rho j\geq d}\sum_{i=0}^{k-1}\alpha_{l,i;j}\Gamma_{l}A_{i}\partial^{j}$
If there is no ambiguity of $(\sigma,\rho)$, we’ll simply write it as $H_{d}(L)$.
If $L\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ and all its homogeneous components $L_{i}$ are HCP from $Hcpc(q)$ in G-form, we extend definition of $H_{d;(\sigma,\rho)}(L)$ in obvious way.
Lemma 4.4.
Suppose $L,M\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ are two operators such that all homogeneous components $L_{i},M_{i}$ are HCPs from $Hcpc(q)$, suppose $(\sigma,\rho)$ is a real pair with $\sigma\geq 0$, $\rho>0$, and $v_{\sigma,\rho}(L),v_{\sigma,\rho}(M)<\infty$. Then
  1. 1.
    If $d_{1}>v_{\sigma,\rho}(L)$, then $H_{d_{1}}(L)=0$.
  2. 2.
    $H_{d}(L+M)=H_{d}(L)+H_{d}(M)$.
  3. 3.
    If $d_{1}>d_{2}$, then
    $v_{\sigma,\rho}(H_{d_{2}}(L)-H_{d_{1}}(L))\leq d_{1}$
    and
    $H_{d_{1}}[H_{d_{2}}(L)]=H_{d_{2}}[H_{d_{1}}(L)]=H_{d_{1}}(L)$
Proof.
1. $d_{1}>v_{\sigma,\rho}(L)$, then there doesn’t exist $(m,u)\in E(L)$, such that $m\sigma+u\rho\geq d_{1}$, hence $H_{d_{1}}(L)=0$.
2, 3 are obvious. ∎
Lemma 4.5.
Suppose $L,M\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ are two operators such that all homogeneous components $L_{i},M_{i}$ are HCPs from $Hcpc(q)$, suppose $(\sigma,\rho)$ is a real pair with $\sigma\geq 0$, $\rho>0$, and $v_{\sigma,\rho}(L),v_{\sigma,\rho}(M)<\infty$. Then
  1. 1.
    If $d_{1}\geq v_{\sigma,\rho}(L)$, and $d_{2}\geq v_{\sigma,\rho}(M)$, then
    $H_{d_{1}+d_{2}}(LM)=H_{d_{1}+d_{2}}[H_{d_{1}}(L)H_{d_{2}}(M)]$
  2. 2.
    Suppose $d_{1}=v_{\sigma,\rho}(L),d_{2}=v_{\sigma,\rho}(M)$. If $H_{d_{1}-\sigma}(L)$ and $H_{d_{2}-\sigma}(M)$ doesn’t contain $A_{i}$, then
    $H_{d_{1}+d_{2}-\sigma}([L,M])=H_{d_{1}+d_{2}-\sigma}([H_{d_{1}-\sigma}(L),H_{d_{2}-\sigma}(M)])$
    with
    $v_{\sigma,\rho}([L,M])\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma$
  3. 3.
    Suppose $d_{1}=v_{\sigma,\rho}(L),d_{2}=v_{\sigma,\rho}(M)$, and $v_{\sigma,\rho}([L,M])\leq d_{1}+d_{2}-\sigma$, we have for any $\epsilon>0$,
    $H_{d_{1}+d_{2}-\sigma+\epsilon}(LM)=H_{d_{1}+d_{2}-\sigma+\epsilon}(ML)$
    In particular
    $H_{d_{1}+d_{2}}(LM)=H_{d_{1}+d_{2}}(ML)$
Proof.
1. If $d_{1}>v_{\sigma,\rho}(L)$ or $d_{2}>v_{\sigma,\rho}(M)$, then $H_{d_{1}}(L)=0$ or $H_{d_{2}}=0$, and by Lemma 5.3 we know there doesn’t exist $(l,j)\in E(LM)$, such that $l\sigma+j\rho>d_{1}+d_{2}$, hence $H_{d_{1}+d_{2}}(LM)=0$. Now let’s consider the case when $d_{1}=v_{\sigma,\rho}(L)$ and $d_{2}=v_{\sigma,\rho}(M)$.
Suppose $L_{1}=H_{d_{1}}(L)$ and $M_{1}=H_{d_{1}}(M)$, put $L_{3}=L-L_{1},M_{3}=M-M_{1}$. This means for any $(m_{3},u_{3})\in E(L_{3})$ and $(n_{3},v_{3})\in E(M_{3})$:
$m_{3}\sigma+u_{3}\rho<d_{1},\quad n_{3}\sigma+v_{3}\rho<d_{2}$
Hence if there exists $(l,j)\in E(L_{1}M_{3})\bigcup E(L_{3}M_{1})\bigcup E(L_{3}M_{3})$, we have $l\sigma+j\rho<d_{1}+d_{2}$. This means $H_{d_{1}+d_{2}}(L_{1}M_{3})=H_{d_{1}+d_{2}}(L_{3}M_{1})=H_{d_{1}+d_{2}}(L_{3}M_{3})=0$. Thus
$H_{d_{1}+d_{2}}(LM)=H_{d_{1}+d_{2}}(L_{1}M_{1}+L_{1}M_{3}+L_{3}M_{1}+L_{3}M_{3})=H_{d_{1}+d_{2}}(L_{1}M_{1})$
2. Assume $L_{1}=H_{d_{1}-\sigma}(L),M_{1}=H_{d_{2}-\sigma}(M)$, put $L_{3}=L-L_{1},M_{3}=M-M_{1}$. Then $v_{\sigma,\rho}(L_{3})<d_{1}-\sigma,v_{\sigma,\rho}(M_{3})<d_{2}-\sigma$ and there doesn’t exist $(m_{3},u_{3})\in E(L_{3})$, $(n_{3},v_{3})\in E(M_{3})$, such that
$m_{3}\sigma+u_{3}\rho=d_{1}-\sigma,n_{3}\sigma+v_{3}\rho=d_{2}-\sigma$
By the same arguments as above (use Lemma 5.3 item 1) there doesn’t exist $(l,j)\in E(L_{1}M_{3})\bigcup E(L_{3}M_{1})\bigcup E(L_{3}M_{3})$, such that
$l\sigma+j\rho\geq d_{1}+d_{2}-\sigma$
Hence $H_{d_{1}+d_{2}-\sigma}(L_{1}M_{3})=H_{d_{1}+d_{2}-\sigma}(L_{3}M_{1})=H_{d_{1}+d_{2}-\sigma}(L_{3}M_{3})=0$. Thus we get
$H_{d_{1}+d_{2}-\sigma}(LM)=H_{d_{1}+d_{2}-\sigma}(L_{1}M_{1}+L_{1}M_{3}+L_{3}M_{1}+L_{3}M_{3})=H_{d_{1}+d_{2}-\sigma}(L_{1}M_{1})$
For the same reason we have
$H_{d_{1}+d_{2}-\sigma}(ML)=H_{d_{1}+d_{2}-\sigma}(M_{1}L_{1})$
So we get $H_{d_{1}+d_{2}-\sigma}([L,M])=H_{d_{1}+d_{2}-\sigma}([L_{1},M_{1}])$.
According to the assumptions, $L_{1}$ and $M_{1}$ doesn’t contain $A_{i}$, then by Lemma 5.3 item 3(a), we know
$v_{\sigma,\rho}([L_{1},M_{1}])\leq d_{1}+d_{2}-\sigma.$
Now suppose $H_{1}=H_{d_{1}+d_{2}-\sigma}([L,M])=H_{d_{1}+d_{2}-\sigma}([L_{1},M_{1}])$, and $H_{3}=H-H_{1}$. So we have $v_{\sigma,\rho}(H_{1})\leq d_{1}+d_{2}-\sigma$ and $v_{\sigma,\rho}(H_{3})\leq d_{1}+d_{2}-\sigma$, hence $v_{\sigma,\rho}([L,M])\leq d_{1}+d_{2}-\sigma$.
3. Since $v_{\sigma,\rho}([L,M])\leq d_{1}+d_{2}-\sigma$, by Lemma 4.4
$H_{d_{1}+d_{2}-\sigma+\epsilon}([L,M])=0$
Hence
$H_{d_{1}+d_{2}-\sigma+\epsilon}(LM)-H_{d_{1}+d_{2}-\sigma+\epsilon}(ML)=0$
Remark 4.4.
Compare this lemma item 2 with Lemma 5.3 item 3(a). Here we give out a more precise estimation: at that time we need $L,M$ are free of $A_{i}$, but here we only need a part of them not containing $A_{i}$.
Combining this Lemma with Lemma 4.3, we get
Corollary 4.2.
Suppose $L,M\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ are two operators such that all homogeneous components $L_{i},M_{i}$ are HCPs from $Hcpc(q)$, suppose $(\sigma,\rho)$ is a real pair with $\sigma\geq 0$, $\rho>0$, and $v_{\sigma,\rho}(L)=v_{\sigma,\rho}(M)=p$. Suppose $L,M$ satisfy the condition that
$H_{2p}([L,M])=0$
Then for any $d>0$, we have
$H_{dp}((L+M)^{d})=\sum_{l=0}^{d}\binom{d}{l}H_{dp}(M^{d-l}L^{l})$
Proof.
Apply Lemma 4.3 for $L,M$. Denote $M^{(0)}=M,M^{(1)}=[L,M],M^{(2)}=[L,[L,M]],\dots$. For $1\leq l\leq d$, since $M^{(l)}=LM^{(l-1)}-M^{(l-1)}L$, then by Lemma 4.5 item 1, we have
$H_{(l+1)p}(M^{(l)})=H_{p}(L)H_{lp}(M^{(l-1)})-H_{lp}(M^{(l-1)})H_{p}(L)$
Since $H_{2p}(M^{(1)})=[L,M]=0$, then step by step we will get $H_{(l+1)p}(M^{(l)})=0$. Suppose $(t_{1},\dots,t_{m})$ and $(i,j)$ are the corresponding index to the term
$f_{i,j,d}(t_{1},\dots,t_{m})M^{(t_{1},\dots,t_{m})}L^{d-i}$
in $(L+M)^{d}$, so by Lemma 4.3, $i-j=m$ and $j=t_{1}+\cdots+t_{m}$. If $j>0$, then at least one of $(t_{1},t_{2},\dots,t_{m})$ are not 0, we have by lemma 4.5 item 1
$H_{(j+m)p}(M^{(t_{1},\dots,t_{m})})=H_{(j+m)p}[H_{(t_{1}+1)p}(M^{(t_{1})})\times\cdots\times H_{(t_{m}+1)p}(M^{(t_{m})})]=0$
Hence for $j>0$
$H_{dp}(M^{(t_{1},\dots,t_{m})}L^{d-i})=0$
So by Lemma 4.4 item 2 and by Lemma 4.3, we have
$H_{dp}((L+M)^{d})=H_{dp}(L^{d}+\sum_{i=1}^{d}\sum_{j=0}^{i-1}\sum_{t_{1}+\ldots+t_{m}=j\atop m=i-j}f_{i,j,d}(t_{1},\dots,t_{m})M^{(t_{1},\dots,t_{m})}L^{d-i})\\ =H_{dp}(L^{d}+\sum_{i=1}^{d}\ f_{i,0,d}(0,\dots,0)M^{(0,\dots,0)}L^{d-i})$
Notice that when $j=0$,$m=i-j=i$. So $M^{(0,\dots,0)}=M^{m}=M^{i}$, and $f_{i,0,d}=\binom{d}{i}g(0,\dots,0)$, with $g(0,\dots,0)=1$. Hence
$H_{dp}((L+M)^{d})=\sum_{l=0}^{d}\binom{d}{l}H_{dp}(M^{d-l}L^{l})$
Remark 4.5.
The condition $H_{2p}([L,M])=0$ holds if
  1. 1.
    $L,M$ doesn’t contain $A_{i}$.
  2. 2.
    One of $L,M$ is $\partial^{ak},a\in\mathbb{N}$.
  3. 3.
    $\exists r\geq 0$, $H_{p-r}(L)$ and $H_{p-r}(M)$ doesn’t contain $A_{i}$.
1 can refer to Lemma 5.3 item 3(a). 2 can refer to Lemma 5.3 item 3(b). 3 can be shown by assuming $L_{1}=H_{p-r}(L),M_{1}=H_{p-r}(M)$, and arguing in the same way like in Lemma 4.5 item 3, so we omit the details here. Notice that when $r=0$ it’s also true.
For the proof of our main theorem in this section we need one more definition.
Definition 4.8.
[…]
$HS_{d;(\sigma,\rho)}^{m}(L):=\sum_{\sigma l+\rho j\geq d;l\leq m}\sum_{i=0}^{k-1}\alpha_{l,i;j}\Gamma_{l}\partial^{j}$
If there is no ambiguity of $(\sigma,\rho)$, we’ll simply write it as $HS_{d}^{m}(L)$.
If $L\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ and all its homogeneous components $L_{i}$ are HCP from $Hcpc(q)$ (in G-form), we extend definition of $HS_{d;(\sigma,\rho)}^{m}(L)$ in obvious way.
By definition,
$Sdeg_{A}(HS^{m}_{d}(L))\leq m$
and we have
Lemma 4.6.
Suppose $L,M\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ are two operators such that all homogeneous components $L_{i},M_{i}$ are HCPs from $Hcpc(q)$, suppose $(\sigma,\rho)$ is a real pair with $\sigma\geq 0$, $\rho>0$, and $d_{1}=v_{\sigma,\rho}(L),d_{2}=v_{\sigma,\rho}(M)$. Then we have:
  1. 1.
    If $d_{1}=d_{2}=d$, then
    $HS_{d}^{m}(L)+HS_{d}^{m}(M)=HS_{d}^{m}(L+M)$
  2. 2.
    For any $d$, we have
    $H_{d}(HS_{d}^{m}(L))=HS_{d}^{m}(H_{d}(L))=HS_{d}^{m}(L)$
  3. 3.
    For any $d$, $Sdeg_{A}(L)\leq a$ iff $HS_{d}^{a}(L)=H_{d}(L)$
  4. 4.
    If $Sdeg_{A}(H_{d_{1}}(L))=a_{1},Sdeg_{A}(H_{d_{2}}(M))=a_{2}$, then
    $HS_{d_{1}+d_{2}}^{a_{1}+a_{2}}(LM)=H_{d_{1}+d_{2}}(HS_{d_{1}}^{a_{1}}(L)HS_{d_{2}}^{a_{2}}(M))$
  5. 5.
    If $E(L)=\{(a_{1},b_{1})\}$, where $a_{1}\sigma+b_{1}\rho=d_{1}$, and $HS_{d_{2}}^{a_{2}}(M)=0$, then
    $HS_{d_{1}+d_{2}}^{a_{1}+a_{2}}(LM)=0$
Proof.
1, 2, 3 are by definitions.
4. By Lemma 4.5 we have $H_{d_{1}+d_{2}}(LM)=H_{d_{1}+d_{2}}(H_{d_{1}}(L)H_{d_{2}}(M))$. Hence we have
$HS_{d_{1}+d_{2}}^{a_{1}+a_{2}}(LM)=HS_{d_{1}+d_{2}}^{a_{1}+a_{2}}(H_{d_{1}+d_{2}}(LM))=HS_{d_{1}+d_{2}}^{a_{1}+a_{2}}(H_{d_{1}}(L)H_{d_{2}}(M))\\ =HS_{d_{1}+d_{2}}^{a_{1}+a_{2}}(HS_{d_{1}}^{a_{1}}(L)HS_{d_{2}}^{a_{2}}(M))=H_{d_{1}+d_{2}}(HS_{d_{1}}^{a_{1}}(L)HS_{d_{2}}^{a_{2}}(M))$
The first equality is by item 2; The second is Lemma 4.5; The last two are by item 3.
5. $HS_{d_{2}}^{a_{2}}(M)=0$, means that for any $(n,v)\in E(H_{d_{2}}(M))$ holds $n>a_{2}$. Denote $M_{0}=\Gamma_{v}D^{n}$. Then
$H_{d_{1}+d_{2}}(LM_{0})=\alpha_{a_{1},b_{1}}\beta_{n,v}H_{d_{1}+d_{2}}(\sum_{l=0}^{n}\binom{n}{l}b_{1}^{l}\Gamma_{n+a_{1}-l}D^{b_{1}+v})=\alpha_{a_{1},b_{1}}\beta_{n,v}\Gamma_{a_{1}+n}D^{b_{1}+v}$
hence $HS_{d_{1}+d_{2}}^{a_{1}+a_{2}}(LM)=0$. ∎
Theorem 4.1.
Assume $(P,Q)\in D_{1}$ be a monic pair of differential operators, $Q$ is normalized, with $\mathop{\rm\bf ord}(Q)=\deg(Q)=q>0$, $\mathop{\rm\bf ord}(P)=\deg(P)=p$. Put $Q^{\prime}=S^{-1}QS=\partial^{q}$, $P^{\prime}=S^{-1}PS$, where $S$ is a Schur operator for $Q$ from proposition 2.5.
Suppose $P^{\prime}$ has the restriction top line, then there doesn’t exist a non-zero polynomial $F\in K[X,Y]$, such that $F(P,Q)=0$.
Remark 4.6.
It can be shown that if the normal form of $P$ with respect to $Q$ has the restriction top line, then the normal form of $Q$ with respect to $P$ has the restriction top line too. We are going to clarify the details of this fact in a subsequent paper.
Proof.
Assume the converse: suppose such $F$ exists. The idea of the proof is to show that the identities of type $i$ holds for $F_{1}$ for all $i\gg 0$. This would imply $F_{1}=0$, a contradiction. [^12] The same idea works in the case of any Burchnall-Chaundy polynomials. For such polynomials it is just an easy exercise to show that theorem is true either if $P^{\prime}$ has the restriction top line or if $P^{\prime}$ has the asymptotic top line.
Arrange the vertices on the restriction top line associated to $(\sigma,1)=(p/q,1)$ as $(0,p)$, $(a_{0},b_{0})$, $(a_{1},b_{1}),\cdots,(a_{n},b_{n}),\cdots$, with $0<a_{0}<a_{1}<\cdots<a_{n}<\cdots$, the coefficient of $(a_{i},b_{i})$ is $t_{i}\in\tilde{K}$ according to Corollary 4.1. Assume $F_{1}(P,Q)=f_{p,q}(F)=k_{1}X^{u_{1}}Y^{v_{1}}+\dots+k_{m}X^{u_{m}}Y^{v_{m}}$, where $u_{1}>u_{2}>\dots>u_{m},k_{i}\neq 0$, $N_{F}=v_{p,q}(F)=u_{i}p+v_{i}q$ for all $1\leq i\leq m$. Suppose $\bar{F}=F-F_{1}$, it’s easy to see $f_{p,q}(\bar{F})\leq N_{F}-1$, so that $H_{N_{F}}(\bar{F})=0$.
Suppose $P^{\prime}=\partial^{p}+L$. Since $P^{\prime}$ has the restriction top line, we know $v_{\sigma,1}(L)=v_{\sigma,1}(P^{\prime})=p$. Denote $\mathbb{D}=\partial^{p}$, and put $\mathbb{L}=L$, $\mathbb{L}_{0}=\Gamma_{a_{0}}\partial^{b_{0}},\mathbb{L}_{1}=f_{\sigma,1}(L)-\mathbb{L}_{0},\mathbb{L}_{2}=L-\mathbb{L}_{0}-\mathbb{L}_{1}$. It’s easy to find
$p=v_{\sigma,1}(\mathbb{D})=v_{\sigma,1}(\mathbb{L}_{0})\geq v_{\sigma,1}(\mathbb{L}_{1}),\quad p\geq v_{\sigma,1}(\mathbb{L}_{2})$ (4.7)
with $H_{p}(\mathbb{L}_{2})=0$, and also
$a_{0}=Sdeg_{A}(\mathbb{L}_{0})<a_{1}$
For $d>0$, consider
$H_{pd}(P^{\prime d})=H_{pd}((\mathbb{D}+\mathbb{L}_{0}+\mathbb{L}_{1}+\mathbb{L}_{2})^{d})=H_{pd}((\mathbb{D}+\mathbb{L}_{0}+\mathbb{L}_{1})^{d})+H_{pd}(\sum_{l=1}^{d}\binom{d}{l}(\mathbb{D}+\mathbb{L}_{0}+\mathbb{L}_{1})^{d-l}\mathbb{L}_{2}^{l}),$
where the last equality follows from corollary 4.2. For any $1\leq l\leq d$, by (4.7) and by Lemma 4.5 item 1 (used $d$ times), we have
$H_{pd}((\mathbb{D}+\mathbb{L}_{0}+\mathbb{L}_{1})^{d-l}\mathbb{L}_{2}^{l})=H_{pd}[(H_{p}(\mathbb{D}+\mathbb{L}_{0}+\mathbb{L}_{1}))^{d-l}(H_{p}(\mathbb{L}_{2}))^{l}]=0$
So we have
$H_{pd}(P^{\prime d})=H_{pd}((\mathbb{D}+\mathbb{L}_{0}+\mathbb{L}_{1}+\mathbb{L}_{2})^{d})$
Since $Q^{\prime}=\partial^{q}$, we have $H_{q}(Q^{\prime})=Q^{\prime}=\partial^{q}$. For the same reason we have
$H_{N_{F}}(P^{\prime u_{j}}Q^{\prime v_{j}})=H_{N_{F}}((\mathbb{D}+\mathbb{L}_{0}+\mathbb{L}_{1})^{u_{j}}\partial^{v_{j}q})$
hence
$H_{N_{F}}(F(P^{\prime},Q^{\prime}))=H_{N_{F}}(F_{1}(P^{\prime},Q^{\prime}))=H_{N_{F}}[\sum_{j=1}^{m}k_{j}((\mathbb{D}+\mathbb{L}_{0}+\mathbb{L}_{1})^{u_{j}}\partial^{v_{j}})]$
Now use Corollary 4.2 for $L:=\mathbb{D},M:=\mathbb{L}_{0}+\mathbb{L}_{1},d=u_{j}$ (notice they satisfy the condition in Remark 4.5 for item 2). So we have for any $1\leq j\leq m$:
$H_{u_{j}p}((\mathbb{D}+\mathbb{L}_{0}+\mathbb{L}_{1})^{u_{j}})=\sum_{l=0}^{u_{j}}\binom{u_{j}}{l}(\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\mathbb{D}^{u_{j}-l}$
Thus
$H_{N_{F}}(F(P^{\prime},Q^{\prime}))=\sum_{j=1}^{m}k_{j}\cdot H_{N_{F}}(\sum_{l=0}^{u_{j}}\binom{u_{j}}{l}(\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})$ (4.8)
To find the coefficient at $\partial^{N_{F}}$ in the equation $F(P^{\prime},Q^{\prime})=0$ (so, this expression should be zero), we need to calculate $HS_{N_{F}}^{0}(F(P^{\prime},Q^{\prime}))$. Since $HS_{p}^{0}(P^{\prime})=\partial^{p}$ and $HS_{q}^{0}(Q^{\prime})=\partial^{q}$, by Lemma 4.6, we have
$HS_{N_{F}}^{0}(F(P^{\prime},Q^{\prime}))=HS_{N_{F}}^{0}(H_{N_{F}}(F(P^{\prime},Q^{\prime})))=\sum_{j=1}^{m}k_{j}\partial^{N_{F}}$
Thus we get the equation of type 0:
$\sum_{j=1}^{m}k_{j}=0$ (4.9)
Now suppose the identities of $10,1,\dots,s-$ 1 type hold, we use induction to prove the identity of type $s$, i.e
$\sum_{j=1}^{m}\binom{u_{j}}{s}k_{j}=0$ (4.10)
Note that
$H_{N_{F}}(F(P^{\prime},Q^{\prime}))=\\ \sum_{l=0}^{s-1}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot H_{N_{F}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})+\sum_{l=s}^{u_{j}}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot H_{N_{F}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})\\ =\sum_{l=s}^{u_{j}}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot H_{N_{F}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})$
To find the coefficient at $\Gamma_{sa_{0}}\partial^{N_{F}-s(p-b)}$, we need to calculate $HS_{N_{F}}^{sa_{0}}(F(P^{\prime},Q^{\prime}))$. Notice that both $\mathbb{L}_{0}$ and $\mathbb{L}_{1}$ lie on $Edg_{u}(P^{\prime})$, this means they doesn’t contain $A_{i}$, hence they satisfy the condition item 3 in Remark 4.5. Use Corollary 4.2 again for $L:=\mathbb{L}_{0},M:=\mathbb{L}_{1}$, we have
$H_{lp}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l})=H_{lp}[\sum_{h=0}^{l}\binom{l}{h}\mathbb{L}_{0}^{l-h}\mathbb{L}_{1}^{h}]$
Since we have $Sdeg_{A}(\mathbb{L}_{0})=a_{0}<a_{1}$, hence $HS_{p}^{a_{0}}(\mathbb{L}_{0})=HS_{p}^{a_{0}}(\mathbb{L})=\mathbb{L}_{0}$, hence $HS_{p}^{a_{0}}(\mathbb{L}_{1})=HS_{p}^{a_{0}}(\mathbb{L})-HS_{p}^{a_{0}}(\mathbb{L}_{0})=0$. Now we can use Lemma 4.6 item 5 (since $v_{\sigma,1}(\mathbb{L}_{1})\leq p$), i.e.
$HS_{lp}^{la_{0}}(\mathbb{L}_{0}^{l-h}\mathbb{L}_{1}^{h})=\begin{cases}0&h\neq 0\\ H_{lp}\mathbb{L}_{0}^{l}&h=0\end{cases}$
For the same reason we have
$HS_{N_{F}}^{sa_{0}}(\mathbb{L}_{0}^{l-h}\mathbb{L}_{1}^{h}\partial^{N_{F}-lp})=\begin{cases}0&h\neq 0\quad\text{or}\quad l>s\\ H_{N_{F}}(\mathbb{L}_{0}^{l}\partial^{N_{F}-pl})&h=0\quad l=s\end{cases}$
So
$HS_{N_{F}}^{sa_{0}}(F(P^{\prime},Q^{\prime}))=HS_{N_{F}}^{sa_{0}}(H_{N_{F}}(F(P^{\prime},Q^{\prime})))\\ =\sum_{j=1}^{m}\binom{u_{j}}{s}k_{j}\cdot HS_{N_{F}}^{sa_{0}}(\sum_{h=0}^{s}\binom{s}{h}\mathbb{L}_{0}^{s-h}\mathbb{L}_{1}^{h}\partial^{N_{F}-sp})+\\ \sum_{l=s+1}^{u_{j}}\sum_{j=1}^{m}\binom{u_{j}}{l}k_{j}\cdot HS_{N_{F}}^{sa_{0}}((\mathbb{L}_{0}+\mathbb{L}_{1})^{l}\partial^{N_{F}-lp})\\ =\sum_{j=1}^{m}\binom{u_{j}}{s}k_{j}\mathbb{L}_{0}^{s}\partial^{N_{F}-sp}$
Hence we get the identity for type $s$. Now we know they hold for any positive integer $i$, we have
$\binom{u_{1}}{i}k_{1}+\cdots+\binom{u_{m}}{i}k_{m}=0$
so choose $u_{2}<i\leq u_{1}$, and consider corresponding equation, we know only $u_{1}>i$, hence only one term left, and we get
$k_{1}\binom{u_{1}}{i}=0$
We get $k_{1}=0$, this is a contradiction. ∎
5 Appendix
In this section we collect all necessary basic technical assertions about the function $v_{\sigma,\rho}$ and the homogeneous highest terms $f_{\sigma,\rho}$ used in the paper.
Lemma 5.1.
Suppose $L,M\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ are two operators such that all homogeneous components $L_{i},M_{i}$ are HCPs from $Hcpc(k)$, suppose $(\sigma,\rho)$ is a real pair with $\sigma\geq 0$, $\rho>0$, and $v_{\sigma,\rho}(L),v_{\sigma,\rho}(M)<\infty$. Then
  1. 1.
    $v_{\sigma,\rho}(L+M)\leq\max\{v_{\sigma,\rho}(L),v_{\sigma,\rho}(M)\}$, and the equality holds if $v_{\sigma,\rho}(L)\neq v_{\sigma,\rho}(M)$.
  2. 2.
    If $v_{\sigma,\rho}(L)\neq v_{\sigma,\rho}(M)$, then
    $f_{\sigma,\rho}(L+M)=\begin{cases}f_{\sigma,\rho}(L)&v_{\sigma,\rho}(L)>v_{\sigma,\rho}(M)\\ f_{\sigma,\rho}(M)&v_{\sigma,\rho}(L)<v_{\sigma,\rho}(M)\end{cases}$
    so that we have $f_{\sigma,\rho}(L+M)=f_{\sigma,\rho}(f_{\sigma,\rho}(L)+f_{\sigma,\rho}(M))$ if $f_{\sigma,\rho}(L),f_{\sigma,\rho}(M)\neq 0$.
  3. 3.
    If $v_{\sigma,\rho}(L)=v_{\sigma,\rho}(M)=v_{\sigma,\rho}(L+M)$, then
    $f_{\sigma,\rho}(L+M)=f_{\sigma,\rho}(L)+f_{\sigma,\rho}(M)$
Proof.
1. Obviously, for any operator $P$ from formulation we have
$v_{\sigma,\rho}(P)\geq\sigma\max\{l|(l,j)\in E(P_{j})\}+\rho j$
[…]
$\max\{l|(l,j)\in E(L_{j}+M_{j})\}\leq\max\{\max\{l|(l,j)\in E(L_{j})\},\max\{l|(l,j)\in E(M_{j})\}\}.$
[…]
$v_{\sigma,\rho}(L)\geq\sigma\max\{\max\{l|(l,j)\in E(L_{j})\},\max\{l|(l,j)\in E(M_{j})\}\}+\rho j\geq\\ \sigma\max\{l|(l,j)\in E(L_{j}+M_{j})\}+\rho j,$
hence $v_{\sigma,\rho}(L)\geq v_{\sigma,\rho}(L+M)$.
[…]
$\sigma\max\{l|(l,j)\in E(L_{j})\}+\rho j>v_{\sigma,\rho}(L)-\varepsilon,$
and if $\varepsilon$ is sufficiently small, then $\sigma\max\{l|(l,j)\in E(L_{j})\}+\rho j>v_{\sigma,\rho}(M)$. Then for such $\varepsilon$ and $j$ we have $v_{\sigma,\rho}(L_{j}+M_{j})=v_{\sigma,\rho}(L_{j})$ and $v_{\sigma,\rho}(L+M)\geq v_{\sigma,\rho}(L_{j}+M_{j})>v_{\sigma,\rho}(L)-\varepsilon$, whence $v_{\sigma,\rho}(L)=v_{\sigma,\rho}(L+M)$.
2. It’s obvious.
3. Just by the definition of $f_{\sigma,\rho}$. ∎
Corollary 5.1.
Suppose $0\neq H\in\hat{D}^{sym}\hat{\otimes}_{K}\tilde{K}$, with all homogeneous components in $H$ are HCPs from $Hcpc(k)$, suppose $(\sigma,\rho)$ is a real pair with $\sigma\geq 0$, $\rho>0$ and $v_{\sigma,\rho}(H)<\infty$. Suppose $H_{1}=f_{\sigma,\rho}(H)$, $H_{2}=H-H_{1}$. Then one of the following is true:
  1. 1.
    $v_{\sigma,\rho}(H_{2})<v_{\sigma,\rho}(H)$;
  2. 2.
    $v_{\sigma,\rho}(H_{2})=v_{\sigma,\rho}(H)$ but $f_{\sigma,\rho}(H_{2})=0$.
Proof.
If $H_{1}\neq 0$, then $v_{\sigma,\rho}(H)=v_{\sigma,\rho}(H_{1})$, so by Lemma 5.1 item 1, we know $v_{\sigma,\rho}(H_{2})\leq v_{\sigma,\rho}(H)$. If $H_{1}=0$, the equality holds.
If $v_{\sigma,\rho}(H_{2})=v_{\sigma,\rho}(H)$, then by the definition of $H_{1}$ we know there doesn’t exist $(l,j)\in E(H_{2})$, such that $\sigma l+\rho j=v_{\sigma,\rho}(H)=v_{\sigma,\rho}(H_{2})$, so by the definition of $f_{\sigma,\rho}$ we get $f_{\sigma,\rho}(H_{2})=0$. ∎
We now want to estimate $v_{\sigma,\rho}(LM)$ and $v_{\sigma,\rho}([L,M])$ with the help of $v_{\sigma,\rho}(L)$ and $v_{\sigma,\rho}(M)$ (cf. similar estimations for $L,M\in A_{1}$ in [10, L.2.7]). We consider first the case when $L,M$ are monomials from $Hcpc(k)$.
Lemma 5.2.
Suppose $L,M\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ are two monomial operators from $Hcpc(k)$, suppose $(\sigma,\rho)$ is a real pair with $\sigma\geq 0$, $\rho>0$. Then
  1. 1.
    $v_{\sigma,\rho}(LM)=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$.
  2. 2.
    $v_{\sigma,\rho}([L,M])\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$. In the following cases we have more precise estimation:
    1. (a)
      In the case of $L$ and $M$ don’t contain $A_{i}$, then
      $v_{\sigma,\rho}([L,M])\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma;$
    2. (b)
      Suppose one of $L,M$ is $g\partial^{b}$, where $b=ck$, $c\in\mathbb{N}$, $g\in K$. Then
      $v_{\sigma,\rho}([L,M])\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma.$
Proof.
If $Sdeg_{A}(L)=-\infty$ or $Sdeg_{A}(M)=-\infty$, then $L$ or $M$ depends only on $B_{j}$, so $LM$ and $ML$ depends only on $B_{j}$ by formulae (2.6)-(2.9), and therefore $v_{\sigma,\rho}(LM)=-\infty$, $v_{\sigma,\rho}([L,M])=-\infty$, and all statements of lemma are trivial. So, we can assume below that $Sdeg_{A}(L,M)\neq-\infty$.
1. Suppose
$L=a_{i_{1},m}\Gamma_{m}A_{i_{1}}D^{u},M=a_{i_{2},n}\Gamma_{n}A_{i_{2}}D^{v}$
Then
$LM=a_{i_{1},m}a_{i_{2},n}\xi^{ui_{2}}\sum_{t=0}^{n}\binom{n}{t}u^{n-t}\Gamma_{t+m}A_{i_{1}+i_{2}}D^{u+v}+\ldots,$ (5.1)
where $\ldots$ here and below in the proof mean terms containing $B_{j}$ (although this equation may contain terms with $B_{j}$, here we are discussing $v_{\sigma,\rho}$, so we don’t have to write them out, and for convenience we will always forget about that in the following).
Hence we know $v_{\sigma,\rho}(LM)=\sup\{(l,j)\in E(LM)\}=(m+n)\sigma+(u+v)\rho=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$.
2. Since $v_{\sigma,\rho}(LM)=v_{\sigma,\rho}(ML)=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$, by lemma 5.1 we know $v_{\sigma,\rho}([L,M])\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$.
Now consider the precise estimation: If $L,M$ both don’t contain $A_{i}$, assume
$L=a_{1}\Gamma_{m}D^{u},M=a_{2}\Gamma_{n}D^{v}$
then
$\begin{cases}LM=a_{1}a_{2}\sum_{t=0}^{n}\binom{n}{t}u^{n-t}\Gamma_{t+m}D^{u+v}+\ldots\\ ML=a_{1}a_{2}\sum_{t=0}^{m}\binom{m}{t}v^{m-t}\Gamma_{t+n}D^{u+v}+\ldots\end{cases}$
Hence
$[L,M]=a_{1}a_{2}(u-v)\Gamma_{m+n-1}D^{u+v}+\cdots,$
where $\ldots$ mean terms with the value of $v_{\sigma,\rho}$ less than $(m+n-1)\sigma+(u+v)\rho=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma$. Thus $v_{\sigma,\rho}([L,M])\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma$.
If one of $L,M$ is $g\partial^{b}$, say, $L=a_{1}\Gamma_{m}A_{i}D^{u}$, $M=g\partial^{ck}$, then
$\begin{cases}LM=a_{1}g\Gamma_{m}A_{i}D^{u+ck}+\ldots\\ ML=a_{1}g\sum_{t=0}^{m}\binom{m}{t}(ck)^{m-t}\Gamma_{t}A_{i}D^{u+ck}\end{cases}$
Hence
$[L,M]=-a_{1}gmckA_{i}\Gamma_{m-1}D^{u+ck}$
Thus $v_{\sigma,\rho}([L,M])\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma$. ∎
Now we come to the general case:
Lemma 5.3.
Suppose $L,M\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ are two operators such that all homogeneous components $L_{i},M_{i}$ are HCPs from $Hcpc(k)$, suppose $(\sigma,\rho)$ is a real pair with $\sigma\geq 0$, $\rho>0$, and $v_{\sigma,\rho}(L),v_{\sigma,\rho}(M)<\infty$. Then
  1. 1.
    For any $(l,j)\in E(LM)$, there exists $(m,u)\in E(L)$ and $(n,v)\in E(M)$, such that
    $l\leq m+n,\quad j\leq u+v$
  2. 2.
    $v_{\sigma,\rho}(LM)\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$. The equality holds if one of the following case is true:
    1. (a)
      $f_{\sigma,\rho}(L)\neq 0,f_{\sigma,\rho}(M)\neq 0$, with $f_{\sigma,\rho}(L)$ and $f_{\sigma,\rho}(M)$ don’t contain $A_{i}$.
    2. (b)
      $f_{\sigma,\rho}(L)\neq 0,f_{\sigma,\rho}(M)=0$, with $f_{\sigma,\rho}(L)$ doesn’t contain $A_{i}$ and $\exists\epsilon>0$ such that all points $(l,j)\in E(M)$ with $\sigma l+\rho j>v_{\sigma,\rho}(M)-\epsilon$ don’t contain $A_{i}$.
    3. (c)
      $f_{\sigma,\rho}(L)=0,f_{\sigma,\rho}(M)\neq 0$ with $f_{\sigma,\rho}(M)$ doesn’t contain $A_{i}$ and $\exists\epsilon>0$ such that all points $(l,j)\in E(L)$ with $\sigma l+\rho j>v_{\sigma,\rho}(L)-\epsilon$ don’t contain $A_{i}$.
    4. (d)
      $f_{\sigma,\rho}(L)=0,f_{\sigma,\rho}(M)=0$, and $\exists\epsilon>0$ such that all points $(l,j)\in E(L)$ with $\sigma l+\rho j>v_{\sigma,\rho}(L)-\epsilon$ don’t contain $A_{i}$ and all points $(l,j)\in E(M)$ with $\sigma l+\rho j>v_{\sigma,\rho}(M)-\epsilon$ don’t contain $A_{i}$.
  3. 3.
    $v_{\sigma,\rho}([L,M])\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$
    In the following cases we have more precise estimation:
    1. (a)
      In the case of $L$ and $M$ don’t contain $A_{i}$, then
      $v_{\sigma,\rho}([L,M])\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma$
    2. (b)
      Suppose $M=g\partial^{n}$, where $n=mk$, $m\in\mathbb{N}$, $g\in K$. Then
      $v_{\sigma,\rho}([L,M])\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma$
Proof.
If $E(LM)=\emptyset$, there is nothing to prove. So, we can assume $E(LM)\neq\emptyset$. In this case $E(L)\neq\emptyset$ and $E(M)\neq\emptyset$, since otherwise $L$ or $M$ would contain only monomials with $B_{j}$, and then $LM$ would contain also only monomials with $B_{j}$ according to formulae (2.6)-(2.9), i.e. $E(LM)=\emptyset$, a contradiction.
1. Suppose the result is not true, hence there exists $(l_{0},j_{0})\in E(LM)$, but for any $(m,u)\in E(L)$ and $(n,v)\in E(M)$, whether $l_{0}>m+n$ or $j_{0}>u+v$ holds. Assume $L_{0},M_{0}$ are monomial elements in $L,M$, $L_{0}=a_{m,i_{1};u}\Gamma_{m}A_{i_{1}}D^{u}$, $M_{0}=a_{n,i_{2};v}\Gamma_{n}A_{i_{2}}D^{v}$ (obviously, it’s sufficient to consider only monomials corresponding to points $(m,n)$, $(u,v)$). Then like in equation 5.1 (Lemma 5.2 item 1) we have
$L_{0}M_{0}=a_{i_{1},m}a_{i_{2},n}\xi^{ui_{2}}\sum_{t=0}^{n}\binom{n}{t}u^{n-t}\Gamma_{t+m}A_{i_{1}+i_{2}}D^{u+v}+\ldots$ (5.2)
Hence for any $(l,j)\in E(L_{0}M_{0})$, $l\leq m+n$ and $j\leq u+v$. This means $(l_{0},j_{0})\notin E(L_{0}M_{0})$ for any monomials of $L,M$, so $(l_{0},j_{0})\notin E(LM)$, this is a contradiction.
2. We know $v_{\sigma,\rho}(LM)=\sup\{\sigma l+\rho j|(l,j)\in E(LM)\}$. Thus for any $\epsilon>0$, there exists $(l,j)\in E(LM)$, such that $v_{\sigma,\rho}(LM)<\sigma l+\rho j+\epsilon$. According to item 1, there exist $(m,u)\in E(L)$ and $(n,v)\in E(M)$, such that $l\leq m+n$ and $j\leq u+v$, thus we have
$v_{\sigma,\rho}(LM)<\sigma l+\rho j+\epsilon\leq(\sigma m+\rho u)+(\sigma n+\rho v)+\epsilon\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)+\epsilon$
So, we get $v_{\sigma,\rho}(LM)\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$.
Now lets discuss when the equality holds:
  1. (a)
    $f_{\sigma,\rho}(L)\neq 0,f_{\sigma,\rho}(M)\neq 0$
    Suppose $u_{0}=\sup\{u|(m,u)\in E(f_{\sigma,\rho}(L))\}$, and $v_{0}=\sup\{v|(n,v)\in E(f_{\sigma,\rho}(M))\}$. Notice that $u_{0}$ is an integer and $u_{0}\leq\frac{v_{\sigma,\rho}(L)}{\rho}$ (because $\rho>0$), so $u_{0}$ is well-defined, so does $v_{0}$. And suppose $m_{0},n_{0}$ are the corresponding integers for $u_{0}$ and $v_{0}$, such that
    $m_{0}\sigma+u_{0}\rho=v_{\sigma,\rho}(L),\quad n_{0}\sigma+v_{0}\rho=v_{\sigma,\rho}(M)$
    Hence $(m_{0},u_{0})\in E(L;\sigma,\rho)$ and $(n_{0},v_{0})\in E(M;\sigma,\rho)$. Suppose $L_{0}=a_{0,m_{0}}\Gamma_{m_{0}}D^{u_{0}},M_{0}=a_{0,n_{0}}\Gamma_{n_{0}}D^{v_{0}}$ are the monomials corresponding to the points $(m_{0},u_{0})\in E(L;\sigma,\rho)$ and $(n_{0},v_{0})\in E(M;\sigma,\rho)$ (they don’t contain $A_{i}$ according to the assumptions).
    Now put $L_{1}=f_{\sigma,\rho}(L),L_{2}=L_{1}-L_{0},L_{3}=L-L_{1}$, then for any $(m,u)\in E(L_{2})$, we have $u<u_{0}$, and for any $(m,u)\in E(L_{3})$, we have $m\sigma+u\rho<m_{0}\sigma+u_{0}\rho$. For the same we assume $M_{1}=f_{\sigma,\rho}(M),M_{2}=M_{1}-M_{0},M_{3}=M-M_{1}$, for any $(n,v)\in E(M_{2})$, we have $v<v_{0}$, and for any $(n,v)\in E(M_{3})$, we have $n\sigma+v\rho<n_{0}\sigma+v_{0}\rho$. Thus we get the decomposition
    $L=L_{0}+L_{2}+L_{3},\quad M=M_{0}+M_{2}+M_{3}$
    Consider the following equation:
    $LM=L_{0}M_{0}+L_{0}(M_{2}+M_{3})+(L_{2}+L_{3})M_{0}+(L_{2}+L_{3})(M_{2}+M_{3})$
    We want to show $(m_{0}+n_{0},u_{0}+v_{0})\in E(LM)$. This can be true if $(m_{0}+n_{0},u_{0}+v_{0})\in E(L_{0}M_{0})$, but doesn’t appear in the rest three terms:
    By formula (5.2) we know $(m_{0}+n_{0},u_{0}+v_{0})\in E(L_{0}M_{0})$.
    On the other hand, in $L_{0}M_{2}$, since for any $(n,v)\in E(M_{2})$ we have $v<v_{0}$, thus for any $(l,j)\in E(L_{0}M_{2})$ we have $j<v_{0}+u_{0}$, hence $(m_{0}+n_{0},u_{0}+v_{0})\notin E(L_{0}M_{2})$. Thus there doesn’t exist $(n,v)\in E(M_{2})$ such that
    $n_{0}\leq n,\quad v_{0}\leq v.$
    Also for $L_{0}M_{3}$, since for any $(n,v)\in E(M_{3})$, we have $n\sigma+v\rho<n_{0}\sigma+v_{0}\rho$, we also have there doesn’t exist $(n,v)\in E(M_{3})$ such that
    $n_{0}\leq n,\quad v_{0}\leq v.$
    Then according to item 1, we know $(m_{0}+n_{0},u_{0}+v_{0})\notin E(L_{0}(M_{2}+M_{3}))$, since, obviously, $E(M_{2}+M_{3})\subseteq E(M_{2})\cup E(M_{3})$. The same arguments work for $(L_{2}+L_{3})(M_{0})$ and $(L_{2}+L_{3})(M_{2}+M_{3})$. So, we get
    $(m_{0}+n_{0},u_{0}+v_{0})\notin E(L_{0}(M_{2}+M_{3}))\cup E((L_{2}+L_{3})(M_{2}+M_{3}))\cup E((L_{2}+L_{3})M_{0}).$
    Hence we have $(m_{0}+n_{0},u_{0}+v_{0})\in E(LM)$, this means
    $v_{\sigma,\rho}(LM)\geq(m_{0}+n_{0})\sigma+(u_{0}+v_{0})\rho=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$
    and together with $v_{\sigma,\rho}(LM)\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$ we get the equality.
  2. (b)
    $f_{\sigma,\rho}(L)\neq 0,f_{\sigma,\rho}(M)=0$
    It’s easy to see the equality holds iff the following is true
    $v_{\frac{\sigma}{\rho},1}(LM)=v_{\frac{\sigma}{\rho},1}(L)+v_{\frac{\sigma}{\rho},1}(M)$
    So here we may assume $\rho=1$.
    Since $f_{\sigma,\rho}(M)=0$, then for any $\epsilon>0$, there exists $(n,v)\in E(M)$, such that
    $n\sigma+v<v_{\sigma,1}(M)<n\sigma+v+\epsilon$
    So we can choose
    $v_{0}=\sup\{v|(n,v)\in E(M),n\sigma+v>v_{\sigma,1}(M)-\epsilon\},$
    where $\epsilon<\epsilon_{0}$ and $\epsilon_{0}$ is the number that all points $(n,v)\in E(M)$ with $\sigma n+v>v_{\sigma,\rho}(M)-\epsilon_{0}$ doesn’t contain $A_{i}$ as in assumption. This $v_{0}$ is well defined since $\{v|(n,v)\in E(M),n\sigma+v>v_{\sigma,1}(M)-\epsilon\}$ is a non-empty set and $v<v_{\sigma,1}(M)$ always holds. And we choose $n_{0}:=\sup\{n|(n,v_{0})\in E(M)\}$, it’s easy to see $n_{0}$ is well-defined and $(n_{0},v_{0})$ satisfies the properties:
    1. (1)
      $(n_{0},v_{0})\in E(M)$, with $v_{\sigma,\rho}(M)-\epsilon<n_{0}\sigma+v_{0}<v_{\sigma,1}(M)$
    2. (2)
      Suppose the monomial corresponding to $(n_{0},v_{0})$ is
      $M_{0}=a_{n_{0},v_{0}}\Gamma_{n_{0}}D^{v_{0}}$
      $M_{1}=\sum_{(n,v)\in E(M)|n\sigma+v>v_{\sigma,1}(M)-\epsilon}a_{n,v}\Gamma_{n}D^{v}$
      ($M_{1}$ is well-defined and it doesn’t contain $A_{i}$). Define $M_{2}=M_{1}-M_{0}$ Then for any $(n,v)\in E(M_{2})$, we have either $n\sigma+v\leq n_{0}\sigma+v_{0}$ or $v<v_{0}$.
    3. (3)
      Suppose $M_{3}=M-M_{1}$, then $v_{\sigma,1}(M_{2})<n_{0}\sigma+v_{0}$.
    Since $f_{\sigma,1}(L)\neq 0$, we can define $L_{0},L_{1},L_{2},L_{3}$ in the same way like in (a). Then again
    $LM=L_{0}M_{0}+L_{0}(M_{2}+M_{3})+(L_{2}+L_{3})M_{0}+(L_{2}+L_{3})(M_{2}+M_{3}).$
    For the same reason we know $(m_{0}+n_{0},u_{0}+v_{0})\in E(LM)$, because $(m_{0}+n_{0},u_{0}+v_{0})\in E(L_{0}M_{0})$, but doesn’t appear in the rest three parts. Thus $(m_{0}+n_{0},u_{0}+v_{0})\in E(LM)$, and
    $v_{\sigma,\rho}(LM)\geq(m_{0}+n_{0})\sigma+(u_{0}+v_{0})\geq v_{\sigma,1}(L)+v_{\sigma,1}(M)-\epsilon.$
    Together with the inequality from item 2) we get the equality.
  3. (c)
    $f_{\sigma,\rho}(L)=0,f_{\sigma,\rho}(M)\neq 0$. This case is analogous to b), so we omit the details.
  4. (d)
    $f_{\sigma,\rho}(L)=0,f_{\sigma,\rho}(M)=0$, in this case just deal with $L,M$ like in (b), the discussion will be the same, we omit the details.
3. The inequality is obvious in view of item 2.
3(a). Assume the converse, i.e. $v_{\sigma,\rho}([L,M])>v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma$, then there exist $(l,j)\in E([L,M])$, such that $l\sigma+j\rho>v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma$.
Suppose $L_{0}=a_{m,u}\Gamma_{m}D^{u},M_{0}=a_{n,v}\Gamma_{n}D^{v}$(according to the assumptions they don’t contain $A_{i}$) are the monomials in $L,M$. Using the calculation in Lemma 5.2 item 2, we have
$[L_{0},M_{0}]=a_{m,u}a_{n,v}(u-v)\Gamma_{m+n-1}D^{u+v}+\cdots$
This means for any $(l_{0},j_{0})\in E([L_{0},M_{0}])$,
$l_{0}\sigma+j_{0}\rho\leq(m+n-1)\sigma+(u+v)\rho\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma,$
but $l\sigma+j\rho>v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)-\sigma$,this means $(l,j)\notin E([L_{0},M_{0}])$ for any $L_{0},M_{0}$, Hence $(l,j)\notin E([L,M])$, a contradiction.
3(b) The arguments are the same as in 3(a), we omit the proof here. ∎
Lemma 5.4.
In the notations of lemma 5.3, if $v_{\sigma,\rho}(LM)=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$, and $f_{\sigma,\rho}(LM)\neq 0$, then we have
$v_{\sigma,\rho}[f_{\sigma,\rho}(LM)]=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$ (5.3)
On the other hand, if (5.3) holds and $v_{\sigma,\rho}(L)\neq-\infty$ and $v_{\sigma,\rho}(M)\neq-\infty$, then $f_{\sigma,\rho}(L)\neq 0,f_{\sigma,\rho}(M)\neq 0$ and $v_{\sigma,\rho}(LM)=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$.
Proof.
If $v_{\sigma,\rho}(LM)=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$ and $f_{\sigma,\rho}(LM)\neq 0$, then there exist $(l,j)\in E(f_{\sigma,\rho}(LM))\subseteq E(LM)$ such that
$\sigma l+\rho j=v_{\sigma,\rho}(f_{\sigma,\rho}(LM))=v_{\sigma,\rho}(LM)=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M).$
Assume now 5.3 holds. Then $f_{\sigma,\rho}(LM)\neq 0$ (as $v_{\sigma,\rho}(L)\neq-\infty$ and $v_{\sigma,\rho}(M)\neq-\infty$). Define $H=LM,H_{1}=f_{\sigma,\rho}(LM),H_{2}=H-H_{1}$ like in Corollary 5.1. By Corollary 5.1 we have
$v_{\sigma,\rho}(H_{2})<v_{\sigma,\rho}(H)\quad\text{or}\quad f_{\sigma,\rho}(H_{2})=0.$
By Lemma 5.1 item 1 we have
$v_{\sigma,\rho}(H_{1})\leq\max\{v_{\sigma,\rho}(H),v_{\sigma,\rho}(H_{2})\}=v_{\sigma,\rho}(H)$
In item 2, we have proved $v_{\sigma,\rho}(H)\leq v_{\sigma,\rho}(L)+v_{\sigma,\rho}(L)$ and equation 5.3 means $v_{\sigma,\rho}(H_{1})=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$. Hence we must have
$v_{\sigma,\rho}(H)=v_{\sigma,\rho}(H_{1})=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M),$
hence there exists $(l,j)\in E(H)$, such that $l\sigma+j\rho=v_{\sigma,\rho}(H)=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$. By item 1 there exist $(m,u)\in E(L)$ and $(n,v)\in E(M)$, such that $l\leq m+n,j\leq u+v$, thus
$v_{\sigma,\rho}(H)=l\sigma+j\rho\leq(m+n)\sigma+(u+v)\rho$ (5.4)
But $(m,u)\in E(L)$ and $(n,v)\in E(M)$, this means $\sigma m+\rho u\leq v_{\sigma}(L)$ and $\sigma n+\rho v\leq v_{\sigma}(M)$, hence
$v_{\sigma,\rho}(H)=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)\geq(m+n)\sigma+(u+v)\rho$ (5.5)
Comparing two equations 5.4 and 5.5, we get $m\sigma+u\rho=v_{\sigma,\rho}(L)$ and $n\sigma+v\rho=v_{\sigma,\rho}(M)$, this means $f_{\sigma,\rho}(L)\neq 0$ and $f_{\sigma,\rho}(M)\neq 0$. ∎
As a result, we have a way to calculate $f_{\sigma,\rho}(LM)$ only by $f_{\sigma,\rho}(L),f_{\sigma,\rho}(M)$ when $v_{\sigma,\rho}(LM)=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$.
Lemma 5.5.
In the notations of lemma 5.3, if $v_{\sigma,\rho}(LM)=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$, then
$f_{\sigma,\rho}(LM)=f_{\sigma,\rho}[f_{\sigma,\rho}(L)f_{\sigma,\rho}(M)].$
Proof.
Assume first $f_{\sigma,\rho}(L)\neq 0,f_{\sigma,\rho}(M)\neq 0$. Put $L_{1}=f_{\sigma,\rho}(L)\neq 0$, $M_{1}=f_{\sigma,\rho}(M)\neq 0$, put $L_{3}=L-L_{1}$ and $M_{3}=M-M_{1}$. Consider the equation
$H=LM=L_{1}M_{1}+L_{1}M_{3}+L_{3}M_{1}+L_{3}M_{3}$
For $L_{3},M_{3}$, we have 4 possibilities:
  1. (1)
    $v_{\sigma,\rho}(L_{3})<v_{\sigma,\rho}(L_{1})$,$v_{\sigma,\rho}(M_{3})<v_{\sigma,\rho}(M_{1})$
  2. (2)
    $v_{\sigma,\rho}(L_{3})<v_{\sigma,\rho}(L_{1})$,$v_{\sigma,\rho}(M_{3})=v_{\sigma,\rho}(M_{1})$, but $f_{\sigma,\rho}(M_{3})=0$.
  3. (3)
    $v_{\sigma,\rho}(L_{3})=v_{\sigma,\rho}(L_{1})$,$v_{\sigma,\rho}(M_{3})<v_{\sigma,\rho}(M_{1})$, but $f_{\sigma,\rho}(L_{3})=0$.
  4. (4)
    $v_{\sigma,\rho}(L_{3})=v_{\sigma,\rho}(L_{1})$,$v_{\sigma,\rho}(M_{3})=v_{\sigma,\rho}(M_{1})$, but $f_{\sigma,\rho}(L_{3})=f_{\sigma,\rho}(M_{3})=0$.
For (1), we know $v_{\sigma,\rho}(L_{1}M_{3})\leq v_{\sigma,\rho}(L_{1})+v_{\sigma,\rho}(M_{3})<v_{\sigma,\rho}(L_{1})+v_{\sigma,\rho}(M_{1})$. By Lemma 5.1 item 2, we have $f_{\sigma,\rho}(L_{1}M_{1}+L_{1}M_{3})=f_{\sigma,\rho}(L_{1}M_{1})$, analogously for $L_{3}M_{1}$ and $L_{3}M_{3}$. We get
$f_{\sigma,\rho}(H)=f_{\sigma,\rho}(L_{1}M_{1})$
For (2), $f_{\sigma,\rho}(M_{3})=0$ means for any $(n_{3},v_{3})\in E(M_{3})$ $\sigma n_{3}+\rho v_{3}<v_{\sigma,\rho}(M_{1})$. We need the following claim:
Claim: There doesn’t exist $(l,j)\in E(L_{1}M_{3})$, such that $l\sigma+j\rho\geq v_{\sigma,\rho}(L_{1})+v_{\sigma,\rho}(M_{1})$.
(Proof of the Claim) Assume the converse, then by item 1, there exist $(m_{1},u_{1})\in E(L_{1})$ and $(n_{3},v_{3})$, such that $l\leq m_{1}+n_{3}$ and $j\leq u_{1}+v_{3}$, but we know $m_{1}\sigma+u_{1}\rho\leq v_{\sigma,\rho}(L_{1})$ and $\sigma n_{3}+\rho v_{3}<v_{\sigma,\rho}(M_{1})$, this is a contradiction.
So this claim shows that $v_{\sigma,\rho}(L_{1}M_{3})<v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$ or $v_{\sigma,\rho}(L_{1}M_{3})=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$, but $f_{\sigma,\rho}(L_{1}M_{3})=0$. Like in (1) we can check $v_{\sigma,\rho}(L_{3}M_{1})<v_{\sigma,\rho}(L_{1}M_{1})$, $v_{\sigma,\rho}(L_{3}M_{3})<v_{\sigma,\rho}(L_{1}M_{1})$. So we get again $f_{\sigma,\rho}(H)=f_{\sigma,\rho}(L_{1}M_{1})$.
Cases (3) and (4) are analogous, we omit the details.
If at least one of $f_{\sigma,\rho}(L)$ and $f_{\sigma,\rho}(M)=0$, then the above arguments show there doesn’t exist $(l,j)\in E(LM)$ such that $l\sigma+j\rho=v_{\sigma,\rho}(LM)=v_{\sigma,\rho}(L)+v_{\sigma,\rho}(M)$, hence $f_{\sigma,\rho}(LM)=0$. ∎